### v1

# RAS: Retrieval-And-Structuring for Knowledge-Intensive LLM Generation

Abstract
Retrieval-augmented language models often struggle with knowledge-intensive tasks due to inefficient retrieval, unstructured knowledge integration, and single-pass architectures. We present Retrieval-And-Structuring (RAS), a novel framework that dynamically constructs and reasons over query-specific knowledge graphs through iterative retrieval and structuring. RAS introduces four key technical innovations: (1) a theme-scoped retrieval mechanism that efficiently narrows the search space while maintaining retrieval quality, (2) an action planning module that determines knowledge needs and generates focused sub-queries, (3) a dynamic knowledge structuring approach that converts retrieved text into an evolving knowledge graph, and (4) a graph-augmented answering component that leverages the accumulated structured information. Our framework achieves state-of-the-art performance, surpassing leading baselines by 6.4% with open-source language models and 7.0% with proprietary models on seven knowledge-intensive generation datasets across all evaluation metrics. Detailed ablation studies verify the contribution of each technical component to the overall system performance.
1 Introduction
Complex reasoning tasks like scientific analysis or multi-hop question answering demand both comprehensive knowledge and structured logical thinking. While large language models (LLMs) have achieved remarkable performance across natural language processing tasks (Devlin et al., 2018; Brown et al., 2020; Chowdhery et al., 2023), they often struggle with knowledge-intensive reasoning due to the absence of precise, logically structured information (Rae et al., 2021; Ling et al., 2024). This limitation has prompted growing research into augmenting LLMs with structured knowledge to enhance their reasoning capabilities (Wang et al., 2021).
Retrieval-augmented generation (RAG) approaches have emerged as a promising direction by providing LLMs with additional context from retrieved passages (Guu et al., 2020; Lewis et al., 2020; Izacard & Grave, 2021; He et al., 2024). However, these methods frequently face hallucination challenges (Maynez et al., 2020; Zhang et al., 2023b), where generated content deviates from retrieved information. This issue stems primarily from the unstructured nature of retrieved passages, which lack explicit logical connections needed for precise reasoning. Recent efforts have explored integrating knowledge graphs (KGs) with LLMs (Sun et al., 2019; Yu et al., 2022; He et al., 2024; Edge et al., 2024), as KGs offer compact, structured representations of entities and relationships that enable more accurate and interpretable reasoning (Hogan et al., 2021; Jiang et al., 2024).
We present Retrieval-And-Structuring (RAS), a novel framework that dynamically constructs and reasons over query-specific knowledge graphs through iterative retrieval and structuring. RAS integrates theme-aware retrieval, dynamic KG construction, and self-refinement (Ji et al., 2023; Asai et al., 2023; Lyu et al., 2024b) to comprehensively prepare knowledge for challenging questions. The framework comprises four key components: (1) an action planner that strategically determines knowledge needs and generates targeted sub-queries, (2) a theme-scoped retrieval module that efficiently combines thematic and semantic relevance, (3) a knowledge structuring module that builds and expands query-specific KGs, and (4) a knowledge-augmented answerer that leverages the gathered structured information.
RAS addresses several critical limitations of existing approaches. First, unlike traditional RAG methods that rely solely on semantic similarity (Lewis et al., 2020; Guu et al., 2020), RAS employs theme-aware retrieval that significantly reduces the search space while preserving retrieval quality. Second, in contrast to static KG approaches (He et al., 2024; Edge et al., 2024), RAS dynamically constructs and evolves KGs based on the specific reasoning requirements of each query. Third, through its iterative planning mechanism, RAS actively identifies knowledge gaps and initiates focused retrievals, moving beyond the single-pass limitations of existing methods (Izacard & Grave, 2021).
On seven knowledge-intensive tasks including open-domain QA, closed-set QA, and long-form generation, RAS outperforms state-of-the-art baselines by 6.4% with open-source LLMs and 7.0% with proprietary models. The main contributions of this work:
We introduce RAS, a novel framework that dynamically constructs query-specific knowledge graphs through iterative retrieval and structuring for knowledge-intensive language generation.
We propose theme-scoped retrieval that combines thematic classification with dense retrieval to efficiently narrow the search space while maintaining retrieval quality.
We develop a dynamic knowledge structuring approach that iteratively builds and enriches knowledge graphs based on reasoning needs.
We validate RAS’s effectiveness through comprehensive experiments and ablation studies across diverse knowledge-intensive tasks.
2 Related Work
Retrieval-Augmented Generation. Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020) enhances performance on knowledge-intensive tasks by incorporating retrieved passages into input of large language models (LLMs), improving factual accuracy and grounding. Traditional approaches retrieved a fixed number of passages once before generation (Shao et al., 2023; Es et al., 2024; Lyu et al., 2024a) or focused on adaptive retrieval (Jiang et al., 2023), where models dynamically query external knowledge when confidence is low, or analyze on the retrieval (Kim et al., 2024b). However, these methods often fail to perform complex or multi-hop reasoning process due to the linear and simple designs, where several recent methods have emerged to improve retrieval quality. Iterative retrieval-generation approaches (Shao et al., 2023; Guan et al., 2024) leverage historical context to enhance subsequent retrievals. Other work introduces targeted retrieval strategies through subqueries (Khattab et al., 2023) or follow-up questions (Yao et al., 2023; Press et al., 2022), progressively enriching the context for more factually grounded generation. SelfRAG (Asai et al., 2023) introduced self-reflective RAG, which allows LLMs to retrieve content on demand and critically evaluate retrievals, demonstrating both the importance of retrieval quality and the capacity of LLMs to seek self-directed information. RPG (Lyu et al., 2024b) further refined this approach by extracting fine-grained paragraphs from retrieved content to improve query relevance. However, existing methods face key limitations. They often either include irrelevant content in the retrieved context or risk discarding essential information through refinement. Additionally, dense retrieval approaches can be computationally expensive when deployed on large corpora, hampering real-time user interaction. To address these challenges, we propose converting retrieved text into knowledge-intensive graphs and maintaining an iteratively enriched, query-specific knowledge graph for knowledge serving. We also introduce theme-scoped retrieval for more focused and efficient context retrieval.
Graph as Context for LLMs. The integration of graphs into modern language models represents a promising research direction (Yasunaga et al., 2021, 2022; Yu, 2022; Ju et al., 2022; Zhu et al., 2024; Gutiérrez et al., 2024). Recent approaches leverage graph-based knowledge representation to enhance LLM performance. For instance, GraphToken (Perozzi et al., 2024) shows that LLM can understand a graph encoded as a token, after a simple p-tuning (Liu et al., 2021) while keeping the LLM’s parameters frozen. Based on the findings, G-Retriever (He et al., 2024) constructs subgraphs by retrieving relevant entities and relations from a global KG, using these as structured knowledge input for LLMs. GraphRAG (Edge et al., 2024) employs an alternative strategy, extracting a comprehensive global KG from the corpus, identifying meaningful clusters through graph community detection, and generating LLM-based summaries of these communities to create a refined retrieval corpus. These methods demonstrate how graphs, with their knowledge-intensive modality and inherent structural properties, can significantly improve LLMs’ contextual understanding. However, current approaches treat graphs as static data structures without refinement based on LLM requirements. Moreover, building global KGs for large-scale corpora remains prohibitively expensive and often impractical – Wikipedia alone contains over 30 million documents per dump. The key challenge lies in developing methods to dynamically construct and utilize graphs on-the-fly for knowledge-intensive generation. To address these limitations, we propose a retrieval-and-structuring paradigm that dynamically builds and maintains query-specific KGs on-the-fly. Our approach iteratively enriches these graphs through LLM’s self-evaluation of information completeness and targeted retrieval via focused subqueries.
3 RAS Framework
Knowledge-intensive language generation requires not just retrieving relevant information, but also structuring and reasoning over it effectively. We propose Retrieval-And-Structuring (RAS), a framework shown in Figure 1. In the following subsections, we detail the technical mechanisms underlying each stage. To facilitate understanding, we showcase a running example of RAS in Figure 26.
3.1 Action Planning
The Action Planning stage is the first critical step in the RAS framework, guiding the retrieval and reasoning process based on the input query and the current knowledge state. It involves two key components: the Action Planner , a language model that assesses the sufficiency of the current knowledge state and generates sub-queries for further retrieval if necessary, and a decision step that determines the next action based on the planner’s output.
(For initial iteration) Formally, given an input query , the Action Planner generates a plan according to:
where can take one of two forms:
[NO_RETRIEVAL]: If determines that can be answered directly without requiring any additional knowledge, the planning process terminates, and the framework proceeds directly to the final Answer Generation stage (§3.4).
[SUBQ] : If assesses that the query cannot be satisfactorily answered with its own knowledge, we start the iteration with the main query as the initial subquery, and move to the next stage (§3.2).
(For iteration 0) At each iteration , the Action Planner reassesses the cumulative retrieved knowledge and the chain of previous sub-queries in the context of the main query :
where is a graph neural network for encoding the evolving KG ; is the subquery at iteration , and is the extracted graph information (triple list) from the retrieved context. The base model we use here is GraphLLM, an architecture derived from previous work (Perozzi et al., 2024; He et al., 2024). We discuss more details in §3.3.
The planner’s output at each iteration can be either:
[SUFFICIENT]: The accumulated knowledge is deemed sufficient to comprehensively address the main query . The iterative retrieval process terminates, and the framework proceeds to the Knowledge-Augmented Generation stage (§3.4).
[SUBQ] : The planner generates a new subquery to guide the retrieval of additional relevant knowledge. The subquery is designed to fill specific gaps in the current knowledge state with respect to answering . The framework continues the iterative process and proceeds to the Theme-Scoped Retrieval stage (§3.2).
The Action Planner serves as a key driver of the RAS framework’s iterative retrieval and refinement process. By dynamically assessing the adequacy of the retrieved knowledge and generating targeted sub-queries, it enables the efficient acquisition of query-relevant information.
3.2 Theme-Scoped Retrieval
This stage efficiently retrieves thematically and semantically relevant context for subquery through a two-step process: (1) theme scoping to identify a narrowed sub-corpus close to ’s thematic distribution, and (2) dense retrieval to find the most relevant passages within . This thematic pre-filtering before semantic matching aims to improve both retrieval efficiency and precision.
3.2.1 Theme Scoping
The Theme Scoping mechanism transforms sub-queries into thematic probability distributions and retrieves relevant document subsets through three key operations:
First, a theme classifier maps the subquery to a posteriori probability distribution (PPD) over thematic classes (Eq. 3). The classifier is trained on DBPedia-298 (Lehmann et al., 2015)111Document label space of DBPedia-298: https://mappings.dbpedia.org/server/ontology/classes/, a multilabel wiki text classification dataset, enabling fine-grained thematic distinctions.
To address the distribution shift between queries and documents, an MLP-based shifter transforms the query theme distribution into an expected document theme distribution (Eq. 4). This shifter is trained on query-document pairs in our constructed dataset HotpotQA-SUBQ (see Appendix B), learning to capture the relationship between query and document thematic patterns.
Finally, using the shifted distribution , we retrieve a theme-specific sub-corpus from the theme-indexed corpus (Eq. 5). The function selects the top- documents based on L2 similarity between and the pre-computed theme distributions in . This focused retrieval process effectively reduces the search space while maintaining thematic relevance for subsequent semantic matching operations.
3.2.2 Dense Retrieval
Within the theme-specific sub-corpus , a dense retriever is employed to identify the most semantically relevant passages for the subquery :
where selects the top- passages based on the cosine similarity between the query embedding and the passage embeddings precomputed in the dense index.
3.3 Knowledge Structuring and Enrichment
This stage is designed to transform the retrieved text passages into a structured graph representation and iteratively expand the query-specific knowledge graph .
3.3.1 Text-to-Triple Conversion
To capture the essential facts from the retrieved passages , a text-to-triples model is employed. We train the model on the full set of WikiOFGraph dataset (Kim et al., 2024a), an LLM-curated high-quality text-triples dataset, enabling it to generate quality and information-rich triples in the format:
where each triple represents a subject-predicate-object fact extracted from the text. This structured representation allows for efficient downstream reasoning and simple external knowledge graph integration.
3.3.2 Iterative Knowledge Enrichment
The extracted triples are then converted into a graph structure , where and denote the sets of nodes and edges, respectively. Each node corresponds to a unique subject or object entity in , while each edge represents a predicate connecting two entities. To enrich the graph with semantic information, the attributes of nodes and edges are obtained through a text encoder (Sentence-BERT (Reimers, 2019) in our case):
These semantic embeddings enable the model to capture the nuanced relationships between entities and facilitate reasoning over the knowledge graph.
To progressively expand the knowledge base in response to the evolving sub-queries, the structured graph at each iteration is merged into a evolving knowledge graph specific to the main query :
After enriching with the new knowledge, the action planner (§3.1) reassesses the current knowledge to determine the next step. Based on and the chain of previous subqueries and their associated graph information, the planner decides whether to generate another focused subquery for additional retrieval (§3.2) or to proceed with answering (§3.4) if the accumulated knowledge is deemed sufficient.
3.4 Knowledge-Augmented Generation
The final stage generates an answer to the main query using the Answerer model conditioned on knowledge graph and subquery chain .
If no retrieval is needed (), the answer is generated directly:
Otherwise, after iterative knowledge enrichment concludes with [SUFFICIENT] plan, the answer is generated using encoded KG and subquery chain:
attends to relevant knowledge in and subquery chain to generate accurate, coherent answers.
Note: In this study, and are two roles played by the same model with different instructions (see §C.3).
4 Experiments
Training Data & Setting. We develop HotpotQA-SUBQ, a dataset derived from HotpotQA (Yang et al., 2018), to train our model’s action planning and answering capabilities. Our dataset creation begins with document filtering: using Claude 3.5 Sonnet (Anthropic, 2024), we identify and retain only the supporting documents necessary for answering the main query, removing irrelevant content. For each supporting document , we then iteratively generate a subquery , considering the main query, previous subqueries, and supporting documents. During iteration , when more supporting documents remain, we create training samples with input and output label “[SUBQ] ”, where represents triples extracted from document , and is the main query. For the final supporting document, we label the sample as “[SUFFICIENT]”. To identify queries that can be answered directly, we test our base LLM (LLaMA-2-7B) on HotpotQA’s main queries without context. For correctly answered queries, we create training samples with the main query as input and “[NO_RETRIEVAL]” as the output label. Additionally, to ensure fair comparison with existing approaches, we incorporate the subset of Arc-Easy (2,147 samples) and ASQA (3,897 samples) from SelfRAG’s training data, resulting in 208k training samples in total. We place detailed training data processing, dataset statistics, and data samples in Appendix B. We use DBPedia-298, HotpotQA-SUBQ, and WikiOFGraph to train our theme classifier , distribution shifter , and text-to-triples model , respectively, as detailed in Appendix C. We present our hyperparameter study of each component in Appendix G.
Knowledge Sources. Due to the large sizes of wiki corpora, we employ faiss (Douze et al., 2024) for efficient vector searching across both theme and dense indices. Following the SelfRAG (Asai et al., 2023), we utilize the Wikipedia 2018 (Izacard et al., 2023) by default, while specifically using the Wikipedia 2020 for PopQA to access more recent information. To optimize retrieval efficiency, we partition the index into five segments. Within each segment, we first perform theme scoping to reduce the candidate document pool to 250,000 documents by default. However, we skip theme scoping for PopQA since our classifier , trained on DBPedia-298 data based on old Wiki dump, performs poorly on newer Wiki 2020 dump, producing near-zero vectors for roughly one-third of recent content.
Test Datasets & Metrics & Compared Baselines. We conduct comprehensive evaluations on diverse knowledge-intensive tasks following previous studies (Asai et al., 2023; Lyu et al., 2024b). The evaluation encompasses three categories of datasets: (1) open-domain short-form generation datasets: TriviaQA (Joshi et al., 2017), PopQA (Mallen et al., 2022), and 2WikiMultihopQA (Ho et al., 2020); (2) closed-set task datasets: PubHealth (Zhang et al., 2023a) and ARC-Challenge (Clark et al., 2018); and (3) long-form generation datasets: ALCE-ASQA (Gao et al., 2023; Stelmakh et al., 2022) and ELI5 (Fan et al., 2019). For evaluation metrics, we maintain consistency with prior work (Asai et al., 2023; Mallen et al., 2022; Lyu et al., 2024b), employing “golden match” accuracy for PopQA and TriviaQA, token-level F1 score for 2WikiMultihopQA, accuracy for PubHealth and ARC-Challenge, and ROUGE-LSum alongside MAUVE score (Pillutla et al., 2021) for ASQA and ELI5. Our comparative analysis includes three baseline categories: models without retrieval augmentation, incorporating Claude 3.5 Sonnet as a state-of-the-art closed-source baseline; models with single retrieval over top-5 documents, including Claude 3.5 Sonnet, and SuRe (Kim et al., 2024b), a leading retrieve-and-summarize framework; and models with self-reflective retrieval, including leading approaches Self-RAG (Asai et al., 2023) and RPG (Lyu et al., 2024b). We place more details of metrics/datasets in Appendix D.
| Short-form | Closed-set | Long-form Generation | ||||||||
| Model | TQA | 2WQA | PopQA | Pub | ARC | ASQA | ELI5 | |||
| (acc) | (F1) | (acc) | (acc) | (acc) | (rouge) | (mauve) | (rouge) | (mauve) | ||
| Baselines without retrieval | ||||||||||
| Llama2 | 30.5 | 18.9 | 14.7 | 34.2 | 21.8 | 15.3 | 19.0 | 18.3 | 32.4 | |
| Alpaca | 54.5 | – | 23.6 | 49.8 | 45.0 | 29.4 | 61.7 | – | – | |
| Llama2 | 38.5 | 20.2 | 14.7 | 29.4 | 29.4 | 12.4 | 16.0 | 18.2 | 41.4 | |
| Alpaca | 61.3 | – | 24.4 | 55.5 | 54.9 | 32.0 | 70.6 | – | – | |
| ChatGPT | 74.3 | 24.8 | 29.3 | 70.1 | 75.3 | 36.2 | 68.8 | 22.8 | 32.6 | |
| Sonnet-3.5 | 78.4 | 40.0 | 30.2 | 83.7 | 88.5 | 37.0 | 39.1 | 21.8 | 26.5 | |
| Baselines with retrieval (#docs=5 by default) | ||||||||||
| Llama2 | 42.5 | 21.0 | 38.2 | 30.0 | 48.0 | 22.1 | 32.0 | 18.6 | 35.3 | |
| Alpaca | 64.1 | – | 46.7 | 40.2 | 48.0 | 33.3 | 57.9 | – | – | |
| Llama2 | 47.0 | 31.2 | 45.7 | 30.2 | 26.0 | 20.5 | 24.7 | 18.6 | 42.3 | |
| Alpaca | 66.9 | – | 46.1 | 51.1 | 57.6 | 36.7 | 56.6 | – | – | |
| RECOMP (Xu et al., 2023) | – | 32.4 | – | – | – | 36.5 | 76.0 | – | – | |
| SuRe (Kim et al., 2024b) | – | 20.6 | – | – | – | 35.8 | 76.2 | – | – | |
| SuRe | 72.3 | 38.1 | 53.6 | 57.2 | 79.6 | 36.0 | 74.2 | – | – | |
| ChatGPT | 65.7 | 32.8 | 50.8 | 54.7 | 75.3 | 39.9 | 79.7 | 20.6 | 57.2 | |
| Sonnet-3.5 | 69.1 | 41.9 | 51.5 | 49.1 | 88.6 | – | – | – | – | |
| Sonnet-3.5 | 72.5 | 53.7 | 57.3 | 53.9 | 87.1 | 38.8 | 61.6 | 20.2 | 32.3 | |
| SuRe | 76.8 | 37.6 | 41.2 | 62.8 | 91.6 | 30.2 | 69.9 | 15.4 | 27.2 | |
| Baselines with self-reflective retrieval | ||||||||||
| SELF-RAG (Asai et al., 2023) | 66.4 | 25.1 | 54.9 | 72.4 | 67.3 | 35.7 | 74.3 | 17.9 | 35.6 | |
| SELF-RAG | 69.3 | – | 55.8 | 74.5 | 73.1 | 37.0 | 71.6 | – | – | |
| RPG (Lyu et al., 2024b) | – | 33.6 | 56.0 | 73.4 | – | 37.6 | 84.4 | 19.1 | 46.4 | |
| Retrieval-And-Structuring (Ours) | ||||||||||
| RAS | RAS | 77.4 | 57.7 | 62.3 | 71.4 | 93.8 | 39.1 | 70.5 | 23.3 | 37.7 |
| RAS | RAS | 73.8 | 43.4 | 61.3 | 75.5 | 62.4 | 37.2 | 95.2 | 19.7 | 47.8 |
| RAS | RAS | 78.7 | 54.7 | 58.8 | 70.8 | 91.2 | 39.1 | 70.5 | 23.3 | 37.7 |
| RAS | RAS | 72.6 | 42.1 | 58.3 | 74.8 | 67.7 | 37.2 | 95.2 | 19.7 | 47.8 |
[…]
| RAS | RAS | 73.8 | 44.2 | 57.7 | 77.5 | 71.5 | 37.6 | 96.2 | 20.1 | 54.4 |
Inference Setting. We evaluate RAS using both our trained open-source model (RAS with LLaMA-2-7B/LLaMA-3-8B and Graph Transformer (Shi et al., 2020) encoder) and the closed-source Claude-3.5-Sonnet model under varied inference settings. For RAS, we maintain consistency with previous work (Asai et al., 2023; Lyu et al., 2024b) by employing zero-shot inference across all datasets. For closed-source models (e.g., RAS), we utilize few-shot inference with two exemplars for ASQA and ELI5, while maintaining zero-shot for the other datasets.
During inference, if the initial planner output is not [NO_RETRIEVAL], we utilize the main query as the initial subquery . For PopQA and TriviaQA evaluation, we follow established settings (Asai et al., 2023; Luo et al., 2023), incorporating top-five web search engine results as initial retrieved context . For ASQA and ELI5, we maintain methodological consistency with prior work (Lyu et al., 2024b; Asai et al., 2023; Gao et al., 2023), utilizing their predetermined five-document context. Under these conditions, we omit plan generation and text retrieval phases, implementing static inference (Asai et al., 2023) with five fixed iterations. For remaining datasets, we establish a maximum iteration count of five. Following previous studies, we employ Contriever-MS MARCO (Izacard et al., 2021) as the primary dense retriever, with BM25 (Robertson et al., 2009) serving as the retrieval mechanism for 2WikiMultihopQA. Across all retrieval processes, we maintain a consistent top- document selection of five. We present more details of inference settings in Appendix E.
| TQA | 2WQA | Pub | ASQA | ||
| (acc) | (F1) | (acc) | (rg) | (mv) | |
| RAS | 72.6 | 42.1 | 74.8 | 37.2 | 95.2 |
| Training Phase | |||||
| No GraphToken | 70.2 | 38.4 | 66.4 | 33.1 | 85.0 |
| No LoRA | 71.5 | 37.8 | 54.8 | 32.8 | 84.8 |
| No Text-to-Triple | 70.4 | 38.2 | 71.4 | 36.2 | 73.8 |
| No Mutli-Task | 68.6 | 39.2 | 65.5 | 36.7 | 88.9 |
| Inference Phase | |||||
| No Retrieval | 56.9 | 27.4 | 69.0 | 31.3 | 70.6 |
| No GraphToken | 68.8 | 38.7 | 67.3 | 36.5 | 93.6 |
| No Planning | 66.7 | 37.8 | 71.5 | – | – |
| No ThemeScope | 72.0 | – | 73.9 | – | – |
4.1 Main Results
Our performance evaluation, presented in Table 1, demonstrates that the LLaMA-2-7B/LLaMA-3-8B model fine-tuned with RAS outperforms existing open-source solutions, including SelfRAG (Asai et al., 2023) and RPG (Lyu et al., 2024b). Notably, compared to the previous SOTA models, RAS shows a 9.7% improvement in short-form question-answering and a 7.9% gain in long-form generation tasks. Additionally, when applied to closed-source Sonnet-3.5 model, RAS consistently achieves superior results compared to single retrieval RAG approaches, including retrieve-and-summarize approach SuRe (Kim et al., 2024b).
We find that sometimes (e.g., on TriviaQA and PubHealth) single-hop retrieval could not boost LLM’s performance, and even makes it worse, which demonstrates the necessity of on-demand retrieval, aligning with previous findings.
Our “role-swapping” study, where Sonnet-3.5 and RAS alternately serve as action planner and answerer, reveals that performance is primarily bounded by the answerer’s inherent capabilities. For example, when Sonnet-3.5 acts as the planner with RAS as the answerer, the system achieves 62.4% accuracy on the ARC-Challenge dataset. However, when Sonnet-3.5 serves as both planner and answerer, accuracy increases to 93.8%. Notably, despite having 60x fewer parameters, our trained RAS performs comparably to Sonnet-3.5 as a planner, and even outperforms it sometimes.
Also, for long-form generation task, our trained model achieved notably high MAUVE scores, which can be majorly attributed to our “text-to-triple” generation step, as demonstrated by our later ablation study in Table 2.
4.2 Ablations & More Analysis
We study the effect of individual component in Table 2.
Effect of Iterative Planning and Retrieval. Comparing the base model with “No Planning” variant shows that iterative planning provides consistent improvements across all metrics (e.g., +8.8% on TQA, +9.0% on 2WQA). This demonstrates the importance of dynamically determining retrieval needs and generating focused sub-queries. Without planning, the model relies on single-pass retrieval, which may miss crucial information needed for complex reasoning. Also, when turning off retrieval, the performance degradation is more obvious, due to the knowledge-intensive nature of those datasets.
Effect of Graph Construction and Encoding. The impact of structured knowledge representation is evident from two ablations. First, “No Text-to-Triple” degrades performance significantly on all metrics (e.g., -9.0% on 2WQA, -22.2% MAUVE on ASQA), showing the value of essential information extraction via converting retrieved text into structured triples. Second, removing the graph encoder (“No GraphToken”) during training or inference consistently hurts performance across datasets, with particularly large drops on PubHealth (-11.2% and -10.0% respectively). This suggests that the graph structure encoding helps the model better leverage the knowledge relationships.
Effect of LoRA and Multi-task Learning. Our experiments reveal that parameter-efficient training strategies significantly impact model performance. Using only graph token optimization without LoRA leads to substantial degradation (-11.8% on average). A similar observation can be made for “No Multi-Task” (e.g., 12.3% accuracy degradation on PubHealth), indicating the significance of jointly training the model on both action planning and answer generation tasks rather than optimizing for each task separately, supporting findings from prior work (Lyu et al., 2024b). The complementary effects suggest that while graph-based knowledge representation is valuable, it needs to be combined with careful parameter tuning and multi-task learning to achieve optimal performance.
Effect of Theme Scoping. The theme-scoped retrieval shows mixed results – while removing it (“No ThemeScoping”) has minimal performance drops on TQA and PubHealth, the efficiency benefits are substantial as shown in Figure 2. This suggests theme scoping successfully narrows the search space without compromising performance.
Efficiency Analysis. Figure 2 compares the computational costs of different retrieval strategies. Theme-scoped RAG reduces processing time by approximately 60% compared to dense-only RAG by first filtering the corpus through lightweight theme classification (with 298 class dimensions). This combination of comparable performance with significantly reduced computational cost demonstrates the effectiveness of our theme-scoped retrieval approach.
Impact of Training Data Volume. Figure 3 demonstrates how training dataset size influences model performance across different tasks. Considering the computational efficiency, we sampled 2,000 instances each from TQA and 2WQA for evaluation, while maintaining the original sizes of other datasets. The results indicate that model performance generally improves with increased training data volume. Notably, our model achieves competitive performance even with limited training data - using only 5% (10K instances) of our full dataset, it surpasses previous state-of-the-art models on TQA, 2WQA, and ELI5. These results suggest both the robustness of our architectural design and the effectiveness of our data curation methodology. However, we observed an exception with the ELI5 dataset, where performance were inconsistent. This irregularity can be attributed to the inclusion of ASQA training data in our training set, following established setting from previous research (Asai et al., 2023; Lyu et al., 2024b). Among our test datasets, ASQA and ELI5 are unique in requiring long-form response generation. The periodic decline in ELI5 performance suggests that the model’s response generation began to align more closely with ASQA’s training data distribution, potentially at the expense of ELI5’s distinct characteristics.
Impact of Graph Information Abundance. Figure 4 demonstrates a strong positive correlation between graph information abundance and model performance across all tasks. Both RAS and RAS show consistent improvements as graph information increases, demonstrating RAS’s strong capability of effective utilizing available graph information for knowledge-intensive generation.
5 Conclusion
We presented RAS (Retrieval-And-Structuring), a novel framework that enhances knowledge-intensive language generation through iterative retrieval and knowledge structuring. Our approach introduces key innovations including theme-scoped retrieval, action planning, dynamic knowledge structuring, and graph-augmented answering. Comprehensive experiments demonstrate RAS’s superior performance over existing approaches across diverse benchmarks. Looking ahead, RAS’s graph-based architecture enables seamless integration with external knowledge graphs and human-in-the-loop curation, making it particularly valuable for specialized domains requiring high accuracy and reliability. We discuss limitations of RAS and future work in Appendix A. Our code can be found at https://github.com/pat-jj/Retrieval-And-Structure.
Impact Statements
The RAS framework has several potential positive impacts on both research and practical applications. First, by improving knowledge-intensive language generation through structured knowledge representation and iterative reasoning, RAS could enhance the reliability and factual accuracy of AI systems in critical domains like education, research assistance, and technical documentation. The theme-scoped retrieval mechanism also offers computational efficiency benefits, potentially reducing the energy consumption and computational resources required for large-scale knowledge retrieval tasks.
However, there are important considerations regarding potential risks and challenges. The framework’s effectiveness in knowledge structuring could potentially amplify existing biases in training data and knowledge sources. Additionally, while RAS improves retrieval efficiency, it may create barriers for low-resource languages or domains where structured knowledge bases and thematic classifications are limited. These concerns necessitate careful consideration of data sources and regular evaluation of the system’s impact across different user groups and application contexts.
Appendix A Limitations & Future Work
While RAS demonstrates strong performance, several key limitations warrant consideration for future improvements:
Knowledge Extraction Quality: The effectiveness of both planning and answer generation depends on the quality of text-to-triple extraction. Despite using a Flan-T5-based model for triple extraction, implementing a more sophisticated approach – such as a dedicated Named Entity Recognition and Relation Extraction (NER-RE) pipeline – could potentially enhance RAS’s reasoning capabilities and knowledge representation.
Theme Scoping Limitations: The current classifier was trained on DBPedia-298, an older dataset that may not adequately represent the thematic distribution of contemporary document collections. This temporal gap particularly affects performance when processing newer corpora, as evidenced by our experiments with Wikipedia 2020 dumps. The development of a huge, up-to-date multilabel document classification dataset would likely improve the effectiveness of theme scoping and, consequently, the overall efficiency gains it provides.
Graph Evolution Strategy: The current approach to evolving knowledge graphs could be enhanced by incorporating more sophisticated graph pruning and merging strategies. Future work could explore dynamic graph summarization techniques to maintain the most relevant information while preventing excessive graph growth during iterations.
Cross-Domain Adaptability: While RAS performs well on the evaluated tasks, its effectiveness across highly specialized domains or multilingual settings remains to be investigated. Future research could focus on developing domain-adaptation techniques and multilingual knowledge structuring approaches.
Interactive Refinement: The current implementation lacks mechanisms for interactive refinement of the constructed knowledge graphs. Future versions could incorporate human feedback loops to improve the quality of knowledge structuring and reasoning paths.
These limitations suggest several promising directions for future research, including improved knowledge extraction techniques, more robust theme classification approaches, and enhanced graph evolution strategies. A particularly promising direction is the integration of external knowledge graphs into RAS’s iterative loop, as its graph-based architecture naturally accommodates the incorporation of established KGs like Wikidata or domain-specific knowledge bases. This could significantly enhance the system’s reasoning capabilities by combining dynamically constructed knowledge with curated structured information. We summarize the potential future directions of extension in Figure 5.
Appendix B HotpotQA-SUBQ Dataset
The source of our dataset is HotpotQA (Yang et al., 2018).222We use the version hotpot_train_v1.1 from https://github.com/hotpotqa/hotpot. Our goal is to create a dataset that can let the model learn the capabilities of subquery generation and answering with intensive knowledge. We name the dataset as HotpotQA-SUBQ.
The construction of HotpotQA-SUBQ includes three steps: (1) filter out irrelevant/unnecessary “supporting docs” (§B.1), (2) generate one subquery for each supporting document (§B.2), and (3) sample labeling (§B.3).
B.1 Supporting Document Filtering
To identify and filter out irrelevant supporting documents from HotpotQA, we employ an instruction-based approach using Claude 3.5 Sonnet, as shown in Figure 6.
The filtering prompt is designed with clear examples that illustrate the criteria for document relevance. For each HotpotQA sample, we enumerate all supporting documents and use Claude to identify only those that contribute directly to answering the question. After filtering out irrelevant documents, we filter the question with no supporting documents. This filtering step reduces noise in the training data and helps focus the model on truly relevant information during sub-query generation.
B.2 Sub-query Generation
For sub-query generation, we use the template as follows to generate one subquery per (filtered) document:
The prompt template enforces document-specificity, goal-orientation, and conciseness in sub-query generation. For iterative querying, we maintain a list of previously generated sub-queries to avoid redundancy and encourage progressive information gathering.
B.3 Sample Labeling
The sample labeling process transforms the filtered and subquery-augmented HotpotQA examples into training instances for both the Action Planner and Answerer components. We formalize this process in Algorithm 1.
The algorithm takes as input the filtered HotpotQA dataset , base language model , and text-to-triple conversion model . For each example, we first attempt direct answer generation using the base LLM (Line 3). If successful, we create [NO_RETRIEVAL] training instances for both the planner and answerer (Lines 4-6).
For examples requiring retrieval, we process each supporting document sequentially:
1.
Convert document text to structured triples using (Line 12)
2.
For intermediate documents ():
Construct input by concatenating previous subquery-graph pairs
Label with [SUBQ] and next subquery
3.
For final document ():
Include all accumulated context
Label planner sample as [SUFFICIENT]
Create answerer training instance with ground truth answer
The input formatting function (Lines 14 and 17) follows the template:
This process generates two datasets:
: Trains the Action Planner to determine retrieval needs and generate targeted subqueries
: Trains the Answerer to synthesize final responses from accumulated graph context
Table 3 illustrates the distribution of our generated datasets. The planner dataset demonstrates a balanced distribution across three label types: 35% [SUFFICIENT], 55% [SUBQ], and 10% [NO_RETRIEVAL]. The answerer dataset incorporates both no-retrieval and sufficient cases from multiple sources. The size of the answerer dataset exceeds the combined total of no-retrieval and sufficient cases from the planner dataset for two reasons: first, we incorporated additional data from ASQA and Arc-Easy datasets; second, we included no-retrieval cases that were initially filtered out in step 1 of our process. This comprehensive approach ensures a robust training set for the answerer component.
Our labeling approach ensures that models learn both the iterative nature of complex question answering and the importance of structured knowledge representation. The complementary training objectives help develop robust reasoning capabilities while maintaining retrievability of supporting evidence.
B.4 Dataset Statistics
| Planner Data | Answerer Data | ||
| # Queries | 129,902 | 78,164 (w/ 3,897 ASQA & 2,147 Arc-Easy) | |
| # Input Tokens | (Mean, Median, Max) | (338, 301, 1,910) | (475, 466, 2,214) |
| # Output Tokens | (Mean, Median, Max) | (13, 12, 62) | (13, 4, 2,332) |
| # Subqueries | (Min, Mean, Max) | (0, 0.8, 5) | (0, 1.2, 6) |
| # [SUFFICIENT] | 45,722 | – | |
| # [SUBQ] | 71,676 | – | |
| # [NO_RETRIEVAL] | 12,504 | – | |
| # [SUBQ] Tokens | (Min, Mean, Max) | (6, 18.3, 62) | – |
| # Nodes | (Mean, Median, Max) | (11.6, 11, 56) | (11.7, 11, 84) |
| # Edges | (Mean, Median, Max) | (21.2, 20, 116) | (21.3, 20, 160) |
B.5 Data Examples
Appendix C Training Details
C.1 Classifier & Shifter Training for Theme Scoping
Classifier Training. For theme classification, we trained a multilabel document classifier on the DBPedia-298 dataset (Lehmann et al., 2015) using a two-stage architecture. The first stage uses the nomic-embed-text-v1 encoder model (Nussbaum et al., 2024) to generate document embeddings, followed by a classification head consisting of a two-layer neural network. The training size and the validation size are 240,932 and 36,008, respectively.
The classifier consists of:
An encoder based on nomic-embed-text-v1 for generating document embeddings
A classification head with:
Input layer: embedding dimension → 512 units
ReLU activation
Dropout (p=0.1)
Output layer: 512 → number of labels (298)
Sigmoid activation for multilabel outputs
We jointly trained both the encoder and classification head with different learning rates:
Classification head: Trained with a higher learning rate (2e-4) to allow faster adaptation to the classification task
Text Encoder: Fine-tuned with a lower learning rate (2e-5) to preserve pretrained knowledge while making subtle adjustments for the domain
This dual learning rate strategy allows the model to benefit from the pretrained knowledge in the encoder while optimizing the classification head for the specific multilabel task. The lower learning rate for the encoder helps prevent catastrophic forgetting of the pretrained representations.
The model was evaluated using micro-averaged metrics on the validation set during training, with monitoring of F1 score, precision, and recall, as shown in Figure 16. We select the best classifier model based on the best validation F1 score achieved during training.
Distribution Shifter Training. The theme distribution shifter maps query theme distributions to expected document theme distributions. Training involves two phases:
Phase 1: PPD Labeling
Using our constructed HotpotQA-SUBQ dataset (in Appendix B) with 204,638 query-doc pairs (where we set training-validation ratio as 9:1):
Input Generation: Process all subqueries and supporting documents through the trained theme classifier to obtain their Posterior Probability Distributions (PPDs)
Pairing: Each subquery is paired with its corresponding supporting document, creating query-document PPD pairs
Data Structure: The labeled dataset consists of:
Input: Query PPDs (298-dimensional probability vectors)
Target: Corresponding document PPDs (298-dimensional probability vectors)
Phase 2: Shifter Training
Train a neural network to map query PPDs to document PPDs:
Architecture: A three-layer MLP with:
Input layer: 298 → 512 units
Hidden layer: 512 → 256 units
Output layer: 256 → 298 units
ReLU activations and dropout (p=0.2) between layers
Softmax activation for final output
Training Objectives: Multiple distribution alignment metrics:
KL divergence loss (primary objective)
Jensen-Shannon divergence monitoring
Wasserstein distance monitoring
L1 and L2 distance monitoring
Training Configuration:
Optimizer: Adam with learning rate 1e-3
Batch size: 32
Early stopping based on validation loss
Numerical stability ensured through epsilon addition (1e-15)
Performance is tracked using validation KL divergence, Jensen-Shannon divergence, and Wasserstein distance to ensure proper distribution alignment, as shown in Figure 16.
C.2 Text-to-Triples Model Training
For our text-to-triple conversion model, we fine-tuned a Flan-T5-Large model (Chung et al., 2024) to transform raw text passages into structured knowledge triples. The model processes input text sequences up to 512 tokens in length and generates structured triples in a standardized format “(S subject P predicate O object)”. 333Our fine-tuned can be found at https://huggingface.co/pat-jj/text2triple-flan-t5 and https://huggingface.co/pat-jj/text2graph-llama-3.2-3b
Our training dataset WikiOFGraph (Kim et al., 2024a), curated by ChatGPT, comprises 5,851,776 text-triple pairs, with an additional 5,000 samples reserved for validation. Each training instance consists of a natural language text passage paired with its corresponding set of comma-separated triples. For example:
We implemented the training using the AdamW optimizer with a learning rate of 2e-5, incorporating linear warmup and decay schedules. The training process ran for 500,000 steps with a batch size of 32 per GPU and gradient accumulation over 4 steps. To optimize training efficiency and memory usage, we employed mixed-precision training using bfloat16 and applied weight decay at 0.01. The maximum source and target sequence lengths were set to 512 and 256 tokens respectively.
For evaluation, we primarily relied on ROUGE-1, ROUGE-2, and ROUGE-L scores to assess the quality of triple generation. We supplemented these metrics with custom triple matching accuracy measures that consider subject matching, predicate normalization, and object entity alignment. Validation metrics were computed at 5,000-step intervals throughout training.
The training curves for various metrics are shown in Figure 16, demonstrating steady improvement in the model’s ability to extract structured knowledge from text. The training of this model was conducted on eight NVIDIA RTX 6000 with 48GB memory.
C.3 Action Planner & Answerer Multi-task Training
We employ GraphLLM (Perozzi et al., 2024; He et al., 2024) as our foundational model architecture and utilize our constructed HotpotQA-SUBQ dataset to train a unified model capable of performing both graph-conditioned action planning and answer generation through multitask learning.
Since the Action Planner and Answerer share an identical architectural foundation, as illustrated in Figure 1, we differentiate their functionality through specialized instruction sets. For the Action Planner , we employ the following instruction template:
For the Answerer , we utilize this distinct instruction set:
For graph encoding, we implement Graph Transformer (Shi et al., 2020), selected for its robust capability in handling non-fully-connected graphs—a common characteristic of text-extracted triples. Our base language models comprise LLaMA-2-7B and LLaMA-3-8B, chosen both to maintain consistency with previous research (Asai et al., 2023; Lyu et al., 2024b) and to investigate our framework’s performance scaling across different model capacities.
Our implementation of GraphLLM differs from G-Retriever (He et al., 2024) primarily due to the distinct nature of our graph structures. While G-Retriever operates on single interconnected graphs, our framework processes multiple potentially disconnected subgraphs, each corresponding to different subqueries. To address this architectural difference, we adopt a sequential encoding strategy: rather than encoding the entire graph at once, we process each subgraph individually using Graph Transformer, followed by mean pooling across all subgraph embeddings to produce the final encoded representation.
The training process utilizes 4 NVIDIA RTX 6000 Ada Generation GPUs, each with 48GB memory. We train all models for 2 epochs using a batch size of 2 and gradient accumulation steps of 2, implementing a peak learning rate of 1e-5 with a 0.15 warmup ratio and 0.01 decay rate. The maximum sequence lengths are set to 300 tokens for generation and 2,500 tokens for input, with training conducted in BFloat16 precision.
Appendix D Evaluation Datasets & Metrics
D.1 Test Datasets
We evaluate RAS on diverse benchmark datasets spanning short-form QA, closed-set tasks, and long-form generation in the zero-shot setting, aligning with (Asai et al., 2023; Lyu et al., 2024b). Below we describe each dataset:
Short-form Generation Datasets:
TriviaQA-unfiltered (TQA) (Joshi et al., 2017): A large-scale QA dataset containing 11,313 question-answer pairs in our test set. The questions are sourced from trivia enthusiasts and cover diverse topics.
2WikiMultiHopQA (2WQA) (Ho et al., 2020): A multi-hop question answering dataset (with 12,576 samples in test set) that requires models to combine information from multiple Wikipedia articles to answer questions.
PopQA (Mallen et al., 2022): A dataset focusing on questions about long-tail entities, containing 1,399 queries where the monthly Wikipedia page views are less than 100. These questions test models’ ability to handle queries about less common entities.
Closed-set Tasks:
PubHealth (Pub) (Zhang et al., 2023a): A dataset for verifying health-related claims. Models must classify statements as ”true” or ”false” based on scientific evidence. The dataset contains 987 test samples.
ARC-Challenge (ARC) (Clark et al., 2018): A multiple-choice science question dataset designed to be challenging for models, requiring multi-hop reasoning and background knowledge. The dataset contains 1,172 test samples.
Long-form Generation Datasets:
ASQA (Stelmakh et al., 2022): A long-form question answering dataset that requires models to generate comprehensive answers with proper citation of sources. Models must balance information completeness with factual accuracy. The dataset contains 948 test samples.
ELI5 (Fan et al., 2019): A dataset derived from the ”Explain Like I’m Five” subreddit, containing questions seeking straightforward explanations of complex topics. The dataset contains 1,000 test samples.
D.2 Evaluation Metrics
We employ different evaluation metrics appropriate for each task category:
Short-form Generation:
For PopQA and TriviaQA, we use the ”golden match” metric (Asai et al., 2023; Min et al., 2019; Guu et al., 2020)444Implementation aligned with https://github.com/AkariAsai/self-rag/blob/main/retrieval_lm/metrics.py, where a prediction is considered correct if it contains any normalized version of the ground truth answers :
(12) where normalizes text by lowercasing, removing articles and punctuation.
For 2WikiMultiHopQA, we follow RPG (Lyu et al., 2024b)555Implementation aligned with https://github.com/haruhi-sudo/RPG/blob/main/retriever/src/evaluation.py to use token-level F1 score between prediction and ground truth :
(13) where precision and recall are computed over normalized token overlap:
(14) (15)
Closed-set Tasks: For both PubHealth and ARC-Challenge, we use accuracy (Asai et al., 2023), computed as:
where is the total number of examples.
Long-form Generation: For ASQA and ELI5, we use multiple metrics666Implementation aligned with https://github.com/princeton-nlp/ALCE/blob/main/eval.py:
All scores are reported as percentages, multiplied by 100.
Appendix E Inference & Evaluation Details
We evaluated RAS using both closed-source and open-source language models. For closed-source evaluation, we used Claude-3.5-Sonnet (Sonnet-3.5) as our base model. For open-source evaluation, we tested with both LLaMA-2-7B and LLaMA-3-8B, as shown in the performance table (Table 1).
For the open-source models, we first fine-tuned the GraphLLM architecture on our HotpotQA-SUBQ dataset (see Section C.3). We then conducted zero-shot knowledge-intensive generation tests. With Claude-3.5-Sonnet, we used few-shot prompting for both action planning and answer generation phases. Below, we detail our approach for each model type.
E.1 Closed-Source Model Settings
For Text-to-Triple Conversion (Figure 20), we instruct the model to extract structured knowledge triples following specific formatting rules. The prompt precisely defines the triple format as and provides comprehensive guidelines to ensure consistent knowledge representation. Key rules include extracting maximal meaningful relationships, maintaining original entity casing, avoiding pronoun usage, and ensuring clear predicate specification. The example demonstrates the conversion of a biographical text into structured triples, showing how complex information about an individual’s life, career, and temporal details can be systematically decomposed into atomic facts. This standardized format enables reliable knowledge accumulation and reasoning in subsequent stages of the RAS framework.
For Action Planning (Figure 21), we design a comprehensive prompt that guides the model in making strategic decisions about information retrieval needs. The prompt instructs the model to output one of three labels based on careful analysis of the current knowledge state: (1) [NO_RETRIEVAL] when the question can be answered directly, either due to the model’s inherent knowledge or the question’s nature, (2) [SUBQ] accompanied by a focused subquery when additional information is needed, or (3) [SUFFICIENT] when the accumulated knowledge is adequate to answer the question. The prompt includes diverse examples demonstrating different scenarios:
(1) Generating follow-up queries based on partial information (2) Creating new queries when relevant information is missing (3) Recognizing when accumulated information is sufficient (4) Identifying questions that don’t require external knowledge (5) Handling common knowledge questions
A key feature of the prompt is its emphasis on query efficiency - it explicitly prohibits generating redundant subqueries that might retrieve already-known information. This design helps maintain the system’s efficiency while ensuring comprehensive knowledge gathering.
For Answer Generation (Figure 22), we design a prompt that focuses on synthesizing precise answers from structured knowledge graphs. The prompt emphasizes selective use of retrieved information - instructing the model to utilize subqueries and graph information only when relevant to the question at hand. A distinctive feature of this prompt is its requirement for definitive answers even under uncertainty or incomplete information, ensuring the model always provides a response.
The prompt includes carefully selected examples demonstrating two key scenarios: direct fact retrieval (Nobel Prize winner) and complex reasoning across multiple knowledge pieces (athlete’s sport classification). These examples illustrate how to effectively combine information from multiple subqueries and their associated graph information to construct accurate answers. The prompt strictly enforces concise answer generation, requiring only the essential information without additional explanation or commentary.
In addition, for ASQA and ELI5 with closed-source models (RAS and all the other baselines (e.g., SuRe) using Sonnet-3.5), we conduct few-shot inference with two in-context learning demonstrations, aligning with previous study’s (Gao et al., 2023) implementation.777https://github.com/princeton-nlp/ALCE/blob/main/run.py888Exemplars can be found at https://github.com/princeton-nlp/ALCE/tree/main/prompts
E.2 Open-Source Model Settings
For inference with open-source model, we use our GraphLLM trained by Hotpot-SUBQ (see Appendix C.3), we use 8 NVIDIA RTX 6000 Ada Generation with 48GB memory and CUDA version 12.4. We use Python 3.10, PyTorch 2.1.2, and torch-geometric 2.6.1 throughout all experiments.
For more efficient text retrieval, we split both theme index and dense index of the corpus into eight splits, and load them on eight GPUs with faiss-gpu 1.7.2.999All the Wikipedia dumps (2018, 2020) we used are downloaded from https://github.com/facebookresearch/atlas
We set maximum new tokens as 100 for PopQA, TriviaQA, and 2WikiMultihopQA, as 50 for PubHealth and ARC-Challenge, and as 300 for ASQA and ELI5, aligning with previous study (Asai et al., 2023; Lyu et al., 2024b). All the generation are configured with batch size of 1.
As we always set the main query as the first subquery, we pre-extract the triples from the context at the first iteration using our Flan-T5-based text-to-triple model for batch triple extraction. Notably, when we experimented with using Claude 3.5 Sonnet for triple extraction using the prompt from Figure 20, we observed an additional 3% average performance gain. This suggests room for improvement in our triple extraction approach - one potential direction would be training a larger model like LLaMA-2-7B on the WikiOFGraph dataset to achieve higher-quality triple extraction.
E.3 Instructions
We use the following task-specific instructions for zero-shot inference:
| Dataset | Instruction |
| ARC-C (baseline) | Given four answer candidates, A, B, C and D, choose the best answer choice. Please answer with the capitalized |
| alphabet only, without adding any extra phrase or period. | |
| ARC-C | Which is true? Output A, B, C, or D. |
| PubHealth (baseline) | Is the following statement correct or not? Say true if it’s correct; otherwise, say false. Don’t capitalize or add |
| periods, just say “true” or “false”. | |
| PubHealth | Is statement ’true’ or ’false’? Only output ’true’ or ’false’. |
| ASQA (baseline) | Instruction: Write an ACCURATE, ENGAGING, and CONCISE answer for the given question using the retrieved |
| graph information (some of which might be irrelevant). Use an unbiased and journalistic tone. | |
| ASQA | Answer the following question. The question may be ambiguous and have multiple correct answers, |
| and in that case, you have to provide a long-form answer including all correct answers. [Long Form] | |
| ELI5 (baseline) | Instruction: Write an ACCURATE, ENGAGING, and CONCISE answer for the given question using the retrieved |
| graph information (some of which might be irrelevant). Use an unbiased and journalistic tone. | |
| ELI5 | Provide a paragraph-length response using simple words to answer the following question. [Long Form] |
Appendix F Qualitative Examples
In this section, we show some qualitative examples and running examples of RAS.
Appendix G Hyperparameter Study
| Hyper-parameter | Studied Values |
| Theme Classifier | |
| Encoder learning rate | {1e-6, 1e-5, 2e-5, 1e-4, 2e-4} |
| Classifier learning rate | {1e-6, 1e-5, 2e-5, 1e-4, 2e-4, 5e-4, 1e-3} |
| Batch size | {4, 8, 16, 32, 64} |
| Hidden layers | {1, 2, 3} |
| Hidden dim | {256, 512, 1024} |
| Dropout rate | {0.1, 0.2, 0.3} |
| Distribution Shifter | |
| Batch size | {4, 8, 16, 32, 64} |
| Learning rate | {0.0001, 0.0005, 0.001, 0.005} |
| Hidden layers | {[256], [512], [512, 256], [1024, 512, 256]} |
| Dropout rate | {0.1, 0.15, 0.2, 0.25, 0.3} |
| Optimizer | {Adam, SGD, RMSprop, AdamW} |
| Loss function | {KLDiv, JSD, MSE, Wasserstein, CrossEntropy} |
| Text-to-Triples Model | |
| Batch size | {2, 4, 8, 16, 32} |
| Learning rate | {1e-5, 2e-5, 5e-5, 1e-4, 2e-4} |
| GraphLLM (Action Planner & Answerer) | |
[…]
| GNN architecture | {GCN, GAT, Graph Transformer} |
| Hidden dimension | {512, 768, 1024, 2048} |
| Number of layers | {2, 3, 4, 5} |
| Number of heads | {4, 6, 8, 12} |
| Dropout rate | {0.05, 0.1, 0.2, 0.3} |
| Projector intermediate dimension | {1024, 2048, 4096} |
| Projector output dimension | 4096 |
[…]
| LoRA rank (r) | {4, 8, 16, 32} |
| LoRA alpha | {8, 16, 32} |
| LoRA dropout | {0.01, 0.05, 0.1, 0.15} |
[…]
| Learning rate | {1e-6, 2e-6, 5e-5, 1e-4} |
| Batch size (training) | {2, 4, 8, 16, 32} |
| Batch size (inference) | 1 |
| Weight decay | {0.001, 0.01, 0.05, 0.1} |
| Gradient accumulation steps | {2, 4, 8, 16} |
| Gradient clipping | {0.1, 0.3, 0.5, 1.0} |
| Warmup ratio | {0.05, 0.1, 0.15, 0.2} |
| Max text length | 2500 |
| Max new tokens | Task-specific (see Appendix E.2) |
| Others | |
| Theme scoping top- | {100,000, 250,000, 500,000 } |
| Dense retrieval top- | 5 |
Appendix H Notation Table
| Notation | Description |
| General | |
| Main input query | |
| Generated answer | |
| Complete document corpus | |
| Theme-indexed document corpus | |
| Action Planning | |
| Action planner model | |
| Plan generated at iteration | |
| Subquery generated at iteration | |
| Graph information (triple list) extracted at iteration | |
| Theme-Scoped Retrieval | |
| Theme classifier function | |
| Theme distribution shifter function | |
| Dense retriever function | |
| Posteriori probability distribution of query | |
| Expected document theme distribution | |
| Theme-specific sub-corpus at iteration | |
| Retrieved text passages at iteration | |
| Number of documents selected in theme scoping | |
| Number of passages selected in dense retrieval | |
| Knowledge Structuring | |
| Text-to-triple conversion model | |
| Text encoder for node/edge attributes | |
| Graph neural network encoder | |
| Knowledge graph at iteration | |
| Final query-specific knowledge graph | |
| Node and edge sets of graph at iteration | |
| Node and edge sets of final graph | |
| Attribute vector of node | |
| Attribute vector of edge | |
| Answer Generation | |
| Answerer model | |
| Model-generated answer before retrieval | |

### v2

@@
 Abstract
-Retrieval-augmented language models often struggle with knowledge-intensive tasks due to inefficient retrieval, unstructured knowledge integration, and single-pass architectures.
-We present Retrieval-And-Structuring (RAS), a novel framework that dynamically constructs and reasons over query-specific knowledge graphs through iterative retrieval and structuring.
-RAS introduces four key technical innovations: (1) a theme-scoped retrieval mechanism that efficiently narrows the search space while maintaining retrieval quality, (2) an action planning module that determines knowledge needs and generates focused sub-queries, (3) a dynamic knowledge structuring approach that converts retrieved text into an evolving knowledge graph, and (4) a graph-augmented answering component that leverages the accumulated structured information.
-Our framework achieves state-of-the-art performance, surpassing leading baselines by 6.4% with open-source language models and 7.0% with proprietary models on seven knowledge-intensive generation datasets across all evaluation metrics.
-Detailed ablation studies verify the contribution of each technical component to the overall system performance.
+Large language models (LLMs) have achieved impressive performance on knowledge-intensive tasks, yet they often struggle with multi-step reasoning due to the unstructured nature of retrieved context.
+While retrieval-augmented generation (RAG) methods provide external information, the lack of explicit organization among retrieved passages limits their effectiveness, leading to brittle reasoning pathways.
+Recent interpretability studies highlighting the importance of structured intermediate reasoning further align with this perspective.
+We propose Retrieval-And-Structuring (RAS), a framework that dynamically constructs query-specific knowledge graphs through iterative retrieval and structured knowledge building.
+RAS interleaves targeted retrieval planning with incremental graph construction, enabling models to assemble and reason over evolving knowledge structures tailored to each query.
+On seven knowledge-intensive benchmarks, RAS consistently outperforms strong baselines, achieving up to 6.4% and 7.0% gains with open-source and proprietary LLMs, respectively.
+Our results demonstrate that dynamic, query-specific knowledge structuring offers a robust path to improving reasoning accuracy and robustness in language model generation.
+Our data and code can be found at https://github.com/pat-jj/RAS.
 1 Introduction
-Complex reasoning tasks like scientific analysis or multi-hop question answering demand both comprehensive knowledge and structured logical thinking.
-While large language models (LLMs) have achieved remarkable performance across natural language processing tasks (Devlin et al., 2018; Brown et al., 2020; Chowdhery et al., 2023), they often struggle with knowledge-intensive reasoning due to the absence of precise, logically structured information (Rae et al., 2021; Ling et al., 2024).
-This limitation has prompted growing research into augmenting LLMs with structured knowledge to enhance their reasoning capabilities (Wang et al., 2021).
-Retrieval-augmented generation (RAG) approaches have emerged as a promising direction by providing LLMs with additional context from retrieved passages (Guu et al., 2020; Lewis et al., 2020; Izacard & Grave, 2021; He et al., 2024).
-However, these methods frequently face hallucination challenges (Maynez et al., 2020; Zhang et al., 2023b), where generated content deviates from retrieved information.
-This issue stems primarily from the unstructured nature of retrieved passages, which lack explicit logical connections needed for precise reasoning.
-Recent efforts have explored integrating knowledge graphs (KGs) with LLMs (Sun et al., 2019; Yu et al., 2022; He et al., 2024; Edge et al., 2024), as KGs offer compact, structured representations of entities and relationships that enable more accurate and interpretable reasoning (Hogan et al., 2021; Jiang et al., 2024).
-We present Retrieval-And-Structuring (RAS), a novel framework that dynamically constructs and reasons over query-specific knowledge graphs through iterative retrieval and structuring.
-RAS integrates theme-aware retrieval, dynamic KG construction, and self-refinement (Ji et al., 2023; Asai et al., 2023; Lyu et al., 2024b) to comprehensively prepare knowledge for challenging questions.
-The framework comprises four key components: (1) an action planner that strategically determines knowledge needs and generates targeted sub-queries, (2) a theme-scoped retrieval module that efficiently combines thematic and semantic relevance, (3) a knowledge structuring module that builds and expands query-specific KGs, and (4) a knowledge-augmented answerer that leverages the gathered structured information.
-RAS addresses several critical limitations of existing approaches.
-First, unlike traditional RAG methods that rely solely on semantic similarity (Lewis et al., 2020; Guu et al., 2020), RAS employs theme-aware retrieval that significantly reduces the search space while preserving retrieval quality.
-Second, in contrast to static KG approaches (He et al., 2024; Edge et al., 2024), RAS dynamically constructs and evolves KGs based on the specific reasoning requirements of each query.
-Third, through its iterative planning mechanism, RAS actively identifies knowledge gaps and initiates focused retrievals, moving beyond the single-pass limitations of existing methods (Izacard & Grave, 2021).
-On seven knowledge-intensive tasks including open-domain QA, closed-set QA, and long-form generation, RAS outperforms state-of-the-art baselines by 6.4% with open-source LLMs and 7.0% with proprietary models.
-The main contributions of this work:
-We introduce RAS, a novel framework that dynamically constructs query-specific knowledge graphs through iterative retrieval and structuring for knowledge-intensive language generation.
-We propose theme-scoped retrieval that combines thematic classification with dense retrieval to efficiently narrow the search space while maintaining retrieval quality.
-We develop a dynamic knowledge structuring approach that iteratively builds and enriches knowledge graphs based on reasoning needs.
-We validate RAS’s effectiveness through comprehensive experiments and ablation studies across diverse knowledge-intensive tasks.
+Complex reasoning tasks such as scientific analysis or multi-hop question answering demand both comprehensive knowledge and structured logical thinking [1].
+While large language models (LLMs) have achieved remarkable performance across a wide range of natural language processing tasks [2, 3, 4], they often struggle with knowledge-intensive reasoning due to the absence of precise, logically organized information [5, 6].
+This limitation has motivated growing research into augmenting LLMs with structured knowledge to enhance their reasoning capabilities [7].
+Retrieval-augmented generation (RAG) approaches have emerged as a promising direction by providing LLMs with additional context from retrieved passages [8, 9, 10, 11].
+However, these methods frequently face hallucination challenges [12, 13], where generated content deviates from retrieved information.
+This issue stems primarily from the unstructured nature of retrieved passages, requiring the model to implicitly bridge gaps during generation.
+Recent research in model interpretability further highlights this issue [14].
+Attribution graph analyses have revealed that LLMs internally attempt multi-hop reasoning by chaining facts across different parts of context, but failures in these reasoning pathways are correlated with hallucinations and factual inconsistencies.
+These insights suggest that explicitly providing structured intermediate knowledge could significantly enhance the reliability and transparency of LLM reasoning.
+Recent efforts have explored integrating knowledge graphs (KGs) with LLMs [15, 16, 11, 17] to provide compact, relational representations that enable more interpretable reasoning [18, 19, 20].
+However, existing approaches largely rely on static, corpus-wide graphs, which may not adapt well to the specific reasoning needs of individual queries.
+Constructing and maintaining large-scale KGs also incurs substantial costs in open-domain settings.
+To address these challenges, we propose Retrieval-And-Structuring (RAS), a framework that dynamically constructs and reasons over query-specific knowledge graphs through iterative retrieval and structured knowledge building.
+RAS interleaves three key components: (1) a planning module that strategically identifies knowledge gaps and generates targeted sub-queries, (2) a retrieval-and-structuring module that extracts factual triples from retrieved passages and incrementally builds a query-specific graph, and (3) a knowledge-augmented answering module that generates outputs conditioned on the accumulated structured knowledge.
+RAS addresses several critical limitations of prior methods.
+Unlike traditional RAG approaches that rely on single-pass retrieval [8, 9, 10], RAS employs iterative planning to actively identify and fill knowledge gaps during inference.
+In contrast to static KG-based approaches [11, 17], RAS dynamically constructs query-specific knowledge graphs tailored to the evolving reasoning needs of each query.
+This targeted construction captures only task-relevant information, avoiding the inefficiencies and noise associated with static, corpus-level graphs, which enables more precise and robust language generation.
+Through extensive evaluations across seven benchmarks spanning open-domain QA, closed-set QA, and long-form generation, RAS consistently outperforms strong baselines by 6.4% with open-source LLMs and 7.0% with proprietary models.
+Our main contributions are:
+We introduce RAS, a novel framework for dynamic knowledge structuring through iterative retrieval and graph construction, tailored to each query.
+We develop a structure-aware planning and answering model that operates over evolving knowledge graphs.
+We demonstrate RAS’s effectiveness through comprehensive experiments and ablation studies across diverse knowledge-intensive tasks.
 2 Related Work
-Retrieval-Augmented Generation.
-Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020) enhances performance on knowledge-intensive tasks by incorporating retrieved passages into input of large language models (LLMs), improving factual accuracy and grounding.
-Traditional approaches retrieved a fixed number of passages once before generation (Shao et al., 2023; Es et al., 2024; Lyu et al., 2024a) or focused on adaptive retrieval (Jiang et al., 2023), where models dynamically query external knowledge when confidence is low, or analyze on the retrieval (Kim et al., 2024b).
-However, these methods often fail to perform complex or multi-hop reasoning process due to the linear and simple designs, where several recent methods have emerged to improve retrieval quality.
-Iterative retrieval-generation approaches (Shao et al., 2023; Guan et al., 2024) leverage historical context to enhance subsequent retrievals.
-Other work introduces targeted retrieval strategies through subqueries (Khattab et al., 2023) or follow-up questions (Yao et al., 2023; Press et al., 2022), progressively enriching the context for more factually grounded generation.
-SelfRAG (Asai et al., 2023) introduced self-reflective RAG, which allows LLMs to retrieve content on demand and critically evaluate retrievals, demonstrating both the importance of retrieval quality and the capacity of LLMs to seek self-directed information.
-RPG (Lyu et al., 2024b) further refined this approach by extracting fine-grained paragraphs from retrieved content to improve query relevance.
-However, existing methods face key limitations.
-They often either include irrelevant content in the retrieved context or risk discarding essential information through refinement.
-Additionally, dense retrieval approaches can be computationally expensive when deployed on large corpora, hampering real-time user interaction.
-To address these challenges, we propose converting retrieved text into knowledge-intensive graphs and maintaining an iteratively enriched, query-specific knowledge graph for knowledge serving.
-We also introduce theme-scoped retrieval for more focused and efficient context retrieval.
+Retrieval-Augmented Generation (RAG) [8, 9] enhances language model performance on knowledge-intensive tasks by incorporating retrieved passages into the model input, improving factual accuracy and grounding.
+Early RAG approaches retrieved a fixed number of passages once before generation [21, 22, 23], while later methods explored adaptive retrieval [24], where models dynamically issue retrieval queries based on uncertainty signals, or retrieval evaluation [25] to improve relevance.
+However, these techniques often struggle with complex or multi-hop reasoning due to their linear designs.
+Recent advances aim to strengthen retrieval quality by iterative refinement.
+Iterative retrieval-generation approaches [21, 26] leverage historical context to guide subsequent retrievals, while targeted retrieval strategies based on subqueries [27] or follow-up questions [28, 29] progressively enrich the retrieved evidence.
+Self-RAG [30] introduced self-reflective retrieval, allowing LLMs to selectively query for missing information and evaluate the quality of retrievals.
+RPG [31] further refined retrieval by extracting fine-grained paragraphs to enhance relevance.
+Despite these advances, existing methods face persistent challenges: retrieved context often contains redundant or irrelevant information, or conversely, discards essential facts during refinement.
+To address these issues, we propose converting retrieved content into a structured graph format and maintaining an iteratively enriched, query-specific knowledge graph, allowing the model to reason over an organized and progressively expanded knowledge structure.
 Graph as Context for LLMs.
-The integration of graphs into modern language models represents a promising research direction (Yasunaga et al., 2021, 2022; Yu, 2022; Ju et al., 2022; Zhu et al., 2024; Gutiérrez et al., 2024).
-Recent approaches leverage graph-based knowledge representation to enhance LLM performance.
-For instance, GraphToken (Perozzi et al., 2024) shows that LLM can understand a graph encoded as a token, after a simple p-tuning (Liu et al., 2021) while keeping the LLM’s parameters frozen.
-Based on the findings, G-Retriever (He et al., 2024) constructs subgraphs by retrieving relevant entities and relations from a global KG, using these as structured knowledge input for LLMs.
-GraphRAG (Edge et al., 2024) employs an alternative strategy, extracting a comprehensive global KG from the corpus, identifying meaningful clusters through graph community detection, and generating LLM-based summaries of these communities to create a refined retrieval corpus.
-These methods demonstrate how graphs, with their knowledge-intensive modality and inherent structural properties, can significantly improve LLMs’ contextual understanding.
-However, current approaches treat graphs as static data structures without refinement based on LLM requirements.
-Moreover, building global KGs for large-scale corpora remains prohibitively expensive and often impractical – Wikipedia alone contains over 30 million documents per dump.
-The key challenge lies in developing methods to dynamically construct and utilize graphs on-the-fly for knowledge-intensive generation.
-To address these limitations, we propose a retrieval-and-structuring paradigm that dynamically builds and maintains query-specific KGs on-the-fly.
-Our approach iteratively enriches these graphs through LLM’s self-evaluation of information completeness and targeted retrieval via focused subqueries.
-3 RAS Framework
-Knowledge-intensive language generation requires not just retrieving relevant information, but also structuring and reasoning over it effectively.
-We propose Retrieval-And-Structuring (RAS), a framework shown in Figure 1.
-In the following subsections, we detail the technical mechanisms underlying each stage.
-To facilitate understanding, we showcase a running example of RAS in Figure 26.
-3.1 Action Planning
-The Action Planning stage is the first critical step in the RAS framework, guiding the retrieval and reasoning process based on the input query and the current knowledge state.
-It involves two key components: the Action Planner , a language model that assesses the sufficiency of the current knowledge state and generates sub-queries for further retrieval if necessary, and a decision step that determines the next action based on the planner’s output.
-(For initial iteration) Formally, given an input query , the Action Planner generates a plan according to:
-where can take one of two forms:
-[NO_RETRIEVAL]: If determines that can be answered directly without requiring any additional knowledge, the planning process terminates, and the framework proceeds directly to the final Answer Generation stage (§3.4).
+Integrating structured graphs into language model workflows has emerged as a powerful approach for enhancing reasoning and factual accuracy [32, 33, 34, 35, 36, 37].
+Graphs provide explicit, relational representations that help models go beyond flat text by making entity connections and multi-hop relationships more tractable.
+Recent work, such as GraphToken [38], shows that LLMs can process serialized graph inputs through token-level prompts [39] without modifying model parameters.
+G-Retriever [11] leverages global knowledge graphs (KGs) to retrieve entity-centric subgraphs, while GraphRAG-style methods [17, 40] construct corpus-level graphs and generate summaries to enhance retrieval quality.
+Despite their benefits, these methods rely heavily on static, global graphs, which often include irrelevant or redundant information and cannot adapt to the needs of individual queries.
+Moreover, building and maintaining large-scale KGs is computationally expensive, particularly in open-domain settings.
+This motivates a shift toward dynamic, inference-time graph construction.
+Our approach embraces this direction by proposing a retrieval-and-structuring framework that builds query-specific knowledge graphs on the fly.
+Rather than relying on fixed knowledge coverage, we iteratively retrieve and structure information into task-relevant graphs that evolve with the reasoning process.
+This dynamic graph construction aligns with recent interpretability findings [14], which suggest that LLM failures often stem from breakdowns in implicit reasoning chains—failures that explicit graph structuring can help mitigate.
+3 Retrieval-And-Structuring (RAS) Framework
+Effective knowledge-intensive language generation requires not only retrieving relevant information, but also structuring and reasoning over it systematically.
+We introduce Retrieval-And-Structuring (RAS), a framework that interleaves iterative retrieval planning with dynamic knowledge graph construction, enabling large language models (LLMs) to reason over progressively organized knowledge tailored to each query.
+Figure 1 illustrates the overall workflow.
+Key Definitions.
+We define the core concepts used in the RAS framework as follows.
+The Main Query () denotes the original task input.
+A Subquery () is a focused retrieval query generated at iteration to obtain supporting evidence.
+Retrieved Passages () are the top- documents selected for each subquery.
+A text-to-triples model converts retrieved text into Extracted Triples (), structured as subject-predicate-object facts.
+These triples are incrementally accumulated into an evolving Knowledge Graph (), representing organized evidence about .
+The model produces an Action Plan () at each step, determining whether to continue retrieval ([SUBQ]), stop retrieval ([SUFFICIENT]), or, initially, answer directly without retrieval ([NO_RETRIEVAL]).
+3.1 Knowledge State-Aware Planning
+The Planning stage initiates and controls the retrieval-and-structuring process by dynamically assessing the current knowledge state.
+Initial Planning. Formally, given an input query , the model generates an initial plan :
+where is the planning instruction (as shown in Figure 15). can take one of two forms:
 [SUBQ] : If assesses that the query cannot be satisfactorily answered with its own knowledge, we start the iteration with the main query as the initial subquery, and move to the next stage (§3.2).
-(For iteration 0) At each iteration , the Action Planner reassesses the cumulative retrieved knowledge and the chain of previous sub-queries in the context of the main query :
-where is a graph neural network for encoding the evolving KG ; is the subquery at iteration , and is the extracted graph information (triple list) from the retrieved context.
-The base model we use here is GraphLLM, an architecture derived from previous work (Perozzi et al., 2024; He et al., 2024).
-We discuss more details in §3.3.
-The planner’s output at each iteration can be either:
+[NO_RETRIEVAL]: If determines that can be answered directly without requiring any additional knowledge, the planning process terminates, and the framework proceeds directly to the final Answering stage (§3.3).
+Iterative Planning. At iteration , given the accumulated knowledge and the subquery-triples history , the model updates the plan:
+where GNN is a graph neural network for encoding and projecting the evolving KG ; is the subquery at iteration , and is the extracted graph information (a list of triples) from the retrieved context .
+The output at each iteration can be either:
+[SUBQ] : The model generates a new subquery to guide the retrieval of additional relevant knowledge.
+The subquery is designed to fill specific gaps in the current knowledge state with respect to answering .
+The framework proceeds to the next stage (§3.2).
 [SUFFICIENT]: The accumulated knowledge is deemed sufficient to comprehensively address the main query .
-The iterative retrieval process terminates, and the framework proceeds to the Knowledge-Augmented Generation stage (§3.4).
-[SUBQ] : The planner generates a new subquery to guide the retrieval of additional relevant knowledge.
-The subquery is designed to fill specific gaps in the current knowledge state with respect to answering .
-The framework continues the iterative process and proceeds to the Theme-Scoped Retrieval stage (§3.2).
-The Action Planner serves as a key driver of the RAS framework’s iterative retrieval and refinement process.
+The iterative retrieval process terminates, and the framework proceeds to the Answering stage (§3.3).
+The Planning stage serves as a key driver of the RAS framework’s iterative retrieval and refinement process.
 By dynamically assessing the adequacy of the retrieved knowledge and generating targeted sub-queries, it enables the efficient acquisition of query-relevant information.
-3.2 Theme-Scoped Retrieval
-This stage efficiently retrieves thematically and semantically relevant context for subquery through a two-step process: (1) theme scoping to identify a narrowed sub-corpus close to ’s thematic distribution, and (2) dense retrieval to find the most relevant passages within .
-This thematic pre-filtering before semantic matching aims to improve both retrieval efficiency and precision.
-3.2.1 Theme Scoping
-The Theme Scoping mechanism transforms sub-queries into thematic probability distributions and retrieves relevant document subsets through three key operations:
-First, a theme classifier maps the subquery to a posteriori probability distribution (PPD) over thematic classes (Eq.
-3).
-The classifier is trained on DBPedia-298 (Lehmann et al., 2015)111Document label space of DBPedia-298: https://mappings.dbpedia.org/server/ontology/classes/, a multilabel wiki text classification dataset, enabling fine-grained thematic distinctions.
-To address the distribution shift between queries and documents, an MLP-based shifter transforms the query theme distribution into an expected document theme distribution (Eq.
-4).
-This shifter is trained on query-document pairs in our constructed dataset HotpotQA-SUBQ (see Appendix B), learning to capture the relationship between query and document thematic patterns.
-Finally, using the shifted distribution , we retrieve a theme-specific sub-corpus from the theme-indexed corpus (Eq.
-5).
-The function selects the top- documents based on L2 similarity between and the pre-computed theme distributions in .
-This focused retrieval process effectively reduces the search space while maintaining thematic relevance for subsequent semantic matching operations.
-3.2.2 Dense Retrieval
-Within the theme-specific sub-corpus , a dense retriever is employed to identify the most semantically relevant passages for the subquery :
-where selects the top- passages based on the cosine similarity between the query embedding and the passage embeddings precomputed in the dense index.
-3.3 Knowledge Structuring and Enrichment
-This stage is designed to transform the retrieved text passages into a structured graph representation and iteratively expand the query-specific knowledge graph .
-3.3.1 Text-to-Triple Conversion
-To capture the essential facts from the retrieved passages , a text-to-triples model is employed.
-We train the model on the full set of WikiOFGraph dataset (Kim et al., 2024a), an LLM-curated high-quality text-triples dataset, enabling it to generate quality and information-rich triples in the format:
+3.2 Text Retrieval and Knowledge Structuring
+This stage transforms retrieved text into structured knowledge and progressively builds the query-specific graph .
+Text Retrieval. We use a text retriever to retrieve the top- semantically relevant passages from the corpus for each subquery :
+We use a standard dense retriever by default but note that RAS is compatible with more advanced information retrieval methods [41, 42, 43].
+Text-to-Triples (Graph) Conversion.
+To extract essential factual information from the retrieved passages , we employ a text-to-triples model .
+This model is trained on the full WikiOfGraph dataset [44], which is a high-quality, LLM-curated text-to-triples corpus.
+Details of the training process are provided in Appendix C.1.
+The model generates structured triples in the following format:
 where each triple represents a subject-predicate-object fact extracted from the text.
-This structured representation allows for efficient downstream reasoning and simple external knowledge graph integration.
-3.3.2 Iterative Knowledge Enrichment
+This structured representation enables efficient downstream reasoning and facilitates integration with external knowledge graphs.
+Although is a lightweight LLM capable of fast inference using techniques such as quantization [45] and optimized inference frameworks like vLLM [46], the text-to-triples conversion can also be precomputed offline when maximal efficiency is required.
+Iterative Knowledge Enrichment to Query-Specific KG.
 The extracted triples are then converted into a graph structure , where and denote the sets of nodes and edges, respectively.
 Each node corresponds to a unique subject or object entity in , while each edge represents a predicate connecting two entities.
-To enrich the graph with semantic information, the attributes of nodes and edges are obtained through a text encoder (Sentence-BERT (Reimers, 2019) in our case):
+To enrich the graph with semantic information, the attributes of nodes and edges are obtained through a text encoder (Sentence-BERT [47]):
 These semantic embeddings enable the model to capture the nuanced relationships between entities and facilitate reasoning over the knowledge graph.
-To progressively expand the knowledge base in response to the evolving sub-queries, the structured graph at each iteration is merged into a evolving knowledge graph specific to the main query :
-After enriching with the new knowledge, the action planner (§3.1) reassesses the current knowledge to determine the next step.
-Based on and the chain of previous subqueries and their associated graph information, the planner decides whether to generate another focused subquery for additional retrieval (§3.2) or to proceed with answering (§3.4) if the accumulated knowledge is deemed sufficient.
-3.4 Knowledge-Augmented Generation
-The final stage generates an answer to the main query using the Answerer model conditioned on knowledge graph and subquery chain .
+To progressively enrich the query-related knowledge in response to the evolving sub-queries, the structured graph at each iteration is merged into an evolving KG specific to the main query :
+After enriching with the new knowledge, the Planning module (§3.1) reassesses the current knowledge to determine the next step.
+Based on and the chain of previous subqueries and their associated graph information, the model decides whether to generate another focused subquery for additional retrieval or to proceed with answering (§3.3) if the accumulated knowledge is sufficient.
+3.3 Knowledge-Augmented Answering
+At the final stage, the model generates an answer to the main query either conditioned on knowledge graph and subquery chain when retrieval-and-structuring was processed, or directly when no retrieval is needed.
 If no retrieval is needed (), the answer is generated directly:
 Otherwise, after iterative knowledge enrichment concludes with [SUFFICIENT] plan, the answer is generated using encoded KG and subquery chain:
-attends to relevant knowledge in and subquery chain to generate accurate, coherent answers.
-Note: In this study, and are two roles played by the same model with different instructions (see §C.3).
+where is the answering instruction (as shown in Figure 16).
+attends to knowledge in and subquery chain to generate accurate, coherent answers.
+This structured conditioning enables systematic reasoning grounded in the assembled knowledge.
+3.4 Sturcture-Aware Multitask Learning
+The RAS framework is trained through a multitask setup that unifies knowledge planning and knowledge-augmented answering under a standard next-token prediction objective.
+Each training instance corresponds to either a planning task or an answering task, selected randomly:
+Planning Task: Input: (encoded ), planning instruction , subquery-triples history, main query . Output: next plan .
+Answering Task: Input: (encoded ), answering instruction , full subquery-triples history, main query . Output: final answer .
+The model used in RAS is Graph LLM, an architecture adapted from prior work [38, 11].
+It is trained to perform both planning and answering tasks via standard next-token prediction, conditioned on the instruction type and input structure.
+To the best of our knowledge, this is the first application of Graph LLMs to general language generation tasks such as open question answering.
 4 Experiments
 Training Data & Setting.
-We develop HotpotQA-SUBQ, a dataset derived from HotpotQA (Yang et al., 2018), to train our model’s action planning and answering capabilities.
-Our dataset creation begins with document filtering: using Claude 3.5 Sonnet (Anthropic, 2024), we identify and retain only the supporting documents necessary for answering the main query, removing irrelevant content.
+We develop HotpotQA-SUBQ, a dataset constructed based on HotpotQA [1], to train our model’s action planning and answering capabilities.
+Our dataset creation begins with document filtering: using Claude-3.5-Sonnet [48], we identify and retain only the supporting documents necessary for answering the main query, removing irrelevant content.
 For each supporting document , we then iteratively generate a subquery , considering the main query, previous subqueries, and supporting documents.
@@
 For correctly answered queries, we create training samples with the main query as input and “[NO_RETRIEVAL]” as the output label.
-Additionally, to ensure fair comparison with existing approaches, we incorporate the subset of Arc-Easy (2,147 samples) and ASQA (3,897 samples) from SelfRAG’s training data, resulting in 208k training samples in total.
+Additionally, to ensure fair comparison with existing approaches, we incorporate the subset of Arc-Easy (2,147 samples) and ASQA (3,897 samples) from Self-RAG’s training data, resulting in 208k training samples in total.
 We place detailed training data processing, dataset statistics, and data samples in Appendix B.
-We use DBPedia-298, HotpotQA-SUBQ, and WikiOFGraph to train our theme classifier , distribution shifter , and text-to-triples model , respectively, as detailed in Appendix C.
-We present our hyperparameter study of each component in Appendix G.
+For efficient inference, we train a text-to-triples model on the WikiOFGraph dataset [44] using LLaMA-3.2-3B-Instruct as the base model (see Appendix C), and deploy it using vLLM for optimized runtime performance.
+We present our hyperparameter study of each component in Appendix H.
 Knowledge Sources.
-Due to the large sizes of wiki corpora, we employ faiss (Douze et al., 2024) for efficient vector searching across both theme and dense indices.
-Following the SelfRAG (Asai et al., 2023), we utilize the Wikipedia 2018 (Izacard et al., 2023) by default, while specifically using the Wikipedia 2020 for PopQA to access more recent information.
+Due to the large sizes of wiki corpora, we employ faiss [49] for efficient vector searching over the dense index.
+Following the Self-RAG [30], we utilize the Wikipedia 2018 [50] by default, while specifically using the Wikipedia 2020 for PopQA to access more recent information.
 To optimize retrieval efficiency, we partition the index into five segments.
-Within each segment, we first perform theme scoping to reduce the candidate document pool to 250,000 documents by default.
-However, we skip theme scoping for PopQA since our classifier , trained on DBPedia-298 data based on old Wiki dump, performs poorly on newer Wiki 2020 dump, producing near-zero vectors for roughly one-third of recent content.
 Test Datasets & Metrics & Compared Baselines.
-We conduct comprehensive evaluations on diverse knowledge-intensive tasks following previous studies (Asai et al., 2023; Lyu et al., 2024b).
-The evaluation encompasses three categories of datasets: (1) open-domain short-form generation datasets: TriviaQA (Joshi et al., 2017), PopQA (Mallen et al., 2022), and 2WikiMultihopQA (Ho et al., 2020); (2) closed-set task datasets: PubHealth (Zhang et al., 2023a) and ARC-Challenge (Clark et al., 2018); and (3) long-form generation datasets: ALCE-ASQA (Gao et al., 2023; Stelmakh et al., 2022) and ELI5 (Fan et al., 2019).
-For evaluation metrics, we maintain consistency with prior work (Asai et al., 2023; Mallen et al., 2022; Lyu et al., 2024b), employing “golden match” accuracy for PopQA and TriviaQA, token-level F1 score for 2WikiMultihopQA, accuracy for PubHealth and ARC-Challenge, and ROUGE-LSum alongside MAUVE score (Pillutla et al., 2021) for ASQA and ELI5.
-Our comparative analysis includes three baseline categories: models without retrieval augmentation, incorporating Claude 3.5 Sonnet as a state-of-the-art closed-source baseline; models with single retrieval over top-5 documents, including Claude 3.5 Sonnet, and SuRe (Kim et al., 2024b), a leading retrieve-and-summarize framework; and models with self-reflective retrieval, including leading approaches Self-RAG (Asai et al., 2023) and RPG (Lyu et al., 2024b).
+We conduct comprehensive evaluations on diverse knowledge-intensive tasks following previous studies [30, 31].
+The evaluation encompasses three categories of datasets: (1) open-domain short-form generation datasets: TriviaQA [51], PopQA [52], and 2WikiMultihopQA [53]; (2) closed-set task datasets: PubHealth [54] and ARC-Challenge [55]; and (3) long-form generation datasets: ALCE-ASQA [56, 57] and ELI5 [58].
+For evaluation metrics, we maintain consistency with prior work [30, 52, 31], employing “golden match” accuracy for PopQA and TriviaQA, token-level F1 score for 2WikiMultihopQA, accuracy for PubHealth and ARC-Challenge, and ROUGE-LSum alongside MAUVE score [59] for ASQA and ELI5.
+Our comparative analysis includes three baseline categories: models without retrieval augmentation, incorporating Claude 3.5 Sonnet as a state-of-the-art closed-source baseline; models with single retrieval over top-5 documents, including Claude 3.5 Sonnet, and SuRe [25], a leading retrieve-and-summarize method; and models with self-reflective retrieval, including leading approaches Self-RAG [30], RPG [31], and Search-R1 [60].
 We place more details of metrics/datasets in Appendix D.
-| Short-form | Closed-set | Long-form Generation | ||||||||
-| Model | TQA | 2WQA | PopQA | Pub | ARC | ASQA | ELI5 | |||
-| (acc) | (F1) | (acc) | (acc) | (acc) | (rouge) | (mauve) | (rouge) | (mauve) | ||
-| Baselines without retrieval | ||||||||||
-| Llama2 | 30.5 | 18.9 | 14.7 | 34.2 | 21.8 | 15.3 | 19.0 | 18.3 | 32.4 | |
-| Alpaca | 54.5 | – | 23.6 | 49.8 | 45.0 | 29.4 | 61.7 | – | – | |
-| Llama2 | 38.5 | 20.2 | 14.7 | 29.4 | 29.4 | 12.4 | 16.0 | 18.2 | 41.4 | |
-| Alpaca | 61.3 | – | 24.4 | 55.5 | 54.9 | 32.0 | 70.6 | – | – | |
-| ChatGPT | 74.3 | 24.8 | 29.3 | 70.1 | 75.3 | 36.2 | 68.8 | 22.8 | 32.6 | |
-| Sonnet-3.5 | 78.4 | 40.0 | 30.2 | 83.7 | 88.5 | 37.0 | 39.1 | 21.8 | 26.5 | |
-| Baselines with retrieval (#docs=5 by default) | ||||||||||
-| Llama2 | 42.5 | 21.0 | 38.2 | 30.0 | 48.0 | 22.1 | 32.0 | 18.6 | 35.3 | |
-| Alpaca | 64.1 | – | 46.7 | 40.2 | 48.0 | 33.3 | 57.9 | – | – | |
-| Llama2 | 47.0 | 31.2 | 45.7 | 30.2 | 26.0 | 20.5 | 24.7 | 18.6 | 42.3 | |
-| Alpaca | 66.9 | – | 46.1 | 51.1 | 57.6 | 36.7 | 56.6 | – | – | |
-| RECOMP (Xu et al., 2023) | – | 32.4 | – | – | – | 36.5 | 76.0 | – | – | |
-| SuRe (Kim et al., 2024b) | – | 20.6 | – | – | – | 35.8 | 76.2 | – | – | |
-| SuRe | 72.3 | 38.1 | 53.6 | 57.2 | 79.6 | 36.0 | 74.2 | – | – | |
-| ChatGPT | 65.7 | 32.8 | 50.8 | 54.7 | 75.3 | 39.9 | 79.7 | 20.6 | 57.2 | |
-| Sonnet-3.5 | 69.1 | 41.9 | 51.5 | 49.1 | 88.6 | – | – | – | – | |
-| Sonnet-3.5 | 72.5 | 53.7 | 57.3 | 53.9 | 87.1 | 38.8 | 61.6 | 20.2 | 32.3 | |
-| SuRe | 76.8 | 37.6 | 41.2 | 62.8 | 91.6 | 30.2 | 69.9 | 15.4 | 27.2 | |
-| Baselines with self-reflective retrieval | ||||||||||
-| SELF-RAG (Asai et al., 2023) | 66.4 | 25.1 | 54.9 | 72.4 | 67.3 | 35.7 | 74.3 | 17.9 | 35.6 | |
-| SELF-RAG | 69.3 | – | 55.8 | 74.5 | 73.1 | 37.0 | 71.6 | – | – | |
-| RPG (Lyu et al., 2024b) | – | 33.6 | 56.0 | 73.4 | – | 37.6 | 84.4 | 19.1 | 46.4 | |
-| Retrieval-And-Structuring (Ours) | ||||||||||
-| RAS | RAS | 77.4 | 57.7 | 62.3 | 71.4 | 93.8 | 39.1 | 70.5 | 23.3 | 37.7 |
-| RAS | RAS | 73.8 | 43.4 | 61.3 | 75.5 | 62.4 | 37.2 | 95.2 | 19.7 | 47.8 |
-| RAS | RAS | 78.7 | 54.7 | 58.8 | 70.8 | 91.2 | 39.1 | 70.5 | 23.3 | 37.7 |
-| RAS | RAS | 72.6 | 42.1 | 58.3 | 74.8 | 67.7 | 37.2 | 95.2 | 19.7 | 47.8 |
-[…]
-| RAS | RAS | 73.8 | 44.2 | 57.7 | 77.5 | 71.5 | 37.6 | 96.2 | 20.1 | 54.4 |
+| Short-form | Closed-set | Long-form Generation | |||||||
+| Model | TQA | 2WQA | PopQA | Pub | ARC | ASQA | ELI5 | ||
+| (acc) | (F1) | (acc) | (acc) | (acc) | (rouge) | (mauve) | (rouge) | (mauve) | |
+| Baselines w/o Retrieval | |||||||||
+| Llama2 | 30.5 | 18.9 | 14.7 | 34.2 | 21.8 | 15.3 | 19.0 | 18.3 | 32.4 |
+| Alpaca | 54.5 | – | 23.6 | 49.8 | 45.0 | 29.4 | 61.7 | – | – |
+| Llama2 | 38.5 | 20.2 | 14.7 | 29.4 | 29.4 | 12.4 | 16.0 | 18.2 | 41.4 |
+| Alpaca | 61.3 | – | 24.4 | 55.5 | 54.9 | 32.0 | 70.6 | – | – |
+| Llama3 | 56.1 | 21.2 | 26.7 | 33.2 | 42.2 | 17.6 | 25.0 | 18.2 | 39.7 |
+| ChatGPT | 74.3 | 24.8 | 29.3 | 70.1 | 75.3 | 36.2 | 68.8 | 22.8 | 32.6 |
+| Sonnet-3.5 | 78.4 | 40.0 | 30.2 | 83.7 | 88.5 | 37.0 | 39.1 | 21.8 | 26.5 |
+| Baselines w/ Retrieval (#docs=5 by default) | |||||||||
+| Llama2 | 42.5 | 21.0 | 38.2 | 30.0 | 48.0 | 22.1 | 32.0 | 18.6 | 35.3 |
+| Alpaca | 64.1 | – | 46.7 | 40.2 | 48.0 | 33.3 | 57.9 | – | – |
+| Llama2 | 47.0 | 31.2 | 45.7 | 30.2 | 26.0 | 20.5 | 24.7 | 18.6 | 42.3 |
+| Llama3 | 60.4 | 33.4 | 48.6 | 36.5 | 40.1 | 23.9 | 52.1 | 18.8 | 40.7 |
+| Alpaca | 66.9 | – | 46.1 | 51.1 | 57.6 | 36.7 | 56.6 | – | – |
+| RECOMP [61] | – | 32.4 | – | – | – | 36.5 | 76.0 | – | – |
+| ChatGPT | 65.7 | 32.8 | 50.8 | 54.7 | 75.3 | 39.9 | 79.7 | 20.6 | 57.2 |
+| Sonnet-3.5 | 69.1 | 41.9 | 51.5 | 49.1 | 88.6 | – | – | – | – |
+| Sonnet-3.5 | 72.5 | 53.7 | 57.3 | 53.9 | 87.1 | 38.8 | 61.6 | 20.2 | 32.3 |
+| SuRe [25] | 51.2 | 20.6 | 39.0 | 36.2 | 52.7 | 35.8 | 76.2 | 16.1 | 26.6 |
+| SuRe | 72.3 | 38.1 | 53.6 | 57.2 | 79.6 | 36.0 | 74.2 | 19.2 | 51.6 |
+| SuRe | 76.8 | 37.6 | 41.2 | 62.8 | 91.6 | 30.2 | 69.9 | 15.4 | 27.2 |
+| Baselines w/ Self-Reflective Retrieval | |||||||||
+| Self-RAG [30] | 66.4 | 25.1 | 54.9 | 72.4 | 67.3 | 35.7 | 74.3 | 17.9 | 35.6 |
+| Self-RAG | 69.3 | – | 55.8 | 74.5 | 73.1 | 37.0 | 71.6 | – | – |
+| RPG [31] | 65.1 | 33.6 | 56.0 | 73.4 | 65.4 | 37.6 | 84.4 | 19.1 | 46.4 |
+| Search-R1 [60] | 63.8 | 35.2 | 45.7 | 69.4 | 64.4 | 35.2 | 72.3 | 18.4 | 37.1 |
+| Retrieval-And-Structuring (RAS) | |||||||||
+| RAS | 77.6 | 57.7 | 62.3 | 71.3 | 93.9 | 39.1 | 70.5 | 23.3 | 37.7 |
+| RAS | 73.8 | 44.2 | 57.7 | 77.6 | 71.4 | 37.6 | 96.2 | 20.1 | 54.4 |
+| RAS | 72.7 | 42.1 | 58.3 | 74.7 | 68.5 | 37.2 | 95.2 | 19.7 | 47.8 |
 Inference Setting.
-We evaluate RAS using both our trained open-source model (RAS with LLaMA-2-7B/LLaMA-3-8B and Graph Transformer (Shi et al., 2020) encoder) and the closed-source Claude-3.5-Sonnet model under varied inference settings.
-For RAS, we maintain consistency with previous work (Asai et al., 2023; Lyu et al., 2024b) by employing zero-shot inference across all datasets.
-For closed-source models (e.g., RAS), we utilize few-shot inference with two exemplars for ASQA and ELI5, while maintaining zero-shot for the other datasets.
-During inference, if the initial planner output is not [NO_RETRIEVAL], we utilize the main query as the initial subquery .
-For PopQA and TriviaQA evaluation, we follow established settings (Asai et al., 2023; Luo et al., 2023), incorporating top-five web search engine results as initial retrieved context .
-For ASQA and ELI5, we maintain methodological consistency with prior work (Lyu et al., 2024b; Asai et al., 2023; Gao et al., 2023), utilizing their predetermined five-document context.
-Under these conditions, we omit plan generation and text retrieval phases, implementing static inference (Asai et al., 2023) with five fixed iterations.
+We evaluate RAS using both our trained open-source models (RAS, based on LLaMA-2-7B/LLaMA-3-8B and a Graph Transformer encoder [65]) and the closed-source Claude-3.5-Sonnet model under varied inference settings.
+For open-source models (RAS), we follow prior work [30, 31] and adopt zero-shot inference across all datasets.
+For the closed-source model (RAS), we apply few-shot inference with two exemplars for ASQA and ELI5, while using zero-shot inference for all other datasets.
+More details are provided in Appendix E.1 and E.2.
+During inference, if the initial planning output is not [NO_RETRIEVAL], we utilize the main query as the initial subquery .
+For PopQA and TriviaQA evaluation, we follow established settings [30, 66], incorporating top-five web search engine results as initial retrieved context .
+For ASQA and ELI5, we maintain methodological consistency with prior work [31, 30, 56], utilizing their predetermined five-document context.
+Under these conditions, we omit plan generation and text retrieval phases, implementing static inference [30] with five fixed iterations.
 For remaining datasets, we establish a maximum iteration count of five.
-Following previous studies, we employ Contriever-MS MARCO (Izacard et al., 2021) as the primary dense retriever, with BM25 (Robertson et al., 2009) serving as the retrieval mechanism for 2WikiMultihopQA.
+Following previous studies, we employ Contriever-MS MARCO [67] as the primary dense retriever, with BM25 [68] serving as the retrieval mechanism for 2WikiMultihopQA.
 Across all retrieval processes, we maintain a consistent top- document selection of five.
 We present more details of inference settings in Appendix E.
-| TQA | 2WQA | Pub | ASQA | ||
-| (acc) | (F1) | (acc) | (rg) | (mv) | |
-| RAS | 72.6 | 42.1 | 74.8 | 37.2 | 95.2 |
-| Training Phase | |||||
-| No GraphToken | 70.2 | 38.4 | 66.4 | 33.1 | 85.0 |
-| No LoRA | 71.5 | 37.8 | 54.8 | 32.8 | 84.8 |
-| No Text-to-Triple | 70.4 | 38.2 | 71.4 | 36.2 | 73.8 |
-| No Mutli-Task | 68.6 | 39.2 | 65.5 | 36.7 | 88.9 |
-| Inference Phase | |||||
-| No Retrieval | 56.9 | 27.4 | 69.0 | 31.3 | 70.6 |
-| No GraphToken | 68.8 | 38.7 | 67.3 | 36.5 | 93.6 |
-| No Planning | 66.7 | 37.8 | 71.5 | – | – |
-| No ThemeScope | 72.0 | – | 73.9 | – | – |
 4.1 Main Results
-Our performance evaluation, presented in Table 1, demonstrates that the LLaMA-2-7B/LLaMA-3-8B model fine-tuned with RAS outperforms existing open-source solutions, including SelfRAG (Asai et al., 2023) and RPG (Lyu et al., 2024b).
+Our performance evaluation, presented in Table 1, demonstrates that the LLaMA-2-7B/LLaMA-3-8B model fine-tuned with RAS outperforms existing open-source solutions, including Self-RAG [30] and RPG [31].
 Notably, compared to the previous SOTA models, RAS shows a 9.7% improvement in short-form question-answering and a 7.9% gain in long-form generation tasks.
-Additionally, when applied to closed-source Sonnet-3.5 model, RAS consistently achieves superior results compared to single retrieval RAG approaches, including retrieve-and-summarize approach SuRe (Kim et al., 2024b).
+Additionally, when applied to closed-source Sonnet-3.5 model, RAS consistently achieves superior results compared to single retrieval RAG approaches, including retrieve-and-summarize approach SuRe [25].
 We find that sometimes (e.g., on TriviaQA and PubHealth) single-hop retrieval could not boost LLM’s performance, and even makes it worse, which demonstrates the necessity of on-demand retrieval, aligning with previous findings.
-Our “role-swapping” study, where Sonnet-3.5 and RAS alternately serve as action planner and answerer, reveals that performance is primarily bounded by the answerer’s inherent capabilities.
-For example, when Sonnet-3.5 acts as the planner with RAS as the answerer, the system achieves 62.4% accuracy on the ARC-Challenge dataset.
-However, when Sonnet-3.5 serves as both planner and answerer, accuracy increases to 93.8%.
-Notably, despite having 60x fewer parameters, our trained RAS performs comparably to Sonnet-3.5 as a planner, and even outperforms it sometimes.
-Also, for long-form generation task, our trained model achieved notably high MAUVE scores, which can be majorly attributed to our “text-to-triple” generation step, as demonstrated by our later ablation study in Table 2.
+Our “role-swapping” study in Figure 3, where Sonnet-3.5 and RAS alternately perform the planning and answering tasks, reveals that performance is primarily bounded by the answering capability.
+For example, when Sonnet-3.5 handles planning and RAS handles answering, the system achieves 62.4% accuracy on the ARC-Challenge dataset.
+However, when Sonnet-3.5 is responsible for both planning and answering, accuracy increases to 93.9%.
+Notably, despite having 60 fewer parameters, our trained RAS achieves planning performance comparable to Sonnet-3.5, and even outperforms it in some cases.
+Also, for long-form generation, our trained model achieved notably high MAUVE scores, which can be majorly attributed to our “text-to-triple” step, as demonstrated by our ablation study in Table 2.
 4.2 Ablations & More Analysis
-We study the effect of individual component in Table 2.
+We analyze the impact of each component of RAS in Table 2 and provide detailed discussion below.
 Effect of Iterative Planning and Retrieval.
@@
 Also, when turning off retrieval, the performance degradation is more obvious, due to the knowledge-intensive nature of those datasets.
+| TQA | 2WQA | Pub | ASQA | ||
+| (acc) | (F1) | (acc) | (rg) | (mv) | |
+| RAS | 72.7 | 42.1 | 74.7 | 37.2 | 95.2 |
+| Training Phase | |||||
+| No GraphEncode | 70.2 | 38.4 | 66.4 | 33.1 | 85.0 |
+| No LoRA | 71.5 | 37.8 | 54.8 | 32.8 | 84.8 |
+| No Text-to-Triple | 70.4 | 38.2 | 71.4 | 36.2 | 73.8 |
+| No Multi-Task | 68.6 | 39.2 | 65.5 | 36.7 | 88.9 |
+| Inference Phase | |||||
+| No Retrieval | 56.9 | 27.4 | 69.0 | 31.3 | 70.6 |
+| No GraphEncode | 68.8 | 38.7 | 67.3 | 36.5 | 93.6 |
+| No Planning | 66.7 | 37.8 | 71.5 | 37.2 | 95.2 |
 Effect of Graph Construction and Encoding.
@@
 Using only graph token optimization without LoRA leads to substantial degradation (-11.8% on average).
-A similar observation can be made for “No Multi-Task” (e.g., 12.3% accuracy degradation on PubHealth), indicating the significance of jointly training the model on both action planning and answer generation tasks rather than optimizing for each task separately, supporting findings from prior work (Lyu et al., 2024b).
+A similar observation can be made for “No Multi-Task” (e.g., 12.3% accuracy degradation on PubHealth), indicating the significance of jointly training the model on both action planning and answer generation tasks rather than optimizing for each task separately, supporting findings from prior work [31].
 The complementary effects suggest that while graph-based knowledge representation is valuable, it needs to be combined with careful parameter tuning and multi-task learning to achieve optimal performance.
-Effect of Theme Scoping.
-The theme-scoped retrieval shows mixed results – while removing it (“No ThemeScoping”) has minimal performance drops on TQA and PubHealth, the efficiency benefits are substantial as shown in Figure 2.
-This suggests theme scoping successfully narrows the search space without compromising performance.
-Efficiency Analysis.
-Figure 2 compares the computational costs of different retrieval strategies.
-Theme-scoped RAG reduces processing time by approximately 60% compared to dense-only RAG by first filtering the corpus through lightweight theme classification (with 298 class dimensions).
-This combination of comparable performance with significantly reduced computational cost demonstrates the effectiveness of our theme-scoped retrieval approach.
+Impact of Graph Information Abundance.
+Figure 3 shows that increasing the amount of structured graph information consistently improves RAS’s performance across tasks.
+For TQA, both RAS and RAS exhibit approximately linear gains as more triples are retained, with no clear saturation even at 100% information.
+For ASQA-MV, RAS benefits from a sharp improvement between 10% and 30% graph information, followed by steady increases, while RAS maintains a smoother and more stable growth pattern throughout.
+These results confirm that RAS effectively leverages structured knowledge at all levels of availability, and that more complete knowledge graphs consistently translate to stronger reasoning and generation quality.
+Additionally, the larger 8B model consistently outperforms the 7B variant under all conditions, suggesting that scaling the model size further enhances RAS’s ability to utilize structured knowledge.
+Interestingly, even partial graphs containing only 30%–50% of triples already deliver substantial gains over low-information baselines, highlighting RAS’s robustness to incomplete or partially retrieved knowledge.
 Impact of Training Data Volume.
@@
 However, we observed an exception with the ELI5 dataset, where performance were inconsistent.
-This irregularity can be attributed to the inclusion of ASQA training data in our training set, following established setting from previous research (Asai et al., 2023; Lyu et al., 2024b).
+This irregularity can be attributed to the inclusion of ASQA training data in our training set, following established setting from previous research [30, 31].
 Among our test datasets, ASQA and ELI5 are unique in requiring long-form response generation.
 The periodic decline in ELI5 performance suggests that the model’s response generation began to align more closely with ASQA’s training data distribution, potentially at the expense of ELI5’s distinct characteristics.
-Impact of Graph Information Abundance.
-Figure 4 demonstrates a strong positive correlation between graph information abundance and model performance across all tasks.
-Both RAS and RAS show consistent improvements as graph information increases, demonstrating RAS’s strong capability of effective utilizing available graph information for knowledge-intensive generation.
+| Triple Extractor | TQA | 2WQA | PopQA |
+| Flan-T5-Large | 70.3 | 40.7 | 56.7 |
+| LLaMA-3.2-3B | 72.7 | 42.1 | 58.3 |
+| Claude-3.5-Sonnet | 73.8 | 44.5 | 60.1 |
+Impact of Triple Extraction Quality We further examine how the quality of triple extraction influences the overall effectiveness of RAS.
+Specifically, we test three text-to-triples models: Flan-T5-Large, LLaMA-3.2-3B model and Claude-3.5-Sonnet where the first two are trained on WikiOfGraph data.
+As shown in Table 3, higher-quality triple extraction leads to consistent performance gains across all evaluated datasets.
+Notably, using Claude-3.5-Sonnet triples improves TriviaQA accuracy from 72.7% to 73.8%, boosts 2WikiMultihopQA F1 from 42.1 to 44.5, and raises PopQA accuracy from 58.3% to 60.1%.
+These results underscore that while RAS achieves strong performance even with lightweight extraction models, its effectiveness can be further enhanced by supplying more precise and complete structured knowledge during inference.
 5 Conclusion
-We presented RAS (Retrieval-And-Structuring), a novel framework that enhances knowledge-intensive language generation through iterative retrieval and knowledge structuring.
-Our approach introduces key innovations including theme-scoped retrieval, action planning, dynamic knowledge structuring, and graph-augmented answering.
-Comprehensive experiments demonstrate RAS’s superior performance over existing approaches across diverse benchmarks.
-Looking ahead, RAS’s graph-based architecture enables seamless integration with external knowledge graphs and human-in-the-loop curation, making it particularly valuable for specialized domains requiring high accuracy and reliability.
-We discuss limitations of RAS and future work in Appendix A.
-Our code can be found at https://github.com/pat-jj/Retrieval-And-Structure.
-Impact Statements
-The RAS framework has several potential positive impacts on both research and practical applications.
-First, by improving knowledge-intensive language generation through structured knowledge representation and iterative reasoning, RAS could enhance the reliability and factual accuracy of AI systems in critical domains like education, research assistance, and technical documentation.
-The theme-scoped retrieval mechanism also offers computational efficiency benefits, potentially reducing the energy consumption and computational resources required for large-scale knowledge retrieval tasks.
-However, there are important considerations regarding potential risks and challenges.
-The framework’s effectiveness in knowledge structuring could potentially amplify existing biases in training data and knowledge sources.
-Additionally, while RAS improves retrieval efficiency, it may create barriers for low-resource languages or domains where structured knowledge bases and thematic classifications are limited.
-These concerns necessitate careful consideration of data sources and regular evaluation of the system’s impact across different user groups and application contexts.
+We presented RAS (Retrieval-And-Structuring), a novel framework that improves knowledge-intensive language generation by interleaving iterative retrieval with dynamic knowledge graph construction.
+RAS introduces key innovations, including structure-aware action planning and graph-augmented answering, enabling models to reason over progressively organized information tailored to each query.
+Extensive experiments across seven benchmarks demonstrate that RAS consistently outperforms existing methods, particularly in tasks requiring multi-step reasoning and factual consistency.
+Furthermore, the modular, graph-based architecture of RAS allows seamless integration with external knowledge sources and human feedback, making it well-suited for high-stakes domains where transparency and adaptability are essential.
+Limitations and directions for future work are discussed in Appendix A.
 Appendix A Limitations & Future Work
@@
 Knowledge Extraction Quality: The effectiveness of both planning and answer generation depends on the quality of text-to-triple extraction.
-Despite using a Flan-T5-based model for triple extraction, implementing a more sophisticated approach – such as a dedicated Named Entity Recognition and Relation Extraction (NER-RE) pipeline – could potentially enhance RAS’s reasoning capabilities and knowledge representation.
-Theme Scoping Limitations: The current classifier was trained on DBPedia-298, an older dataset that may not adequately represent the thematic distribution of contemporary document collections.
-This temporal gap particularly affects performance when processing newer corpora, as evidenced by our experiments with Wikipedia 2020 dumps.
-The development of a huge, up-to-date multilabel document classification dataset would likely improve the effectiveness of theme scoping and, consequently, the overall efficiency gains it provides.
+Despite using an end-to-end text-to-triples model for triple extraction, implementing a more sophisticated approach – such as a dedicated Named Entity Recognition and Relation Extraction (NER-RE) pipeline – could potentially enhance RAS’s reasoning capabilities and knowledge representation.
 Graph Evolution Strategy: The current approach to evolving knowledge graphs could be enhanced by incorporating more sophisticated graph pruning and merging strategies.
@@
 Future versions could incorporate human feedback loops to improve the quality of knowledge structuring and reasoning paths.
-These limitations suggest several promising directions for future research, including improved knowledge extraction techniques, more robust theme classification approaches, and enhanced graph evolution strategies.
+These limitations suggest several promising directions for future research, including improved knowledge extraction techniques and enhanced graph evolution strategies.
 A particularly promising direction is the integration of external knowledge graphs into RAS’s iterative loop, as its graph-based architecture naturally accommodates the incorporation of established KGs like Wikidata or domain-specific knowledge bases.
@@
 Appendix B HotpotQA-SUBQ Dataset
-The source of our dataset is HotpotQA (Yang et al., 2018).222We use the version hotpot_train_v1.1 from https://github.com/hotpotqa/hotpot.
+The source of our dataset is HotpotQA [1].111We use the version hotpot_train_v1.1 from https://github.com/hotpotqa/hotpot.
 Our goal is to create a dataset that can let the model learn the capabilities of subquery generation and answering with intensive knowledge.
@@
 B.3 Sample Labeling
-The sample labeling process transforms the filtered and subquery-augmented HotpotQA examples into training instances for both the Action Planner and Answerer components.
+The sample labeling process transforms the filtered and subquery-augmented HotpotQA examples into training instances for both the Planning and Answering components.
 We formalize this process in Algorithm 1.
@@
 For each example, we first attempt direct answer generation using the base LLM (Line 3).
-If successful, we create [NO_RETRIEVAL] training instances for both the planner and answerer (Lines 4-6).
+If successful, we create [NO_RETRIEVAL] training instances for both the planning and answering (Lines 4-6).
 For examples requiring retrieval, we process each supporting document sequentially:
@@
 Include all accumulated context
-Label planner sample as [SUFFICIENT]
-Create answerer training instance with ground truth answer
+Label planning sample as [SUFFICIENT]
+Create answering training instance with ground truth answer
 The input formatting function (Lines 14 and 17) follows the template:
 This process generates two datasets:
-: Trains the Action Planner to determine retrieval needs and generate targeted subqueries
-: Trains the Answerer to synthesize final responses from accumulated graph context
+: Trains the planning ability to determine retrieval needs and generate targeted subqueries
+: Trains the answering ability to synthesize final responses from accumulated graph context
 Table 4 illustrates the distribution of our generated datasets.
-The planner dataset demonstrates a balanced distribution across three label types: 35% [SUFFICIENT], 55% [SUBQ], and 10% [NO_RETRIEVAL].
-The answerer dataset incorporates both no-retrieval and sufficient cases from multiple sources.
-The size of the answerer dataset exceeds the combined total of no-retrieval and sufficient cases from the planner dataset for two reasons: first, we incorporated additional data from ASQA and Arc-Easy datasets; second, we included no-retrieval cases that were initially filtered out in step 1 of our process.
-This comprehensive approach ensures a robust training set for the answerer component.
+The planning dataset demonstrates a balanced distribution across three label types: 35% [SUFFICIENT], 55% [SUBQ], and 10% [NO_RETRIEVAL].
+The answering dataset incorporates both no-retrieval and sufficient cases from multiple sources.
+The size of the answering dataset exceeds the combined total of no-retrieval and sufficient cases from the planning dataset for two reasons: first, we incorporated additional data from ASQA and Arc-Easy datasets; second, we included no-retrieval cases that were initially filtered out in step 1 of our process.
+This comprehensive approach ensures a robust training set for the answering component.
 Our labeling approach ensures that models learn both the iterative nature of complex question answering and the importance of structured knowledge representation.
@@
 B.4 Dataset Statistics
-| Planner Data | Answerer Data | ||
+| Planning Data | Answering Data | ||
 | # Queries | 129,902 | 78,164 (w/ 3,897 ASQA & 2,147 Arc-Easy) | |
@@
 Appendix C Training Details
-C.1 Classifier & Shifter Training for Theme Scoping
-Classifier Training.
-For theme classification, we trained a multilabel document classifier on the DBPedia-298 dataset (Lehmann et al., 2015) using a two-stage architecture.
-The first stage uses the nomic-embed-text-v1 encoder model (Nussbaum et al., 2024) to generate document embeddings, followed by a classification head consisting of a two-layer neural network.
-The training size and the validation size are 240,932 and 36,008, respectively.
-The classifier consists of:
-An encoder based on nomic-embed-text-v1 for generating document embeddings
-A classification head with:
-Input layer: embedding dimension → 512 units
-ReLU activation
-Dropout (p=0.1)
-Output layer: 512 → number of labels (298)
-Sigmoid activation for multilabel outputs
-We jointly trained both the encoder and classification head with different learning rates:
-Classification head: Trained with a higher learning rate (2e-4) to allow faster adaptation to the classification task
-Text Encoder: Fine-tuned with a lower learning rate (2e-5) to preserve pretrained knowledge while making subtle adjustments for the domain
-This dual learning rate strategy allows the model to benefit from the pretrained knowledge in the encoder while optimizing the classification head for the specific multilabel task.
-The lower learning rate for the encoder helps prevent catastrophic forgetting of the pretrained representations.
-The model was evaluated using micro-averaged metrics on the validation set during training, with monitoring of F1 score, precision, and recall, as shown in Figure 16.
-We select the best classifier model based on the best validation F1 score achieved during training.
-Distribution Shifter Training. The theme distribution shifter maps query theme distributions to expected document theme distributions. Training involves two phases:
-Phase 1: PPD Labeling
-Using our constructed HotpotQA-SUBQ dataset (in Appendix B) with 204,638 query-doc pairs (where we set training-validation ratio as 9:1):
-Input Generation: Process all subqueries and supporting documents through the trained theme classifier to obtain their Posterior Probability Distributions (PPDs)
-Pairing: Each subquery is paired with its corresponding supporting document, creating query-document PPD pairs
-Data Structure: The labeled dataset consists of:
-Input: Query PPDs (298-dimensional probability vectors)
-Target: Corresponding document PPDs (298-dimensional probability vectors)
-Phase 2: Shifter Training
-Train a neural network to map query PPDs to document PPDs:
-Architecture: A three-layer MLP with:
-Input layer: 298 → 512 units
-Hidden layer: 512 → 256 units
-Output layer: 256 → 298 units
-ReLU activations and dropout (p=0.2) between layers
-Softmax activation for final output
-Training Objectives: Multiple distribution alignment metrics:
-KL divergence loss (primary objective)
-Jensen-Shannon divergence monitoring
-Wasserstein distance monitoring
-L1 and L2 distance monitoring
-Training Configuration:
-Optimizer: Adam with learning rate 1e-3
-Batch size: 32
-Early stopping based on validation loss
-Numerical stability ensured through epsilon addition (1e-15)
-Performance is tracked using validation KL divergence, Jensen-Shannon divergence, and Wasserstein distance to ensure proper distribution alignment, as shown in Figure 16.
-C.2 Text-to-Triples Model Training
-For our text-to-triple conversion model, we fine-tuned a Flan-T5-Large model (Chung et al., 2024) to transform raw text passages into structured knowledge triples.
+C.1 Text-to-Triples Model Training
+For our text-to-triple conversion model, we fine-tuned a Flan-T5-Large model [69]222https://huggingface.co/pat-jj/text2triple-flan-t5 and a LLaMA-3.2-3B-Instruct model [70]333https://huggingface.co/pat-jj/text2graph-llama-3.2-3b to transform raw text passages into structured knowledge triples.
 The model processes input text sequences up to 512 tokens in length and generates structured triples in a standardized format “(S subject P predicate O object)”.
-333Our fine-tuned can be found at https://huggingface.co/pat-jj/text2triple-flan-t5 and https://huggingface.co/pat-jj/text2graph-llama-3.2-3b
-Our training dataset WikiOFGraph (Kim et al., 2024a), curated by ChatGPT, comprises 5,851,776 text-triple pairs, with an additional 5,000 samples reserved for validation.
+Our training dataset WikiOFGraph [44], curated by ChatGPT, comprises 5,851,776 text-triple pairs, with an additional 5,000 samples reserved for validation.
 Each training instance consists of a natural language text passage paired with its corresponding set of comma-separated triples.
@@
 The training of this model was conducted on eight NVIDIA RTX 6000 with 48GB memory.
-C.3 Action Planner & Answerer Multi-task Training
-We employ GraphLLM (Perozzi et al., 2024; He et al., 2024) as our foundational model architecture and utilize our constructed HotpotQA-SUBQ dataset to train a unified model capable of performing both graph-conditioned action planning and answer generation through multitask learning.
-Since the Action Planner and Answerer share an identical architectural foundation, as illustrated in Figure 1, we differentiate their functionality through specialized instruction sets.
-For the Action Planner , we employ the following instruction template:
-For the Answerer , we utilize this distinct instruction set:
-For graph encoding, we implement Graph Transformer (Shi et al., 2020), selected for its robust capability in handling non-fully-connected graphs—a common characteristic of text-extracted triples.
-Our base language models comprise LLaMA-2-7B and LLaMA-3-8B, chosen both to maintain consistency with previous research (Asai et al., 2023; Lyu et al., 2024b) and to investigate our framework’s performance scaling across different model capacities.
-Our implementation of GraphLLM differs from G-Retriever (He et al., 2024) primarily due to the distinct nature of our graph structures.
+C.2 Multitask Training of Planning and Answering Capabilities
+We employ Graph LLM [38, 11] as our foundational model architecture and utilize our constructed HotpotQA-SUBQ dataset to train a unified model capable of performing both graph-conditioned action planning and answer generation through multitask learning.
+Since both tasks share an identical architectural foundation, as illustrated in Figure 1, we differentiate their functionality through specialized instruction sets.
+For Planning, we employ the following instruction template:
+For the Answering, we utilize this distinct instruction set:
+For graph encoding, we implement Graph Transformer [65], selected for its robust capability in handling non-fully-connected graphs—a common characteristic of text-extracted triples.
+Our base language models comprise LLaMA-2-7B and LLaMA-3-8B, chosen both to maintain consistency with previous research [30, 31] and to investigate our framework’s performance scaling across different model capacities.
+Our implementation of GraphLLM differs from G-Retriever [11] primarily due to the distinct nature of our graph structures.
 While G-Retriever operates on single interconnected graphs, our framework processes multiple potentially disconnected subgraphs, each corresponding to different subqueries.
@@
 D.1 Test Datasets
-We evaluate RAS on diverse benchmark datasets spanning short-form QA, closed-set tasks, and long-form generation in the zero-shot setting, aligning with (Asai et al., 2023; Lyu et al., 2024b).
-Below we describe each dataset:
+We evaluate RAS on diverse benchmark datasets spanning short-form QA, closed-set tasks, and long-form generation in the zero-shot setting, aligning with [30, 31]. Below we describe each dataset:
 Short-form Generation Datasets:
-TriviaQA-unfiltered (TQA) (Joshi et al., 2017): A large-scale QA dataset containing 11,313 question-answer pairs in our test set.
-The questions are sourced from trivia enthusiasts and cover diverse topics.
-2WikiMultiHopQA (2WQA) (Ho et al., 2020): A multi-hop question answering dataset (with 12,576 samples in test set) that requires models to combine information from multiple Wikipedia articles to answer questions.
-PopQA (Mallen et al., 2022): A dataset focusing on questions about long-tail entities, containing 1,399 queries where the monthly Wikipedia page views are less than 100.
+TriviaQA-unfiltered (TQA) [51]: A large-scale QA dataset containing 11,313 question-answer pairs in our test set. The questions are sourced from trivia enthusiasts and cover diverse topics.
+2WikiMultiHopQA (2WQA) [53]: A multi-hop question answering dataset (with 12,576 samples in test set) that requires models to combine information from multiple Wikipedia articles to answer questions.
+PopQA [52]: A dataset focusing on questions about long-tail entities, containing 1,399 queries where the monthly Wikipedia page views are less than 100.
 These questions test models’ ability to handle queries about less common entities.
 Closed-set Tasks:
-PubHealth (Pub) (Zhang et al., 2023a): A dataset for verifying health-related claims.
-Models must classify statements as ”true” or ”false” based on scientific evidence.
-The dataset contains 987 test samples.
-ARC-Challenge (ARC) (Clark et al., 2018): A multiple-choice science question dataset designed to be challenging for models, requiring multi-hop reasoning and background knowledge.
+PubHealth (Pub) [54]: A dataset for verifying health-related claims. Models must classify statements as "true" or "false" based on scientific evidence. The dataset contains 987 test samples.
+ARC-Challenge (ARC) [55]: A multiple-choice science question dataset designed to be challenging for models, requiring multi-hop reasoning and background knowledge.
 The dataset contains 1,172 test samples.
 Long-form Generation Datasets:
-ASQA (Stelmakh et al., 2022): A long-form question answering dataset that requires models to generate comprehensive answers with proper citation of sources.
+ASQA [57]: A long-form question answering dataset that requires models to generate comprehensive answers with proper citation of sources.
 Models must balance information completeness with factual accuracy.
 The dataset contains 948 test samples.
-ELI5 (Fan et al., 2019): A dataset derived from the ”Explain Like I’m Five” subreddit, containing questions seeking straightforward explanations of complex topics.
-The dataset contains 1,000 test samples.
+ELI5 [58]: A dataset derived from the "Explain Like I’m Five" subreddit, containing questions seeking straightforward explanations of complex topics. The dataset contains 1,000 test samples.
 D.2 Evaluation Metrics
@@
 Short-form Generation:
-For PopQA and TriviaQA, we use the ”golden match” metric (Asai et al., 2023; Min et al., 2019; Guu et al., 2020)444Implementation aligned with https://github.com/AkariAsai/self-rag/blob/main/retrieval_lm/metrics.py, where a prediction is considered correct if it contains any normalized version of the ground truth answers :
+For PopQA and TriviaQA, we use the "golden match" metric [30, 71, 8]444Implementation aligned with https://github.com/AkariAsai/self-rag/blob/main/retrieval_lm/metrics.py, where a prediction is considered correct if it contains any normalized version of the ground truth answers :
 (9) where normalizes text by lowercasing, removing articles and punctuation.
-For 2WikiMultiHopQA, we follow RPG (Lyu et al., 2024b)555Implementation aligned with https://github.com/haruhi-sudo/RPG/blob/main/retriever/src/evaluation.py to use token-level F1 score between prediction and ground truth :
+For 2WikiMultiHopQA, we follow RPG [31]555Implementation aligned with https://github.com/haruhi-sudo/RPG/blob/main/retriever/src/evaluation.py to use token-level F1 score between prediction and ground truth :
 (10) where precision and recall are computed over normalized token overlap:
 (11) (12)
-Closed-set Tasks: For both PubHealth and ARC-Challenge, we use accuracy (Asai et al., 2023), computed as:
+Closed-set Tasks: For both PubHealth and ARC-Challenge, we use accuracy [30], computed as:
 where is the total number of examples.
@@
 This standardized format enables reliable knowledge accumulation and reasoning in subsequent stages of the RAS framework.
-For Action Planning (Figure 21), we design a comprehensive prompt that guides the model in making strategic decisions about information retrieval needs.
+For Planning (Figure 19), we design a comprehensive prompt that guides the model in making strategic decisions about information retrieval needs.
 The prompt instructs the model to output one of three labels based on careful analysis of the current knowledge state: (1) [NO_RETRIEVAL] when the question can be answered directly, either due to the model’s inherent knowledge or the question’s nature, (2) [SUBQ] accompanied by a focused subquery when additional information is needed, or (3) [SUFFICIENT] when the accumulated knowledge is adequate to answer the question.
@@
 This design helps maintain the system’s efficiency while ensuring comprehensive knowledge gathering.
-For Answer Generation (Figure 22), we design a prompt that focuses on synthesizing precise answers from structured knowledge graphs.
+For Answering (Figure 20), we design a prompt that focuses on synthesizing precise answers from structured knowledge graphs.
 The prompt emphasizes selective use of retrieved information - instructing the model to utilize subqueries and graph information only when relevant to the question at hand.
@@
 The prompt strictly enforces concise answer generation, requiring only the essential information without additional explanation or commentary.
-In addition, for ASQA and ELI5 with closed-source models (RAS and all the other baselines (e.g., SuRe) using Sonnet-3.5), we conduct few-shot inference with two in-context learning demonstrations, aligning with previous study’s (Gao et al., 2023) implementation.777https://github.com/princeton-nlp/ALCE/blob/main/run.py888Exemplars can be found at https://github.com/princeton-nlp/ALCE/tree/main/prompts
+In addition, for ASQA and ELI5 with closed-source models (RAS and all the other baselines (e.g., SuRe) using Sonnet-3.5), we conduct few-shot inference with two in-context learning demonstrations, aligning with previous study’s [56] implementation.777https://github.com/princeton-nlp/ALCE/blob/main/run.py888Exemplars can be found at https://github.com/princeton-nlp/ALCE/tree/main/prompts
 E.2 Open-Source Model Settings
@@
 We use Python 3.10, PyTorch 2.1.2, and torch-geometric 2.6.1 throughout all experiments.
-For more efficient text retrieval, we split both theme index and dense index of the corpus into eight splits, and load them on eight GPUs with faiss-gpu 1.7.2.999All the Wikipedia dumps (2018, 2020) we used are downloaded from https://github.com/facebookresearch/atlas
-We set maximum new tokens as 100 for PopQA, TriviaQA, and 2WikiMultihopQA, as 50 for PubHealth and ARC-Challenge, and as 300 for ASQA and ELI5, aligning with previous study (Asai et al., 2023; Lyu et al., 2024b).
+For more efficient text retrieval, we split the dense index of the corpus into eight splits, and load them on eight GPUs with faiss-gpu 1.7.2.999All the Wikipedia dumps (2018, 2020) we used are downloaded from https://github.com/facebookresearch/atlas
+We set maximum new tokens as 100 for PopQA, TriviaQA, and 2WikiMultihopQA, as 50 for PubHealth and ARC-Challenge, and as 300 for ASQA and ELI5, aligning with previous study [30, 31].
 All the generation are configured with batch size of 1.
-As we always set the main query as the first subquery, we pre-extract the triples from the context at the first iteration using our Flan-T5-based text-to-triple model for batch triple extraction.
-Notably, when we experimented with using Claude 3.5 Sonnet for triple extraction using the prompt from Figure 20, we observed an additional 3% average performance gain.
-This suggests room for improvement in our triple extraction approach - one potential direction would be training a larger model like LLaMA-2-7B on the WikiOFGraph dataset to achieve higher-quality triple extraction.
+As we always set the main query as the first subquery, we pre-extract the triples from the context at the first iteration using our text-to-triple model for batch triple extraction.
 E.3 Instructions
@@
 | ELI5 | Provide a paragraph-length response using simple words to answer the following question. [Long Form] |
-Appendix F Qualitative Examples
+Appendix F Supplementary Results and Analysis
+Higher Performance with More Contexts?
+We further examine the performance of Sonnet-3.5 when increasing the number of retrieved documents. Table 6 compares Sonnet-3.5 under different retrieval depths with RAS-Sonnet.
+| Method | 2WikiMultihopQA (F1) | PopQA (Acc) | PubHealth (Acc) | ARC (Acc) |
+| Sonnet-3.5 (#docs=5) | 53.7 | 57.3 | 53.9 | 87.1 |
+| Sonnet-3.5 (#docs=10) | 55.1 | 58.5 | 56.6 | 87.9 |
+| Sonnet-3.5 (#docs=15) | 55.4 | 58.0 | 56.9 | 86.9 |
+| RAS (Sonnet) | 57.7 | 62.3 | 71.4 | 93.8 |
+Increasing retrieval depth yields diminishing returns for Sonnet-3.5, while RAS-Sonnet achieves significantly higher performance with structured, targeted context.
+This highlights the limitations of passive context scaling and the benefits of RAS’s organized retrieval strategy.
+Error Analysis on Graph Construction
+We present two representative cases where the quality of retrieved structured knowledge significantly impacts final answer generation.
+These examples are distinct from the qualitative examples shown in Section G and highlight both successful and failed reasoning paths.
+Case 1: Accurate Reasoning with Well-Structured Graph
+Question: Who directed the film adaptation of the novel “Tinker Tailor Soldier Spy”?
+Subqueries and Graphs:
+[SUBQ] Who directed the film “Tinker Tailor Soldier Spy”?
+(S> Tinker Tailor Soldier Spy | P> Film adaptation director | O> Tomas Alfredson),
+(S> Tinker Tailor Soldier Spy | P> Release year | O> 2011) -
+[SUBQ] Who is Tomas Alfredson?
+(S> Tomas Alfredson | P> Occupation | O> Film director)
+Answer: Tomas Alfredson (Correct)
+Analysis: The subquery identifies the correct adaptation and its director. The structured triple precisely encodes this relation, allowing the model to reason accurately.
+Case 2: Reasoning Failure Due to Spurious or Missing Triples
+Question: What invention is Nikola Tesla best known for?
+Subqueries and Graphs:
+[SUBQ] What are Nikola Tesla’s major inventions?
+(S> Nikola Tesla | P> Known for | O> Earthquake machine),
+(S> Nikola Tesla | P> Proposed | O> Death ray),
+(S> Nikola Tesla | P> Created | O> Tesla coil) -
+[SUBQ] What is a Tesla coil used for?
+(S> Tesla coil | P> Used in | O> High-frequency experiments)
+Answer: Earthquake machine or death ray (Incorrect)
+Analysis: The graph omits the key triple (Nikola Tesla | Invented | Alternating current system), instead surfacing fringe inventions.
+This misleads the model into emphasizing less impactful or speculative work, leading to an incorrect answer.
+These examples reveal that RAS makes reasoning pathways more interpretable by structuring retrieved knowledge explicitly.
+When failures occur, they can often be traced to identifiable issues in the extracted triples, such as missing key relations or overemphasis on fringe facts, rather than opaque model behavior.
+This transparency not only facilitates error analysis but also suggests concrete avenues for improving upstream components like triple extraction or subquery generation, making RAS a flexible foundation for future refinement.
+Appendix G Qualitative Examples
 In this section, we show some qualitative examples and running examples of RAS.
-Appendix G Hyperparameter Study
+Appendix H Hyperparameter Study
 | Hyper-parameter | Studied Values |
-| Theme Classifier | |
-| Encoder learning rate | {1e-6, 1e-5, 2e-5, 1e-4, 2e-4} |
-| Classifier learning rate | {1e-6, 1e-5, 2e-5, 1e-4, 2e-4, 5e-4, 1e-3} |
-| Batch size | {4, 8, 16, 32, 64} |
-| Hidden layers | {1, 2, 3} |
-| Hidden dim | {256, 512, 1024} |
-| Dropout rate | {0.1, 0.2, 0.3} |
-| Distribution Shifter | |
-| Batch size | {4, 8, 16, 32, 64} |
-| Learning rate | {0.0001, 0.0005, 0.001, 0.005} |
-| Hidden layers | {[256], [512], [512, 256], [1024, 512, 256]} |
-| Dropout rate | {0.1, 0.15, 0.2, 0.25, 0.3} |
-| Optimizer | {Adam, SGD, RMSprop, AdamW} |
-| Loss function | {KLDiv, JSD, MSE, Wasserstein, CrossEntropy} |
 | Text-to-Triples Model | |
@@
 | Others | |
-| Theme scoping top- | {100,000, 250,000, 500,000 } |
 | Dense retrieval top- | 5 |
-Appendix H Notation Table
-| Notation | Description |
-| General | |
-| Main input query | |
-| Generated answer | |
-| Complete document corpus | |
-| Theme-indexed document corpus | |
-| Action Planning | |
-| Action planner model | |
-| Plan generated at iteration | |
-| Subquery generated at iteration | |
-| Graph information (triple list) extracted at iteration | |
-| Theme-Scoped Retrieval | |
-| Theme classifier function | |
-| Theme distribution shifter function | |
-| Dense retriever function | |
-| Posteriori probability distribution of query | |
-| Expected document theme distribution | |
-| Theme-specific sub-corpus at iteration | |
-| Retrieved text passages at iteration | |
-| Number of documents selected in theme scoping | |
-| Number of passages selected in dense retrieval | |
-| Knowledge Structuring | |
-| Text-to-triple conversion model | |
-| Text encoder for node/edge attributes | |
-| Graph neural network encoder | |
-| Knowledge graph at iteration | |
-| Final query-specific knowledge graph | |
-| Node and edge sets of graph at iteration | |
-| Node and edge sets of final graph | |
-| Attribute vector of node | |
-| Attribute vector of edge | |
-| Answer Generation | |
-| Answerer model | |
-| Model-generated answer before retrieval | |
+Appendix I Broader Impacts, Safeguards, and Assets
+Broader Impacts.
+The RAS framework presents several promising impacts for both research and practical applications.
+By transforming unstructured retrieved content into structured, query-specific knowledge graphs, RAS improves reasoning transparency and factual reliability—qualities that are especially valuable in education, scientific research, and technical writing.
+The ability to explicitly organize retrieved knowledge into interpretable structures also opens up new possibilities for human-in-the-loop AI systems, enabling users to trace and verify model reasoning paths.
+However, several ethical considerations remain.
+RAS’s reliance on structured knowledge representations means its effectiveness may be limited in low-resource languages or domains with sparse data.
+Additionally, biases present in training data can be preserved or even amplified in structured forms.
+Addressing these concerns will require rigorous evaluation across diverse user groups and careful curation of source corpora.
+Promoting fairness, transparency, and inclusivity should be core design goals for future RAS-based systems.
+Safeguards.
+To mitigate potential harms, we propose several safeguards during the development and deployment of RAS-based systems.
+First, developers should routinely audit training and retrieval corpora for biases and factual inconsistencies, and apply data augmentation or counterfactual generation techniques where appropriate.
+Second, regular evaluations should be conducted across a diverse range of tasks, languages, and user demographics to assess any unintended performance disparities.
+Third, when applied in sensitive domains, human-in-the-loop oversight should be incorporated to validate the structured knowledge outputs and final generations.
+Lastly, transparency measures, such as providing users with access to intermediate structured knowledge (e.g., the constructed knowledge graph), can help foster trust and enable error analysis.
+Assets Used.
+To support the development and evaluation of the RAS framework, we utilized a wide range of publicly available datasets, open-source tools, and pretrained models.
+All datasets were selected based on relevance to knowledge-intensive tasks, including question answering and long-form generation, and were licensed for academic research or distributed under permissive open-source terms such as MIT, Apache 2.0, and CC BY.
+Our training and evaluation relied on benchmark datasets such as HotpotQA, ASQA, TriviaQA, and ARC, while structured knowledge was derived using models trained on WikiOfGraph.
+We used open-source infrastructure including FAISS for retrieval, vLLM for inference optimization, and Graph LLM as our model backbone.
+For baseline comparisons, we incorporated implementations from Self-RAG, RPG, and ALCE.
+Closed-source models like Claude-3.5-Sonnet were used for performance comparison and triple extraction under API constraints.
+Full licensing and usage details are provided in Table 8.
+| Asset | Source / URL | License | Usage |
+| HotpotQA | https://hotpotqa.github.io/ | MIT | Base QA dataset |
+| HotpotQA-SUBQ | We constructed from HotpotQA | MIT (inherited) | Subquery training dataset |
+| ASQA | https://github.com/google-research/language/tree/master/language/asqa | Apache 2.0 | Training and evaluation |
+| Arc-Easy | https://allenai.org/data/arc | AI2 Research License | Training (subset) |
+| TriviaQA | https://nlp.cs.washington.edu/triviaqa/ | CC BY-SA 4.0 | Evaluation |
+| PopQA | https://huggingface.co/datasets/akariasai/PopQA | CC BY 4.0 | Evaluation |
+| 2WikiMultihopQA | https://github.com/Alab-NII/2wikimultihop | Apache 2.0 | Evaluation |
+| PubHealth | https://huggingface.co/datasets/bigbio/pubhealth | MIT | Evaluation |
+| ARC-Challenge | https://allenai.org/data/arc | AI2 Research License | Evaluation |
+| ELI5 | https://huggingface.co/datasets/eli5 | Apache 2.0 | Evaluation |
+| Wikipedia 2018/2020 | https://github.com/facebookresearch/atlas | Creative Commons Public Licenses | Knowledge source for retrieval |
+| WikiOfGraph | https://github.com/daehuikim/WikiOFGraph | CC BY 4.0 | Triple model training |
+| Claude-3.5-Sonnet | https://www.anthropic.com | Proprietary | Default closed source model |
+| LLaMA-2/3 | https://ai.meta.com/llama | Meta LLaMA Community License | Default open source model |
+| LLaMA-3.2-3B | Derived from Meta LLaMA | Meta LLaMA Community License | Triple extraction |
+| Flan-T5-Large | https://huggingface.co/google/flan-t5-large | Apache 2.0 | Triple extractor |
+| Graph LLM | Modified from https://github.com/XiaoxinHe/G-Retriever | Research use | Model backbone |
+| Faiss | https://github.com/facebookresearch/faiss | MIT | Dense retrieval |
+| vLLM | https://github.com/vllm-project/vllm | Apache 2.0 | Optimized inference |
+| Sentence-BERT | https://www.sbert.net | Apache 2.0 | Graph node/edge embedding |
+| Self-RAG Code | https://github.com/AkariAsai/self-rag | MIT | Baseline eval |
+| RPG Code | https://github.com/haruhi-sudo/RPG | Research use | Baseline eval |
+| ALCE Code | https://github.com/princeton-nlp/ALCE | MIT | Eval metrics |

### v3

@@
 Recent interpretability studies highlighting the importance of structured intermediate reasoning further align with this perspective.
-We propose Retrieval-And-Structuring (RAS), a framework that dynamically constructs query-specific knowledge graphs through iterative retrieval and structured knowledge building.
+We propose Retrieval-And-Structuring (RAS), a framework that dynamically constructs question-specific knowledge graphs through iterative retrieval and structured knowledge building.
 RAS interleaves targeted retrieval planning with incremental graph construction, enabling models to assemble and reason over evolving knowledge structures tailored to each query.
-On seven knowledge-intensive benchmarks, RAS consistently outperforms strong baselines, achieving up to 6.4% and 7.0% gains with open-source and proprietary LLMs, respectively.
-Our results demonstrate that dynamic, query-specific knowledge structuring offers a robust path to improving reasoning accuracy and robustness in language model generation.
-Our data and code can be found at https://github.com/pat-jj/RAS.
+On seven knowledge-intensive benchmarks, RAS consistently outperforms strong baselines, achieving up to 8.7% and 7.0% gains with proprietary and open-source LLMs, respectively.
+Our results demonstrate that dynamic, question-specific knowledge structuring offers a robust path to improving reasoning accuracy and robustness in language model generation.
 1 Introduction
-Complex reasoning tasks such as scientific analysis or multi-hop question answering demand both comprehensive knowledge and structured logical thinking [1].
-While large language models (LLMs) have achieved remarkable performance across a wide range of natural language processing tasks [2, 3, 4], they often struggle with knowledge-intensive reasoning due to the absence of precise, logically organized information [5, 6].
-This limitation has motivated growing research into augmenting LLMs with structured knowledge to enhance their reasoning capabilities [7].
-Retrieval-augmented generation (RAG) approaches have emerged as a promising direction by providing LLMs with additional context from retrieved passages [8, 9, 10, 11].
-However, these methods frequently face hallucination challenges [12, 13], where generated content deviates from retrieved information.
-This issue stems primarily from the unstructured nature of retrieved passages, requiring the model to implicitly bridge gaps during generation.
-Recent research in model interpretability further highlights this issue [14].
-Attribution graph analyses have revealed that LLMs internally attempt multi-hop reasoning by chaining facts across different parts of context, but failures in these reasoning pathways are correlated with hallucinations and factual inconsistencies.
-These insights suggest that explicitly providing structured intermediate knowledge could significantly enhance the reliability and transparency of LLM reasoning.
-Recent efforts have explored integrating knowledge graphs (KGs) with LLMs [15, 16, 11, 17] to provide compact, relational representations that enable more interpretable reasoning [18, 19, 20].
-However, existing approaches largely rely on static, corpus-wide graphs, which may not adapt well to the specific reasoning needs of individual queries.
-Constructing and maintaining large-scale KGs also incurs substantial costs in open-domain settings.
-To address these challenges, we propose Retrieval-And-Structuring (RAS), a framework that dynamically constructs and reasons over query-specific knowledge graphs through iterative retrieval and structured knowledge building.
-RAS interleaves three key components: (1) a planning module that strategically identifies knowledge gaps and generates targeted sub-queries, (2) a retrieval-and-structuring module that extracts factual triples from retrieved passages and incrementally builds a query-specific graph, and (3) a knowledge-augmented answering module that generates outputs conditioned on the accumulated structured knowledge.
-RAS addresses several critical limitations of prior methods.
-Unlike traditional RAG approaches that rely on single-pass retrieval [8, 9, 10], RAS employs iterative planning to actively identify and fill knowledge gaps during inference.
-In contrast to static KG-based approaches [11, 17], RAS dynamically constructs query-specific knowledge graphs tailored to the evolving reasoning needs of each query.
-This targeted construction captures only task-relevant information, avoiding the inefficiencies and noise associated with static, corpus-level graphs, which enables more precise and robust language generation.
-Through extensive evaluations across seven benchmarks spanning open-domain QA, closed-set QA, and long-form generation, RAS consistently outperforms strong baselines by 6.4% with open-source LLMs and 7.0% with proprietary models.
+Complex reasoning tasks such as scientific analysis or multi-hop question answering demand both comprehensive knowledge and structured logical thinking (Yang et al., 2018).
+While large language models (LLMs) have achieved remarkable performance across a wide range of natural language processing tasks (Devlin et al., 2018; Brown et al., 2020), they often struggle with knowledge-intensive reasoning due to the absence of precise, logically organized information (Rae et al., 2021; Ling et al., 2024).
+This limitation has motivated growing research into augmenting LLMs with structured knowledge to enhance their reasoning capabilities (Wang et al., 2021).
+Retrieval-augmented generation (RAG) approaches provide LLMs with additional context from retrieved passages (Guu et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021; He et al., 2024), but often face hallucination challenges (Maynez et al., 2020; Zhang et al., 2023b), where generated content deviates from retrieved information.
+This issue stems from the unstructured nature of passages, which forces the model to implicitly bridge logical gaps.
+Briefly, interpretability analyses have suggested that LLMs attempt to chain facts across context internally, and failures in these implicit reasoning chains correlate with hallucinations (Lindsey et al., 2025).
+These findings reinforce the need for explicitly structured intermediate knowledge to guide reasoning.
+Recent efforts have integrated knowledge graphs (KGs) with LLMs (Sun et al., 2019; Yu et al., 2022; He et al., 2024; Edge et al., 2024), providing compact relational representations that support more interpretable reasoning (Hogan et al., 2021; Jiang et al., 2024; Sun et al., 2023).
+However, existing approaches typically rely on static, corpus-wide graphs.
+This design introduces two limitations.
+First, global KGs are costly to build and maintain: indexing a corpus like Wikipedia 2018 can require millions of LLM calls and cost tens of thousands to millions of USD (see Appendix G).
+Second, global graphs often blend evidence from many documents, leading to ambiguous or even contradictory relations.
+For example, a global KG might simultaneously encode that Geoffrey Hinton is linked to deep learning, to cognitive neuroscience, and to critiques of large models—without clarifying which aspect is relevant to user query’s focus.
+Similarly, biomedical KGs may contain both positive and negative associations between a drug and a disease, reflecting conflicting studies.
+In contrast, a question-specific KG built from targeted documents resolves these conflicts by grounding relations in a coherent, query-relevant context.
+These limitations highlight the need for knowledge graphs that are constructed on demand, tailored to the query, and structured to support reasoning.
+To this end, we propose Retrieval-And-Structuring (RAS), a framework that dynamically constructs and reasons over question-specific knowledge graphs through iterative retrieval and structured knowledge building.
+The RAS process unfolds in three steps: (1) a planning step that identifies knowledge gaps and generates targeted sub-queries, (2) a retrieval-and-structuring step that extracts factual triples from retrieved passages and incrementally builds a question-specific graph, and (3) a knowledge-augmented answering step that produces final outputs conditioned on the accumulated structured knowledge.
+RAS addresses several limitations of prior methods.
+Unlike traditional RAG, which performs single-pass retrieval (Guu et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021), RAS iteratively plans and fills knowledge gaps at inference.
+In contrast to static KG-based approaches (He et al., 2024; Edge et al., 2024), RAS dynamically constructs question-specific graphs tailored to each question, capturing only task-relevant information.
+This design avoids both the inefficiency of offline indexing and the noise of global graphs, enabling precise and robust reasoning.
+Through extensive evaluations across seven benchmarks spanning open-domain QA, closed-set QA, and long-form generation, RAS consistently outperforms strong baselines by 7.0% with open-source LLMs and 8.7% with proprietary models.
 Our main contributions are:
-We introduce RAS, a novel framework for dynamic knowledge structuring through iterative retrieval and graph construction, tailored to each query.
-We develop a structure-aware planning and answering model that operates over evolving knowledge graphs.
-We demonstrate RAS’s effectiveness through comprehensive experiments and ablation studies across diverse knowledge-intensive tasks.
+We propose RAS, a framework that dynamically builds question-specific knowledge graphs through iterative retrieval and structuring.
+We design a unified graph structure-aware model that jointly plans retrieval and generates answers over evolving knowledge graphs.
+We show consistent gains across seven benchmarks, with up to 8.7% improvement over strong RAG baselines, while maintaining efficiency and scalability.
 2 Related Work
-Retrieval-Augmented Generation (RAG) [8, 9] enhances language model performance on knowledge-intensive tasks by incorporating retrieved passages into the model input, improving factual accuracy and grounding.
-Early RAG approaches retrieved a fixed number of passages once before generation [21, 22, 23], while later methods explored adaptive retrieval [24], where models dynamically issue retrieval queries based on uncertainty signals, or retrieval evaluation [25] to improve relevance.
-However, these techniques often struggle with complex or multi-hop reasoning due to their linear designs.
-Recent advances aim to strengthen retrieval quality by iterative refinement.
-Iterative retrieval-generation approaches [21, 26] leverage historical context to guide subsequent retrievals, while targeted retrieval strategies based on subqueries [27] or follow-up questions [28, 29] progressively enrich the retrieved evidence.
-Self-RAG [30] introduced self-reflective retrieval, allowing LLMs to selectively query for missing information and evaluate the quality of retrievals.
-RPG [31] further refined retrieval by extracting fine-grained paragraphs to enhance relevance.
-Despite these advances, existing methods face persistent challenges: retrieved context often contains redundant or irrelevant information, or conversely, discards essential facts during refinement.
-To address these issues, we propose converting retrieved content into a structured graph format and maintaining an iteratively enriched, query-specific knowledge graph, allowing the model to reason over an organized and progressively expanded knowledge structure.
+Retrieval-Augmented Generation (RAG).
+RAG enhances language model performance on knowledge-intensive tasks by incorporating retrieved passages into the model input (Guu et al., 2020; Lewis et al., 2020), improving factual accuracy and grounding.
+Early approaches retrieved a fixed number of passages once before generation (Shao et al., 2023; Es et al., 2024; Lyu et al., 2024a), while later methods explored adaptive retrieval (Jiang et al., 2023) or retrieval evaluation (Kim et al., 2024b) to improve relevance.
+Iterative retrieval-generation approaches (Shao et al., 2023; Guan et al., 2024) and targeted subquery strategies (Khattab et al., 2023; Yao et al., 2023; Press et al., 2022; Trivedi et al., 2023) progressively enrich the evidence.
+Self-RAG (Asai et al., 2023) introduced self-reflective retrieval, and RPG (Lyu et al., 2024b) extracted fine-grained paragraphs.
+More recent work (Jiang et al., 2025c; Mei et al., 2025) further improves RAG by reinforcement learning over search behaviors.
+Despite these advances, retrieved context often contains redundancy or misses critical facts.
+Our work departs from these by converting retrieved content into a structured, evolving graph aligned with the query.
 Graph as Context for LLMs.
-Integrating structured graphs into language model workflows has emerged as a powerful approach for enhancing reasoning and factual accuracy [32, 33, 34, 35, 36, 37].
-Graphs provide explicit, relational representations that help models go beyond flat text by making entity connections and multi-hop relationships more tractable.
-Recent work, such as GraphToken [38], shows that LLMs can process serialized graph inputs through token-level prompts [39] without modifying model parameters.
-G-Retriever [11] leverages global knowledge graphs (KGs) to retrieve entity-centric subgraphs, while GraphRAG-style methods [17, 40] construct corpus-level graphs and generate summaries to enhance retrieval quality.
-Despite their benefits, these methods rely heavily on static, global graphs, which often include irrelevant or redundant information and cannot adapt to the needs of individual queries.
-Moreover, building and maintaining large-scale KGs is computationally expensive, particularly in open-domain settings.
-This motivates a shift toward dynamic, inference-time graph construction.
-Our approach embraces this direction by proposing a retrieval-and-structuring framework that builds query-specific knowledge graphs on the fly.
-Rather than relying on fixed knowledge coverage, we iteratively retrieve and structure information into task-relevant graphs that evolve with the reasoning process.
-This dynamic graph construction aligns with recent interpretability findings [14], which suggest that LLM failures often stem from breakdowns in implicit reasoning chains—failures that explicit graph structuring can help mitigate.
+Graphs offer explicit, relational structures that help models go beyond flat text by making multi-hop relationships more tractable (Yasunaga et al., 2021; 2022; Yu, 2022; Ju et al., 2022; Zhu et al., 2024; Gutiérrez et al., 2024).
+GraphToken (Perozzi et al., 2024) shows LLMs can process serialized graphs directly (Liu et al., 2021).
+G-Retriever (He et al., 2024) leverages global KGs for entity-centric subgraph retrieval, while GraphRAG-style methods (Edge et al., 2024; Jiang et al., 2025b) construct large corpus-level graphs with community summarization.
+These methods rely on static graphs, which are costly to construct (Appendix G) and often introduce irrelevant noise.
+By contrast, RAS builds question-specific knowledge graphs dynamically, eliminating prohibitive offline costs and providing denser, task-relevant context tailored to each reasoning trajectory.
+This design aligns with findings that many LLM errors stem from failed implicit reasoning chains (Lindsey et al., 2025), which explicit, query-focused structuring can mitigate.
 3 Retrieval-And-Structuring (RAS) Framework
 Effective knowledge-intensive language generation requires not only retrieving relevant information, but also structuring and reasoning over it systematically.
-We introduce Retrieval-And-Structuring (RAS), a framework that interleaves iterative retrieval planning with dynamic knowledge graph construction, enabling large language models (LLMs) to reason over progressively organized knowledge tailored to each query.
+We introduce Retrieval-And-Structuring (RAS), a framework that interleaves iterative retrieval planning with dynamic question-specific knowledge graph construction, enabling large language models (LLMs) to reason over progressively organized knowledge tailored to each query.
 Figure 1 illustrates the overall workflow.
@@
 We define the core concepts used in the RAS framework as follows.
-The Main Query () denotes the original task input.
+The Main question () denotes the original task input.
 A subquery () is a focused retrieval query generated at iteration to obtain supporting evidence.
-Retrieved Passages () are the top- documents selected for each subquery.
-A text-to-triples model converts retrieved text into Extracted Triples (), structured as subject-predicate-object facts.
-These triples are incrementally accumulated into an evolving Knowledge Graph (), representing organized evidence about .
-The model produces an Action Plan () at each step, determining whether to continue retrieval ([SUBQ]), stop retrieval ([SUFFICIENT]), or, initially, answer directly without retrieval ([NO_RETRIEVAL]).
-3.1 Knowledge State-Aware Planning
-The Planning stage initiates and controls the retrieval-and-structuring process by dynamically assessing the current knowledge state.
+At the first iteration, .
+Retrieved text () is a set of the top- documents retrieved by .
+A text-to-triples model converts retrieved text into triples (), structured as subject-predicate-object facts.
+These triples are incrementally accumulated into an evolving question-specific knowledge graph (), representing organized evidence related to .
+The model produces an plan () at each step, determining whether to continue retrieval ([SUBQ]), stop retrieval ([SUFFICIENT]), or, initially, answer directly without retrieval ([NO_RETRIEVAL]).
+3.1 Knowledge-Aware Planning
+The planning step initiates and controls the retrieval-and-structuring process by dynamically assessing the current knowledge state.
 Initial Planning. Formally, given an input query , the model generates an initial plan :
 where is the planning instruction (as shown in Figure 15). can take one of two forms:
-[SUBQ] : If assesses that the query cannot be satisfactorily answered with its own knowledge, we start the iteration with the main query as the initial subquery, and move to the next stage (§3.2).
+[SUBQ] : If assesses that the query cannot be satisfactorily answered with its own knowledge, we start the iteration with the main question as the initial subquery, and move to the next stage (§3.2).
 [NO_RETRIEVAL]: If determines that can be answered directly without requiring any additional knowledge, the planning process terminates, and the framework proceeds directly to the final Answering stage (§3.3).
@@
 The framework proceeds to the next stage (§3.2).
-[SUFFICIENT]: The accumulated knowledge is deemed sufficient to comprehensively address the main query .
+[SUFFICIENT]: The accumulated knowledge is deemed sufficient to comprehensively address the main question .
 The iterative retrieval process terminates, and the framework proceeds to the Answering stage (§3.3).
-The Planning stage serves as a key driver of the RAS framework’s iterative retrieval and refinement process.
+Planning serves as a key driver of the RAS framework’s iterative retrieval and refinement process.
 By dynamically assessing the adequacy of the retrieved knowledge and generating targeted sub-queries, it enables the efficient acquisition of query-relevant information.
-3.2 Text Retrieval and Knowledge Structuring
-This stage transforms retrieved text into structured knowledge and progressively builds the query-specific graph .
+3.2 Text Retrieval and Structuring
+Once [SUBQ] is detected, we use the subquery to retrieve the text and transform it into structured knowledge , which is progressively merged to the question-specific graph .
 Text Retrieval. We use a text retriever to retrieve the top- semantically relevant passages from the corpus for each subquery :
-We use a standard dense retriever by default but note that RAS is compatible with more advanced information retrieval methods [41, 42, 43].
-Text-to-Triples (Graph) Conversion.
+We use a standard dense retriever by default but note that RAS is compatible with more advanced information retrieval methods (Chaudhary et al., 2023; Kang et al., 2024; Jiang et al., 2025a).
+Text-to-Triples Conversion.
 To extract essential factual information from the retrieved passages , we employ a text-to-triples model .
-This model is trained on the full WikiOfGraph dataset [44], which is a high-quality, LLM-curated text-to-triples corpus.
-Details of the training process are provided in Appendix C.1.
+This model is trained on the full WikiOfGraph dataset (Kim et al., 2024a), which is a high-quality, LLM-curated text-to-triples corpus.
+Details of the training process are provided in Appendix D.1.
 The model generates structured triples in the following format:
@@
 This structured representation enables efficient downstream reasoning and facilitates integration with external knowledge graphs.
-Although is a lightweight LLM capable of fast inference using techniques such as quantization [45] and optimized inference frameworks like vLLM [46], the text-to-triples conversion can also be precomputed offline when maximal efficiency is required.
-Iterative Knowledge Enrichment to Query-Specific KG.
+Although is a lightweight LLM capable of fast inference using techniques such as quantization (Dettmers et al., 2022) and optimized inference frameworks like vLLM (Kwon et al., 2023), the text-to-triples conversion can be precomputed offline as well when maximal efficiency is required.
+Iterative Knowledge Enrichment to Question-Specific KG.
 The extracted triples are then converted into a graph structure , where and denote the sets of nodes and edges, respectively.
 Each node corresponds to a unique subject or object entity in , while each edge represents a predicate connecting two entities.
-To enrich the graph with semantic information, the attributes of nodes and edges are obtained through a text encoder (Sentence-BERT [47]):
-These semantic embeddings enable the model to capture the nuanced relationships between entities and facilitate reasoning over the knowledge graph.
-To progressively enrich the query-related knowledge in response to the evolving sub-queries, the structured graph at each iteration is merged into an evolving KG specific to the main query :
-After enriching with the new knowledge, the Planning module (§3.1) reassesses the current knowledge to determine the next step.
+To enrich the graph with semantic information, the attributes of nodes and edges are obtained through Sentence-BERT (Reimers, 2019):
+These semantic embeddings enable the model to capture the nuanced relationships between entities and facilitate reasoning over the KG.
+To progressively enrich the question-related knowledge in response to the evolving sub-queries, the structured graph at each iteration is merged into an evolving KG specific to the main question :
+After enriching with the new knowledge, we plan (§3.1) for the next step.
 Based on and the chain of previous subqueries and their associated graph information, the model decides whether to generate another focused subquery for additional retrieval or to proceed with answering (§3.3) if the accumulated knowledge is sufficient.
 3.3 Knowledge-Augmented Answering
-At the final stage, the model generates an answer to the main query either conditioned on knowledge graph and subquery chain when retrieval-and-structuring was processed, or directly when no retrieval is needed.
+When answering is triggered, the model generates an answer to the main question either conditioned on knowledge graph and subquery chain when retrieval-and-structuring was processed, or directly when no retrieval is needed.
 If no retrieval is needed (), the answer is generated directly:
-Otherwise, after iterative knowledge enrichment concludes with [SUFFICIENT] plan, the answer is generated using encoded KG and subquery chain:
+Otherwise, after iterative knowledge enrichment concludes with [SUFFICIENT] plan or the maximum iteration is reached, the answer is generated using encoded KG and subquery chain:
 where is the answering instruction (as shown in Figure 16).
@@
 3.4 Sturcture-Aware Multitask Learning
-The RAS framework is trained through a multitask setup that unifies knowledge planning and knowledge-augmented answering under a standard next-token prediction objective.
+The RAS framework is trained through a multitask setup that unifies knowledge-aware planning and knowledge-augmented answering under a standard next-token prediction objective.
 Each training instance corresponds to either a planning task or an answering task, selected randomly:
-Planning Task: Input: (encoded ), planning instruction , subquery-triples history, main query . Output: next plan .
-Answering Task: Input: (encoded ), answering instruction , full subquery-triples history, main query . Output: final answer .
-The model used in RAS is Graph LLM, an architecture adapted from prior work [38, 11].
-It is trained to perform both planning and answering tasks via standard next-token prediction, conditioned on the instruction type and input structure.
-To the best of our knowledge, this is the first application of Graph LLMs to general language generation tasks such as open question answering.
+Planning: Input: current encoded question-specific KG , planning instruction (shown in Figure 15), subquery-triples history (), main question . Output: next plan .
+Answering: Input: final encoded question-specific KG , answering instruction (shown in Figure 16), subquery-triples history (), main question . Output: final answer .
 4 Experiments
+4.1 Settings
 Training Data & Setting.
-We develop HotpotQA-SUBQ, a dataset constructed based on HotpotQA [1], to train our model’s action planning and answering capabilities.
-Our dataset creation begins with document filtering: using Claude-3.5-Sonnet [48], we identify and retain only the supporting documents necessary for answering the main query, removing irrelevant content.
-For each supporting document , we then iteratively generate a subquery , considering the main query, previous subqueries, and supporting documents.
-During iteration , when more supporting documents remain, we create training samples with input and output label “[SUBQ] ”, where represents triples extracted from document , and is the main query.
+We develop HotpotQA-SUBQ, a dataset constructed based on HotpotQA (Yang et al., 2018), to train our model’s action planning and answering capabilities.
+Our dataset creation begins with document filtering: using Claude-3.5-Sonnet (Anthropic, 2024), we identify and retain only the supporting documents necessary for answering the main question, removing irrelevant content.
+For each supporting document , we then iteratively generate a subquery , considering the main question, previous subqueries, and supporting documents.
+During iteration , when more supporting documents remain, we create training samples with input and output label “[SUBQ] ”, where represents triples extracted from document , and is the main question.
 For the final supporting document, we label the sample as “[SUFFICIENT]”.
 To identify queries that can be answered directly, we test our base LLM (LLaMA-2-7B) on HotpotQA’s main queries without context.
-For correctly answered queries, we create training samples with the main query as input and “[NO_RETRIEVAL]” as the output label.
+For correctly answered queries, we create training samples with the main question as input and “[NO_RETRIEVAL]” as the output label.
 Additionally, to ensure fair comparison with existing approaches, we incorporate the subset of Arc-Easy (2,147 samples) and ASQA (3,897 samples) from Self-RAG’s training data, resulting in 208k training samples in total.
-We place detailed training data processing, dataset statistics, and data samples in Appendix B.
-For efficient inference, we train a text-to-triples model on the WikiOFGraph dataset [44] using LLaMA-3.2-3B-Instruct as the base model (see Appendix C), and deploy it using vLLM for optimized runtime performance.
-We present our hyperparameter study of each component in Appendix H.
+We place detailed training data processing, dataset statistics, and data samples in Appendix C.
+For efficient inference, we train a text-to-triples model on the WikiOFGraph dataset (Kim et al., 2024a) using LLaMA-3.2-3B-Instruct as the base model (see Appendix D), and deploy it using vLLM for optimized runtime performance.
+We present our hyperparameter study of each component in Appendix J.
 Knowledge Sources.
-Due to the large sizes of wiki corpora, we employ faiss [49] for efficient vector searching over the dense index.
-Following the Self-RAG [30], we utilize the Wikipedia 2018 [50] by default, while specifically using the Wikipedia 2020 for PopQA to access more recent information.
+We employ faiss (Douze et al., 2024) for efficient vector searching over the dense index.
+Following the Self-RAG (Asai et al., 2023), we utilize the Wikipedia 2018 (Izacard et al., 2023) by default, while specifically using the Wikipedia 2020 for PopQA to access more recent information.
 To optimize retrieval efficiency, we partition the index into five segments.
 Test Datasets & Metrics & Compared Baselines.
-We conduct comprehensive evaluations on diverse knowledge-intensive tasks following previous studies [30, 31].
-The evaluation encompasses three categories of datasets: (1) open-domain short-form generation datasets: TriviaQA [51], PopQA [52], and 2WikiMultihopQA [53]; (2) closed-set task datasets: PubHealth [54] and ARC-Challenge [55]; and (3) long-form generation datasets: ALCE-ASQA [56, 57] and ELI5 [58].
-For evaluation metrics, we maintain consistency with prior work [30, 52, 31], employing “golden match” accuracy for PopQA and TriviaQA, token-level F1 score for 2WikiMultihopQA, accuracy for PubHealth and ARC-Challenge, and ROUGE-LSum alongside MAUVE score [59] for ASQA and ELI5.
-Our comparative analysis includes three baseline categories: models without retrieval augmentation, incorporating Claude 3.5 Sonnet as a state-of-the-art closed-source baseline; models with single retrieval over top-5 documents, including Claude 3.5 Sonnet, and SuRe [25], a leading retrieve-and-summarize method; and models with self-reflective retrieval, including leading approaches Self-RAG [30], RPG [31], and Search-R1 [60].
-We place more details of metrics/datasets in Appendix D.
-| Short-form | Closed-set | Long-form Generation | |||||||
-| Model | TQA | 2WQA | PopQA | Pub | ARC | ASQA | ELI5 | ||
-| (acc) | (F1) | (acc) | (acc) | (acc) | (rouge) | (mauve) | (rouge) | (mauve) | |
-| Baselines w/o Retrieval | |||||||||
-| Llama2 | 30.5 | 18.9 | 14.7 | 34.2 | 21.8 | 15.3 | 19.0 | 18.3 | 32.4 |
-| Alpaca | 54.5 | – | 23.6 | 49.8 | 45.0 | 29.4 | 61.7 | – | – |
-| Llama2 | 38.5 | 20.2 | 14.7 | 29.4 | 29.4 | 12.4 | 16.0 | 18.2 | 41.4 |
-| Alpaca | 61.3 | – | 24.4 | 55.5 | 54.9 | 32.0 | 70.6 | – | – |
-| Llama3 | 56.1 | 21.2 | 26.7 | 33.2 | 42.2 | 17.6 | 25.0 | 18.2 | 39.7 |
-| ChatGPT | 74.3 | 24.8 | 29.3 | 70.1 | 75.3 | 36.2 | 68.8 | 22.8 | 32.6 |
-| Sonnet-3.5 | 78.4 | 40.0 | 30.2 | 83.7 | 88.5 | 37.0 | 39.1 | 21.8 | 26.5 |
-| Baselines w/ Retrieval (#docs=5 by default) | |||||||||
-| Llama2 | 42.5 | 21.0 | 38.2 | 30.0 | 48.0 | 22.1 | 32.0 | 18.6 | 35.3 |
-| Alpaca | 64.1 | – | 46.7 | 40.2 | 48.0 | 33.3 | 57.9 | – | – |
-| Llama2 | 47.0 | 31.2 | 45.7 | 30.2 | 26.0 | 20.5 | 24.7 | 18.6 | 42.3 |
-| Llama3 | 60.4 | 33.4 | 48.6 | 36.5 | 40.1 | 23.9 | 52.1 | 18.8 | 40.7 |
-| Alpaca | 66.9 | – | 46.1 | 51.1 | 57.6 | 36.7 | 56.6 | – | – |
-| RECOMP [61] | – | 32.4 | – | – | – | 36.5 | 76.0 | – | – |
-| ChatGPT | 65.7 | 32.8 | 50.8 | 54.7 | 75.3 | 39.9 | 79.7 | 20.6 | 57.2 |
-| Sonnet-3.5 | 69.1 | 41.9 | 51.5 | 49.1 | 88.6 | – | – | – | – |
-| Sonnet-3.5 | 72.5 | 53.7 | 57.3 | 53.9 | 87.1 | 38.8 | 61.6 | 20.2 | 32.3 |
-| SuRe [25] | 51.2 | 20.6 | 39.0 | 36.2 | 52.7 | 35.8 | 76.2 | 16.1 | 26.6 |
-| SuRe | 72.3 | 38.1 | 53.6 | 57.2 | 79.6 | 36.0 | 74.2 | 19.2 | 51.6 |
-| SuRe | 76.8 | 37.6 | 41.2 | 62.8 | 91.6 | 30.2 | 69.9 | 15.4 | 27.2 |
-| Baselines w/ Self-Reflective Retrieval | |||||||||
-| Self-RAG [30] | 66.4 | 25.1 | 54.9 | 72.4 | 67.3 | 35.7 | 74.3 | 17.9 | 35.6 |
-| Self-RAG | 69.3 | – | 55.8 | 74.5 | 73.1 | 37.0 | 71.6 | – | – |
-| RPG [31] | 65.1 | 33.6 | 56.0 | 73.4 | 65.4 | 37.6 | 84.4 | 19.1 | 46.4 |
-| Search-R1 [60] | 63.8 | 35.2 | 45.7 | 69.4 | 64.4 | 35.2 | 72.3 | 18.4 | 37.1 |
-| Retrieval-And-Structuring (RAS) | |||||||||
-| RAS | 77.6 | 57.7 | 62.3 | 71.3 | 93.9 | 39.1 | 70.5 | 23.3 | 37.7 |
-| RAS | 73.8 | 44.2 | 57.7 | 77.6 | 71.4 | 37.6 | 96.2 | 20.1 | 54.4 |
-| RAS | 72.7 | 42.1 | 58.3 | 74.7 | 68.5 | 37.2 | 95.2 | 19.7 | 47.8 |
+We conduct comprehensive evaluations on diverse knowledge-intensive tasks following previous studies (Asai et al., 2023; Lyu et al., 2024b).
+The evaluation encompasses three categories of datasets: (1) open-domain short-form generation datasets: TriviaQA (Joshi et al., 2017), PopQA (Mallen et al., 2022), and 2WikiMultihopQA (Ho et al., 2020); (2) closed-set task datasets: PubHealth (Zhang et al., 2023a) and ARC-Challenge (Clark et al., 2018); and (3) long-form generation datasets: ALCE-ASQA (Gao et al., 2023; Stelmakh et al., 2022) and ELI5 (Fan et al., 2019).
+For evaluation metrics, we maintain consistency with prior work (Asai et al., 2023; Mallen et al., 2022; Lyu et al., 2024b), employing “golden match” accuracy for PopQA and TriviaQA, token-level F1 score for 2WikiMultihopQA, accuracy for PubHealth and ARC-Challenge, and ROUGE-LSum alongside MAUVE score (Pillutla et al., 2021) for ASQA and ELI5.
+Our comparative analysis includes three baseline categories: models without retrieval augmentation, incorporating Claude 3.5 Sonnet as a state-of-the-art closed-source baseline; models with single retrieval over top-5 documents, including Claude 3.5 Sonnet, and SuRe (Kim et al., 2024b), a leading retrieve-and-summarize method; and models with self-reflective retrieval, including leading approaches Self-RAG (Asai et al., 2023), RPG (Lyu et al., 2024b) .
+We place more details of datasets and metrics in Appendices E.1 and E.2, respectively.
+| Short-form | Closed-set | Long-form Generation | ||||||||
+| Model/Method | TQA | 2WQA | PopQA | Pub | ARC | ASQA | ELI5 | |||
+| (acc) | (F1) | (acc) | (acc) | (acc) | (rouge) | (mauve) | (rouge) | (mauve) | ||
+| Closed-source | w/o Retrieval | |||||||||
+| ChatGPT | 74.3 | 24.8 | 29.3 | 70.1 | 75.3 | 36.2 | 68.8 | 22.8 | 32.6 | |
+| Sonnet-3.5 | 78.4 | 40.0 | 30.2 | 83.7 | 88.5 | 37.0 | 39.1 | 21.8 | 26.5 | |
+| w/ Single Retrieval (#docs=5) | ||||||||||
+| Sonnet-3.5 | 69.1 | 41.9 | 51.5 | 49.1 | 88.6 | n/a | n/a | n/a | n/a | |
+| Sonnet-3.5 | 72.5 | 53.7 | 57.3 | 53.9 | 87.1 | 38.8 | 61.6 | 20.2 | 32.3 | |
+| SuRe(Kim et al., 2024b) | 72.3 | 38.1 | 53.6 | 57.2 | 79.6 | 36.0 | 74.2 | 19.2 | 51.6 | |
+| SuRe | 76.8 | 37.6 | 41.2 | 62.8 | 91.6 | 30.2 | 69.9 | 15.4 | 27.2 | |
+| w/ Self-Reflective Retrieval | ||||||||||
+| ReAct (Yao et al., 2023) | 73.4 | 53.7 | 55.0 | 62.2 | 89.2 | 38.8 | 61.6 | 20.2 | 32.3 | |
+| IRCoT (Trivedi et al., 2023) | 74.7 | 54.9 | 53.2 | 59.4 | 92.0 | 38.8 | 61.6 | 20.2 | 32.3 | |
+| Retrieval-And-Structuring (ours) | ||||||||||
+| RAS | 77.6 | 57.7 | 62.3 | 71.3 | 93.9 | 39.1 | 70.5 | 23.3 | 37.7 | |
+| Open-source | w/o Retrieval | |||||||||
+| Llama2 | 30.5 | 18.9 | 14.7 | 34.2 | 21.8 | 15.3 | 19.0 | 18.3 | 32.4 | |
+| Llama2 | 38.5 | 20.2 | 14.7 | 29.4 | 29.4 | 12.4 | 16.0 | 18.2 | 41.4 | |
+| Llama3 | 56.1 | 21.2 | 26.7 | 33.2 | 42.2 | 17.6 | 25.0 | 18.2 | 39.7 | |
+| w/ Single Retrieval (#docs=5) | ||||||||||
+| Llama2 | 42.5 | 21.0 | 38.2 | 30.0 | 48.0 | 22.1 | 32.0 | 18.6 | 35.3 | |
+| Llama2 | 47.0 | 31.2 | 45.7 | 30.2 | 26.0 | 20.5 | 24.7 | 18.6 | 42.3 | |
+| Llama3 | 60.4 | 33.4 | 48.6 | 36.5 | 40.1 | 23.9 | 52.1 | 18.8 | 40.7 | |
+| SuRe (Kim et al., 2024b) | 51.2 | 20.6 | 39.0 | 36.2 | 52.7 | 35.8 | 76.2 | 16.1 | 26.6 | |
+| w/ Self-Reflective Retrieval | ||||||||||
+| Self-RAG (Asai et al., 2023) | 66.4 | 25.1 | 54.9 | 72.4 | 67.3 | 35.7 | 74.3 | 17.9 | 35.6 | |
+| Self-RAG | 69.3 | 26.9 | 55.8 | 74.5 | 73.1 | 37.0 | 71.6 | 18.7 | 38.5 | |
+| RPG (Lyu et al., 2024b) | 65.1 | 33.6 | 56.0 | 73.4 | 65.4 | 37.6 | 84.4 | 19.1 | 46.4 | |
+| ReAct (Yao et al., 2023) | 64.0 | 25.0 | 42.7 | 52.4 | 59.0 | 22.1 | 32.0 | 18.6 | 35.3 | |
+| IRCoT (Trivedi et al., 2023) | 61.5 | 27.6 | 44.3 | 59.6 | 61.6 | 22.1 | 32.0 | 18.6 | 35.3 | |
+| Retrieval-And-Structuring (ours) | ||||||||||
+| RAS | 72.7 | 42.1 | 58.3 | 74.7 | 68.5 | 37.2 | 95.2 | 19.7 | 47.8 | |
+| RAS | 73.8 | 44.2 | 57.7 | 77.6 | 71.4 | 37.6 | 96.2 | 20.1 | 54.4 | |
 Inference Setting.
-We evaluate RAS using both our trained open-source models (RAS, based on LLaMA-2-7B/LLaMA-3-8B and a Graph Transformer encoder [65]) and the closed-source Claude-3.5-Sonnet model under varied inference settings.
-For open-source models (RAS), we follow prior work [30, 31] and adopt zero-shot inference across all datasets.
-For the closed-source model (RAS), we apply few-shot inference with two exemplars for ASQA and ELI5, while using zero-shot inference for all other datasets.
-More details are provided in Appendix E.1 and E.2.
-During inference, if the initial planning output is not [NO_RETRIEVAL], we utilize the main query as the initial subquery .
-For PopQA and TriviaQA evaluation, we follow established settings [30, 66], incorporating top-five web search engine results as initial retrieved context .
-For ASQA and ELI5, we maintain methodological consistency with prior work [31, 30, 56], utilizing their predetermined five-document context.
-Under these conditions, we omit plan generation and text retrieval phases, implementing static inference [30] with five fixed iterations.
+We evaluate RAS using both our trained open-source models (RAS, based on LLaMA-2-7B/LLaMA-3-8B and a Graph Transformer encoder (Shi et al., 2020)) and the closed-source Claude-3.5-Sonnet & Claude-4.5-Sonnet (Anthropic, 2025) under varied inference settings.
+For open-source models (RAS), we follow prior work (Asai et al., 2023; Lyu et al., 2024b) and adopt zero-shot inference across all datasets.
+For the closed-source model (RAS, RAS), we apply few-shot inference with two exemplars for ASQA and ELI5, while using zero-shot inference for all other datasets.
+More details are provided in Appendices F.1 and F.2.
+For PopQA and TriviaQA evaluation, we follow established settings (Asai et al., 2023; Luo et al., 2023a), incorporating top-five web search engine results as initial retrieved context .
+For ASQA and ELI5, we maintain methodological consistency with prior work (Lyu et al., 2024b; Asai et al., 2023; Gao et al., 2023), utilizing their predetermined five-document context.
+Under these conditions, we omit plan generation and text retrieval phases, implementing static inference (Asai et al., 2023) with five fixed iterations.
 For remaining datasets, we establish a maximum iteration count of five.
-Following previous studies, we employ Contriever-MS MARCO [67] as the primary dense retriever, with BM25 [68] serving as the retrieval mechanism for 2WikiMultihopQA.
+Following previous studies, we employ Contriever-MS MARCO (Izacard et al., 2021) as the primary dense retriever, with BM25 (Robertson et al., 2009) serving as the retrieval mechanism for 2WikiMultihopQA.
 Across all retrieval processes, we maintain a consistent top- document selection of five.
-We present more details of inference settings in Appendix E.
-4.1 Main Results
-Our performance evaluation, presented in Table 1, demonstrates that the LLaMA-2-7B/LLaMA-3-8B model fine-tuned with RAS outperforms existing open-source solutions, including Self-RAG [30] and RPG [31].
-Notably, compared to the previous SOTA models, RAS shows a 9.7% improvement in short-form question-answering and a 7.9% gain in long-form generation tasks.
-Additionally, when applied to closed-source Sonnet-3.5 model, RAS consistently achieves superior results compared to single retrieval RAG approaches, including retrieve-and-summarize approach SuRe [25].
+We present more details of inference settings in Appendix F.
+4.2 Results
+Main Result.
+Our performance evaluation, presented in Table 1, demonstrates that the LLaMA-2-7B/LLaMA-3-8B model fine-tuned with RAS outperforms existing SFT-based open-source solutions, including Self-RAG (Asai et al., 2023) and RPG (Lyu et al., 2024b).
+Notably, compared to the earlier SOTA models, RAS shows a 9.7% improvement in short-form question-answering and a 7.9% gain in long-form generation tasks.
+Additionally, when applied to Claude-3.5-Sonnet, RAS consistently achieves superior results compared to single retrieval RAG approaches, including retrieve-and-summarize approach SuRe (Kim et al., 2024b).
 We find that sometimes (e.g., on TriviaQA and PubHealth) single-hop retrieval could not boost LLM’s performance, and even makes it worse, which demonstrates the necessity of on-demand retrieval, aligning with previous findings.
-Our “role-swapping” study in Figure 3, where Sonnet-3.5 and RAS alternately perform the planning and answering tasks, reveals that performance is primarily bounded by the answering capability.
-For example, when Sonnet-3.5 handles planning and RAS handles answering, the system achieves 62.4% accuracy on the ARC-Challenge dataset.
-However, when Sonnet-3.5 is responsible for both planning and answering, accuracy increases to 93.9%.
-Notably, despite having 60 fewer parameters, our trained RAS achieves planning performance comparable to Sonnet-3.5, and even outperforms it in some cases.
-Also, for long-form generation, our trained model achieved notably high MAUVE scores, which can be majorly attributed to our “text-to-triple” step, as demonstrated by our ablation study in Table 2.
-4.2 Ablations & More Analysis
+Although RAS integrates planning and answering in a unified framework, these components can be decoupled for greater flexibility.
+Our “role-swapping” study in Figure 4 demonstrates that performance is primarily constrained by answering capability rather than planning.
+When Sonnet-3.5 handles planning while RAS performs answering, the system achieves 62.4% accuracy on ARC-C, compared to 93.9% when Sonnet-3.5 performs both tasks.
+Despite having 60 fewer parameters, RAS achieves planning performance comparable to (and sometimes exceeding) Sonnet-3.5.
+| TQA | 2WQA | Pub | ASQA | ||
+| (acc) | (F1) | (acc) | (rg) | (mv) | |
+| RAS | 72.7 | 42.1 | 74.7 | 37.2 | 95.2 |
+| Training Phase | |||||
+| w/o GraphEncode | 70.2 | 38.4 | 66.4 | 33.1 | 85.0 |
+| w/o LoRA | 71.5 | 37.8 | 54.8 | 32.8 | 84.8 |
+| w/o Text-to-Triple | 70.4 | 38.2 | 71.4 | 36.2 | 73.8 |
+| w/o Multi-Task | 68.6 | 39.2 | 65.5 | 36.7 | 88.9 |
+| Inference Phase | |||||
+| w/o Retrieval | 56.9 | 27.4 | 69.0 | 31.3 | 70.6 |
+| w/o GraphEncode | 68.8 | 38.7 | 67.3 | 36.5 | 93.6 |
+| w/o Planning | 66.7 | 37.8 | 71.5 | 37.2 | 95.2 |
 We analyze the impact of each component of RAS in Table 2 and provide detailed discussion below.
@@
 Also, when turning off retrieval, the performance degradation is more obvious, due to the knowledge-intensive nature of those datasets.
-| TQA | 2WQA | Pub | ASQA | ||
-| (acc) | (F1) | (acc) | (rg) | (mv) | |
-| RAS | 72.7 | 42.1 | 74.7 | 37.2 | 95.2 |
-| Training Phase | |||||
-| No GraphEncode | 70.2 | 38.4 | 66.4 | 33.1 | 85.0 |
-| No LoRA | 71.5 | 37.8 | 54.8 | 32.8 | 84.8 |
-| No Text-to-Triple | 70.4 | 38.2 | 71.4 | 36.2 | 73.8 |
-| No Multi-Task | 68.6 | 39.2 | 65.5 | 36.7 | 88.9 |
-| Inference Phase | |||||
-| No Retrieval | 56.9 | 27.4 | 69.0 | 31.3 | 70.6 |
-| No GraphEncode | 68.8 | 38.7 | 67.3 | 36.5 | 93.6 |
-| No Planning | 66.7 | 37.8 | 71.5 | 37.2 | 95.2 |
 Effect of Graph Construction and Encoding.
@@
 Using only graph token optimization without LoRA leads to substantial degradation (-11.8% on average).
-A similar observation can be made for “No Multi-Task” (e.g., 12.3% accuracy degradation on PubHealth), indicating the significance of jointly training the model on both action planning and answer generation tasks rather than optimizing for each task separately, supporting findings from prior work [31].
+A similar observation can be made for “No Multi-Task” (e.g., 12.3% accuracy degradation on PubHealth), indicating the significance of jointly training the model on both action planning and answer generation tasks rather than optimizing for each task separately, supporting findings from prior work (Lyu et al., 2024b).
 The complementary effects suggest that while graph-based knowledge representation is valuable, it needs to be combined with careful parameter tuning and multi-task learning to achieve optimal performance.
@@
 However, we observed an exception with the ELI5 dataset, where performance were inconsistent.
-This irregularity can be attributed to the inclusion of ASQA training data in our training set, following established setting from previous research [30, 31].
+This irregularity can be attributed to the inclusion of ASQA training data in our training set, following established setting from previous research (Asai et al., 2023; Lyu et al., 2024b).
 Among our test datasets, ASQA and ELI5 are unique in requiring long-form response generation.
 The periodic decline in ELI5 performance suggests that the model’s response generation began to align more closely with ASQA’s training data distribution, potentially at the expense of ELI5’s distinct characteristics.
-| Triple Extractor | TQA | 2WQA | PopQA |
-| Flan-T5-Large | 70.3 | 40.7 | 56.7 |
-| LLaMA-3.2-3B | 72.7 | 42.1 | 58.3 |
-| Claude-3.5-Sonnet | 73.8 | 44.5 | 60.1 |
-Impact of Triple Extraction Quality We further examine how the quality of triple extraction influences the overall effectiveness of RAS.
-Specifically, we test three text-to-triples models: Flan-T5-Large, LLaMA-3.2-3B model and Claude-3.5-Sonnet where the first two are trained on WikiOfGraph data.
-As shown in Table 3, higher-quality triple extraction leads to consistent performance gains across all evaluated datasets.
-Notably, using Claude-3.5-Sonnet triples improves TriviaQA accuracy from 72.7% to 73.8%, boosts 2WikiMultihopQA F1 from 42.1 to 44.5, and raises PopQA accuracy from 58.3% to 60.1%.
-These results underscore that while RAS achieves strong performance even with lightweight extraction models, its effectiveness can be further enhanced by supplying more precise and complete structured knowledge during inference.
+| Triple Extractor | Performance | Efficiency | ||
+| TQA | 2WQA | PopQA | (tokens/s) | |
+| Flan-T5-Large | 70.3 | 40.7 | 56.7 | 166.0 |
+| LLaMA-3.2-3B | 72.7 | 42.1 | 58.3 | 4,885.3 |
+| Claude-3.5-Sonnet | 73.8 | 44.5 | 60.1 | 68.2 |
+Impact of Triple Extractor Selection We examine how triple extraction quality affects RAS performance using three models: Flan-T5-Large, LLaMA-3.2-3B, and Claude-3.5-Sonnet.
+As shown in Table 3, higher-quality extraction consistently improves results across datasets, with Claude-3.5-Sonnet achieving the best accuracy but at significantly lower efficiency (1.5e-2 sec/token vs 2.0e-4 for LLaMA-3.2-3B).
+These results demonstrate the accuracy-efficiency tradeoff in triple extraction, with LLaMA-3.2-3B providing the optimal balance for our experiments.
 5 Conclusion
-We presented RAS (Retrieval-And-Structuring), a novel framework that improves knowledge-intensive language generation by interleaving iterative retrieval with dynamic knowledge graph construction.
-RAS introduces key innovations, including structure-aware action planning and graph-augmented answering, enabling models to reason over progressively organized information tailored to each query.
-Extensive experiments across seven benchmarks demonstrate that RAS consistently outperforms existing methods, particularly in tasks requiring multi-step reasoning and factual consistency.
-Furthermore, the modular, graph-based architecture of RAS allows seamless integration with external knowledge sources and human feedback, making it well-suited for high-stakes domains where transparency and adaptability are essential.
-Limitations and directions for future work are discussed in Appendix A.
+We presented RAS, a framework that dynamically constructs question-specific knowledge graphs through iterative retrieval and structured reasoning.
+Experiments across seven benchmarks show consistent improvements of up to 6.4% and 7.0% with open-source and proprietary LLMs, respectively.
+The modular architecture enables transparent reasoning and seamless integration with external knowledge sources.
+Limitations and future work are discussed in Appendix A.
+The statement of the LLM usage is placed in Appendix B.
+Broader Impacts, safeguards, and used assets are discussed in Appendix K.
+Our data and code can be found at: https://github.com/pat-jj/RAS.
+Ethics Statement
+This work does not involve human subjects, personal data, or sensitive user information.
+Our research focuses on algorithmic improvements for knowledge-intensive language model reasoning.
+We are mindful of potential ethical concerns, such as risks of misinformation or biased outputs when deploying retrieval-augmented LLMs.
+To mitigate these risks, we provide transparent descriptions of our methods, ablation studies on failure cases, and open-sourcing of code and models to facilitate community scrutiny.
+We also discuss broader impacts and safeguards in Appendix K.
+Acknowledgment
+Research was supported in part by the AI Institute for Molecular Discovery, Synthetic Strategy, and Manufacturing: Molecule Maker Lab Institute (MMLI), funded by U.S.
+National Science Foundation under Awards No.
+2019897 and 2505932, NSF IIS 25-37827, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No.
+2118329.
+The research has used the Delta/DeltaAI advanced computing and data resource, supported in part by the University of Illinois Urbana-Champaign and through allocation #250851 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants OAC #2320345, #2138259, #2138286, #2138307, #2137603, and #2138296.
 Appendix A Limitations & Future Work
@@
 This could significantly enhance the system’s reasoning capabilities by combining dynamically constructed knowledge with curated structured information.
-We summarize the potential future directions of extension in Figure 5.
-Appendix B HotpotQA-SUBQ Dataset
-The source of our dataset is HotpotQA [1].111We use the version hotpot_train_v1.1 from https://github.com/hotpotqa/hotpot.
+Appendix B The Use of Large Language Models
+In this work, large language models (LLMs) were used solely for drafting and polishing the manuscript.
+Their role was limited to grammar correction, phrasing refinement, and enhancing readability.
+All research ideas, methodological innovations, theoretical analyses, and experimental work were entirely conceived, designed, and carried out by the authors.
+The authors take full responsibility for the scientific content of the paper.
+Appendix C HotpotQA-SUBQ Dataset
+The source of our dataset is HotpotQA (Yang et al., 2018).111We use the version hotpot_train_v1.1 from https://github.com/hotpotqa/hotpot.
 Our goal is to create a dataset that can let the model learn the capabilities of subquery generation and answering with intensive knowledge.
 We name the dataset as HotpotQA-SUBQ.
-The construction of HotpotQA-SUBQ includes three steps: (1) filter out irrelevant/unnecessary “supporting docs” (§B.1), (2) generate one subquery for each supporting document (§B.2), and (3) sample labeling (§B.3).
-B.1 Supporting Document Filtering
+The construction of HotpotQA-SUBQ includes three steps: (1) filter out irrelevant/unnecessary “supporting docs” (§C.1), (2) generate one subquery for each supporting document (§C.2), and (3) sample labeling (§C.3).
+C.1 Supporting Document Filtering
 To identify and filter out irrelevant supporting documents from HotpotQA, we employ an instruction-based approach using Claude 3.5 Sonnet, as shown in Figure 6.
@@
 This filtering step reduces noise in the training data and helps focus the model on truly relevant information during sub-query generation.
-B.2 Sub-query Generation
+C.2 Sub-query Generation
 For sub-query generation, we use the template as follows to generate one subquery per (filtered) document:
@@
 For iterative querying, we maintain a list of previously generated sub-queries to avoid redundancy and encourage progressive information gathering.
-B.3 Sample Labeling
+C.3 Sample Labeling
 The sample labeling process transforms the filtered and subquery-augmented HotpotQA examples into training instances for both the Planning and Answering components.
@@
 The complementary training objectives help develop robust reasoning capabilities while maintaining retrievability of supporting evidence.
-B.4 Dataset Statistics
+C.4 Dataset Statistics
 | Planning Data | Answering Data | ||
@@
 | # Edges | (Mean, Median, Max) | (21.2, 20, 116) | (21.3, 20, 160) |
-B.5 Data Examples
-Appendix C Training Details
-C.1 Text-to-Triples Model Training
-For our text-to-triple conversion model, we fine-tuned a Flan-T5-Large model [69]222https://huggingface.co/pat-jj/text2triple-flan-t5 and a LLaMA-3.2-3B-Instruct model [70]333https://huggingface.co/pat-jj/text2graph-llama-3.2-3b to transform raw text passages into structured knowledge triples.
+C.5 Data Examples
+Appendix D Training Details
+D.1 Text-to-Triples Model Training
+For our text-to-triple conversion model, we fine-tuned a Flan-T5-Large model (Chung et al., 2024) and a LLaMA-3.2-3B-Instruct model (Grattafiori et al., 2024) to transform raw text passages into structured knowledge triples.
 The model processes input text sequences up to 512 tokens in length and generates structured triples in a standardized format “(S subject P predicate O object)”.
-Our training dataset WikiOFGraph [44], curated by ChatGPT, comprises 5,851,776 text-triple pairs, with an additional 5,000 samples reserved for validation.
+Our training dataset WikiOFGraph (Kim et al., 2024a), curated by ChatGPT, comprises 5,851,776 text-triple pairs, with an additional 5,000 samples reserved for validation.
 Each training instance consists of a natural language text passage paired with its corresponding set of comma-separated triples.
@@
 The training of this model was conducted on eight NVIDIA RTX 6000 with 48GB memory.
-C.2 Multitask Training of Planning and Answering Capabilities
-We employ Graph LLM [38, 11] as our foundational model architecture and utilize our constructed HotpotQA-SUBQ dataset to train a unified model capable of performing both graph-conditioned action planning and answer generation through multitask learning.
-Since both tasks share an identical architectural foundation, as illustrated in Figure 1, we differentiate their functionality through specialized instruction sets.
+D.2 Multitask Training of Planning and Answering Capabilities
+We employ Graph LLM (Perozzi et al., 2024; He et al., 2024) as our foundational model architecture and utilize our constructed HotpotQA-SUBQ dataset to train a unified model capable of performing both graph-conditioned action planning and answer generation through multitask learning.
+Since both tasks share an identical architectural foundation, as illustrated in Figure LABEL:fig:overviewLABEL:fig:overview, we differentiate their functionality through specialized instruction sets.
 For Planning, we employ the following instruction template:
 For the Answering, we utilize this distinct instruction set:
-For graph encoding, we implement Graph Transformer [65], selected for its robust capability in handling non-fully-connected graphs—a common characteristic of text-extracted triples.
-Our base language models comprise LLaMA-2-7B and LLaMA-3-8B, chosen both to maintain consistency with previous research [30, 31] and to investigate our framework’s performance scaling across different model capacities.
-Our implementation of GraphLLM differs from G-Retriever [11] primarily due to the distinct nature of our graph structures.
+For graph encoding, we implement Graph Transformer (Shi et al., 2020), selected for its robust capability in handling non-fully-connected graphs—a common characteristic of text-extracted triples.
+Our base language models comprise LLaMA-2-7B and LLaMA-3-8B, chosen both to maintain consistency with previous research (Asai et al., 2023; Lyu et al., 2024b) and to investigate our framework’s performance scaling across different model capacities.
+Our implementation of GraphLLM differs from G-Retriever (He et al., 2024) primarily due to the distinct nature of our graph structures.
 While G-Retriever operates on single interconnected graphs, our framework processes multiple potentially disconnected subgraphs, each corresponding to different subqueries.
@@
 The maximum sequence lengths are set to 300 tokens for generation and 2,500 tokens for input, with training conducted in BFloat16 precision.
-Appendix D Evaluation Datasets & Metrics
-D.1 Test Datasets
-We evaluate RAS on diverse benchmark datasets spanning short-form QA, closed-set tasks, and long-form generation in the zero-shot setting, aligning with [30, 31]. Below we describe each dataset:
+Appendix E Evaluation Datasets & Metrics
+E.1 Test Datasets
+We evaluate RAS on diverse benchmark datasets spanning short-form QA, closed-set tasks, and long-form generation in the zero-shot setting, aligning with (Asai et al., 2023; Lyu et al., 2024b).
+Below we describe each dataset:
 Short-form Generation Datasets:
-TriviaQA-unfiltered (TQA) [51]: A large-scale QA dataset containing 11,313 question-answer pairs in our test set. The questions are sourced from trivia enthusiasts and cover diverse topics.
-2WikiMultiHopQA (2WQA) [53]: A multi-hop question answering dataset (with 12,576 samples in test set) that requires models to combine information from multiple Wikipedia articles to answer questions.
-PopQA [52]: A dataset focusing on questions about long-tail entities, containing 1,399 queries where the monthly Wikipedia page views are less than 100.
+TriviaQA-unfiltered (TQA) (Joshi et al., 2017): A large-scale QA dataset containing 11,313 question-answer pairs in our test set.
+The questions are sourced from trivia enthusiasts and cover diverse topics.
+2WikiMultiHopQA (2WQA) (Ho et al., 2020): A multi-hop question answering dataset (with 12,576 samples in test set) that requires models to combine information from multiple Wikipedia articles to answer questions.
+PopQA (Mallen et al., 2022): A dataset focusing on questions about long-tail entities, containing 1,399 queries where the monthly Wikipedia page views are less than 100.
 These questions test models’ ability to handle queries about less common entities.
 Closed-set Tasks:
-PubHealth (Pub) [54]: A dataset for verifying health-related claims. Models must classify statements as "true" or "false" based on scientific evidence. The dataset contains 987 test samples.
-ARC-Challenge (ARC) [55]: A multiple-choice science question dataset designed to be challenging for models, requiring multi-hop reasoning and background knowledge.
+PubHealth (Pub) (Zhang et al., 2023a): A dataset for verifying health-related claims.
+Models must classify statements as "true" or "false" based on scientific evidence.
+The dataset contains 987 test samples.
+ARC-Challenge (ARC) (Clark et al., 2018): A multiple-choice science question dataset designed to be challenging for models, requiring multi-hop reasoning and background knowledge.
 The dataset contains 1,172 test samples.
 Long-form Generation Datasets:
-ASQA [57]: A long-form question answering dataset that requires models to generate comprehensive answers with proper citation of sources.
+ASQA (Stelmakh et al., 2022): A long-form question answering dataset that requires models to generate comprehensive answers with proper citation of sources.
 Models must balance information completeness with factual accuracy.
 The dataset contains 948 test samples.
-ELI5 [58]: A dataset derived from the "Explain Like I’m Five" subreddit, containing questions seeking straightforward explanations of complex topics. The dataset contains 1,000 test samples.
-D.2 Evaluation Metrics
+ELI5 (Fan et al., 2019): A dataset derived from the "Explain Like I’m Five" subreddit, containing questions seeking straightforward explanations of complex topics.
+The dataset contains 1,000 test samples.
+E.2 Evaluation Metrics
 We employ different evaluation metrics appropriate for each task category:
 Short-form Generation:
-For PopQA and TriviaQA, we use the "golden match" metric [30, 71, 8]444Implementation aligned with https://github.com/AkariAsai/self-rag/blob/main/retrieval_lm/metrics.py, where a prediction is considered correct if it contains any normalized version of the ground truth answers :
+For PopQA and TriviaQA, we use the "golden match" metric (Asai et al., 2023; Min et al., 2019; Guu et al., 2020)222Implementation aligned with https://github.com/AkariAsai/self-rag/blob/main/retrieval_lm/metrics.py, where a prediction is considered correct if it contains any normalized version of the ground truth answers :
 (9) where normalizes text by lowercasing, removing articles and punctuation.
-For 2WikiMultiHopQA, we follow RPG [31]555Implementation aligned with https://github.com/haruhi-sudo/RPG/blob/main/retriever/src/evaluation.py to use token-level F1 score between prediction and ground truth :
+For 2WikiMultiHopQA, we follow RPG (Lyu et al., 2024b)333Implementation aligned with https://github.com/haruhi-sudo/RPG/blob/main/retriever/src/evaluation.py to use token-level F1 score between prediction and ground truth :
 (10) where precision and recall are computed over normalized token overlap:
 (11) (12)
-Closed-set Tasks: For both PubHealth and ARC-Challenge, we use accuracy [30], computed as:
+Closed-set Tasks: For both PubHealth and ARC-Challenge, we use accuracy (Asai et al., 2023), computed as:
 where is the total number of examples.
-Long-form Generation: For ASQA and ELI5, we use multiple metrics666Implementation aligned with https://github.com/princeton-nlp/ALCE/blob/main/eval.py:
+Long-form Generation: For ASQA and ELI5, we use multiple metrics444Implementation aligned with https://github.com/princeton-nlp/ALCE/blob/main/eval.py:
 All scores are reported as percentages, multiplied by 100.
-Appendix E Inference & Evaluation Details
+Appendix F Inference & Evaluation Details
 We evaluated RAS using both closed-source and open-source language models.
@@
 For open-source evaluation, we tested with both LLaMA-2-7B and LLaMA-3-8B, as shown in the performance table (Table 1).
-For the open-source models, we first fine-tuned the GraphLLM architecture on our HotpotQA-SUBQ dataset (see Section C.2).
+For the open-source models, we first fine-tuned the GraphLLM architecture on our HotpotQA-SUBQ dataset (see Section D.2).
 We then conducted zero-shot knowledge-intensive generation tests.
@@
 Below, we detail our approach for each model type.
-E.1 Closed-Source Model Settings
+F.1 Closed-Source Model Settings
 For Text-to-Triple Conversion (Figure 18), we instruct the model to extract structured knowledge triples following specific formatting rules.
@@
 The prompt strictly enforces concise answer generation, requiring only the essential information without additional explanation or commentary.
-In addition, for ASQA and ELI5 with closed-source models (RAS and all the other baselines (e.g., SuRe) using Sonnet-3.5), we conduct few-shot inference with two in-context learning demonstrations, aligning with previous study’s [56] implementation.777https://github.com/princeton-nlp/ALCE/blob/main/run.py888Exemplars can be found at https://github.com/princeton-nlp/ALCE/tree/main/prompts
-E.2 Open-Source Model Settings
-For inference with open-source model, we use our GraphLLM trained by Hotpot-SUBQ (see Appendix C.2), we use 8 NVIDIA RTX 6000 Ada Generation with 48GB memory and CUDA version 12.4.
+In addition, for ASQA and ELI5 with closed-source models (RAS and all the other baselines (e.g., SuRe) using Sonnet-3.5), we conduct few-shot inference with two in-context learning demonstrations, aligning with previous study’s (Gao et al., 2023) implementation.555https://github.com/princeton-nlp/ALCE/blob/main/run.py666Exemplars can be found at https://github.com/princeton-nlp/ALCE/tree/main/prompts
+F.2 Open-Source Model Settings
+For inference with open-source model, we use our GraphLLM trained by Hotpot-SUBQ (see Appendix D.2), we use 8 NVIDIA RTX 6000 Ada Generation with 48GB memory and CUDA version 12.4.
 We use Python 3.10, PyTorch 2.1.2, and torch-geometric 2.6.1 throughout all experiments.
-For more efficient text retrieval, we split the dense index of the corpus into eight splits, and load them on eight GPUs with faiss-gpu 1.7.2.999All the Wikipedia dumps (2018, 2020) we used are downloaded from https://github.com/facebookresearch/atlas
-We set maximum new tokens as 100 for PopQA, TriviaQA, and 2WikiMultihopQA, as 50 for PubHealth and ARC-Challenge, and as 300 for ASQA and ELI5, aligning with previous study [30, 31].
+For more efficient text retrieval, we split the dense index of the corpus into eight splits, and load them on eight GPUs with faiss-gpu 1.7.2.777All the Wikipedia dumps (2018, 2020) we used are downloaded from https://github.com/facebookresearch/atlas
+We set maximum new tokens as 100 for PopQA, TriviaQA, and 2WikiMultihopQA, as 50 for PubHealth and ARC-Challenge, and as 300 for ASQA and ELI5, aligning with previous study (Asai et al., 2023; Lyu et al., 2024b).
 All the generation are configured with batch size of 1.
-As we always set the main query as the first subquery, we pre-extract the triples from the context at the first iteration using our text-to-triple model for batch triple extraction.
-E.3 Instructions
+As we always set the main question as the first subquery, we pre-extract the triples from the context at the first iteration using our text-to-triple model for batch triple extraction.
+F.3 Instructions
 We use the following task-specific instructions for zero-shot inference:
@@
 | ELI5 | Provide a paragraph-length response using simple words to answer the following question. [Long Form] |
-Appendix F Supplementary Results and Analysis
+Appendix G Cost & Efficiency Comparison with Other Graph-based RAG Methods
+Static vs. Dynamic Structuring.
+Static corpus-level graph methods such as GraphRAG (Edge et al., 2024) and HippoRAG (Gutiérrez et al., 2024) require building and maintaining large knowledge graphs offline.
+This involves millions of LLM calls for entity/relation extraction and summarization, plus heavy clustering and embedding computation.
+At the scale of Wikipedia 2018 (roughly 5M articles, 3B tokens), these pipelines incur substantial costs, days of GPU/CPU compute, and 50 GB storage overhead.
+In contrast, our RAS constructs question-specific knowledge graphs dynamically at inference, eliminating expensive offline indexing.
+The only extra step is lightweight triple extraction from retrieved passages, which we implement with a small LLaMA-3B model accelerated by vLLM (Kwon et al., 2023), achieving throughput of nearly 5K tokens/s (Table 3).
+Cost Estimation Methodology.
+We estimate costs based on empirical data from Microsoft’s GraphRAG analysis 888https://techcommunity.microsoft.com/blog/azure-ai-foundry-blog/graphrag-costs-explained-what-you-need-to-know/4207978 and scale to Wikipedia 2018.
+Cost estimates assume current API pricing and may vary significantly based on implementation details, model selection, and optimization strategies.
+Computation of Indexing Cost.
+We estimate costs using the following assumptions:
+Corpus size: Wikipedia 2018 3B tokens, or 5M articles with 600 tokens/article.
+GraphRAG pipeline: Based on Microsoft’s empirical analysis showing $0.000011 per word for GPT-4o-mini.
+For 3B tokens (2.25B words), base processing cost: $25k–$30k
+Entity/relation extraction with few-shot prompting increases token usage by 3–5×: $75k–$150k
+Community summarization for 10k communities at $0.02 per community: $200–$500
+Additional overhead (clustering, embeddings, storage): $10k–$25k
+Total GraphRAG: $85k–$175k for GPT-4o-mini; $250k–$500k for GPT-4o
+HippoRAG pipeline: More efficient extraction but still corpus-wide processing.
+Entity extraction: 3B tokens at reduced complexity $15k–$30k
+Synonym linking and embedding overhead: $5k–$15k
+Graph construction and indexing: $5k–$10k
+Total HippoRAG: $25k–$55k
+RAS pipeline: No offline corpus-wide indexing. Triple extraction uses local LLaMA-3B model.
+Offline indexing cost: $0 (no corpus preprocessing required)
+Runtime triple extraction: Local model inference (negligible marginal cost)
+Total RAS: $0 offline costs
+Cost Sensitivity Analysis.
+Costs vary significantly based on:
+Model selection: GPT-4o costs 3× more than GPT-4o-mini
+Chunk size: Smaller chunks (300 vs 1200 tokens) can reduce costs by 30–50%
+Prompt optimization: Well-tuned prompts can reduce token usage by 20–40%
+Processing parameters: Conservative estimates assume non-optimized settings
+Quantitative Comparison.
+Table 6 summarizes estimated costs with uncertainty ranges.
+| Method | Offline LLM Cost | Compute Time | Storage | Ready for RAG? |
+| GraphRAG | $85k–$500k (model-dependent) | >7 days (distributed) | 50–70 GB | Yes (community reports) |
+| HippoRAG | $25k–$55k (optimized extraction) | >3 days (cluster) | 30–50 GB | Yes (node-passage matrix) |
+| RAS (ours) | Negligible | Inference-time only | Small (triple cache) | Yes (dynamic graph) |
+Efficiency Visualization.
+We visualize the contrast in offline indexing cost ranges. RAS requires no precomputation, while GraphRAG and HippoRAG incur substantial upfront costs.
+Takeaway.
+RAS achieves the structured reasoning benefits of graphs without any offline indexing costs, compared to GraphRAG ($85k–$500k) and HippoRAG ($25k–$55k).
+By eliminating upfront investment requirements and enabling pay-per-query scaling through local model inference, RAS makes graph-enhanced RAG accessible for scenarios where large-scale preprocessing is impractical or cost-prohibitive.
+Appendix H Supplementary Results and Analysis
 Higher Performance with More Contexts?
@@
 We present two representative cases where the quality of retrieved structured knowledge significantly impacts final answer generation.
-These examples are distinct from the qualitative examples shown in Section G and highlight both successful and failed reasoning paths.
+These examples are distinct from the qualitative examples shown in Section I and highlight both successful and failed reasoning paths.
 Case 1: Accurate Reasoning with Well-Structured Graph
@@
 This transparency not only facilitates error analysis but also suggests concrete avenues for improving upstream components like triple extraction or subquery generation, making RAS a flexible foundation for future refinement.
-Appendix G Qualitative Examples
+Appendix I Qualitative Examples
 In this section, we show some qualitative examples and running examples of RAS.
-Appendix H Hyperparameter Study
+Appendix J Hyperparameter Study
 | Hyper-parameter | Studied Values |
@@
 | GraphLLM (Action Planner & Answerer) | |
-[…]
+| GNN Setting (for Graph Token) | |
 | GNN architecture | {GCN, GAT, Graph Transformer} |
@@
 | Projector output dimension | 4096 |
-[…]
+| LoRA Setting | |
 | LoRA rank (r) | {4, 8, 16, 32} |
@@
 | LoRA dropout | {0.01, 0.05, 0.1, 0.15} |
-[…]
+| General Setting | |
 | Learning rate | {1e-6, 2e-6, 5e-5, 1e-4} |
@@
 | Max text length | 2500 |
-| Max new tokens | Task-specific (see Appendix E.2) |
+| Max new tokens | Task-specific (see Appendix F.2) |
 | Others | |
 | Dense retrieval top- | 5 |
-Appendix I Broader Impacts, Safeguards, and Assets
+Appendix K Broader Impacts, Safeguards, and Assets
 Broader Impacts.
 The RAS framework presents several promising impacts for both research and practical applications.
-By transforming unstructured retrieved content into structured, query-specific knowledge graphs, RAS improves reasoning transparency and factual reliability—qualities that are especially valuable in education, scientific research, and technical writing.
+By transforming unstructured retrieved content into structured, question-specific knowledge graphs, RAS improves reasoning transparency and factual reliability—qualities that are especially valuable in education, scientific research, and technical writing.
 The ability to explicitly organize retrieved knowledge into interpretable structures also opens up new possibilities for human-in-the-loop AI systems, enabling users to trace and verify model reasoning paths.